### v2

# On the convergence of loss and uncertainty-based active learning algorithms

Abstract
We consider the convergence rates of loss and uncertainty-based active learning algorithms under various assumptions. Firstly, we establish a set of conditions that ensure convergence rates when applied to linear classifiers and linearly separable datasets. This includes demonstrating convergence rate guarantees for loss-based sampling with various loss functions. Secondly, we introduce a framework that allows us to derive convergence rate bounds for loss-based sampling by leveraging known convergence rate bounds for stochastic gradient descent algorithms. Lastly, we propose a new algorithm that combines point sampling and stochastic Polyak’s step size. We establish a condition on the sampling process, ensuring a convergence rate guarantee for this algorithm, particularly in the case of smooth convex loss functions. Our numerical results showcase the efficiency of the proposed algorithm.
1 Introduction
In practice, when training machine learning models for prediction tasks (classification or regression), one often has an abundance of unlabeled data available, but obtaining the corresponding labels may come at a high cost. This may especially be the case in fields like computer vision, natural language processing, and speech recognition. Active learning algorithms are designed to efficiently learn a prediction model by employing a label acquisition, with the goal of minimizing the number of labels used to train an accurate prediction model.
Various label acquisition strategies have been proposed, each aiming to select informative points for the underlying model training task. For instance, active learning strategies include query-by-committee (Seung et al., 1992), expected model change (Settles et al., 2007), expected error reduction (Roy and McCallum, 2001), expected variance reduction (Wang et al., 2016), and mutual information maximization (Kirsch et al., 2019; Kirsch and Gal, 2022).
A common label acquisition strategy involves estimating uncertainty, which can be viewed as self-disagreement about predictions made by a given model. Algorithms using an uncertainty acquisition strategy are referred to as uncertainty-based active learning algorithms. Recently, a loss-based active learning approach gained attention in research Yoo and Kweon (2019), Lahlou et al. (2022), Nguyen et al. (2021), Luo et al. (2021), and is now applied at scale in industry, such as for training integrity violation classifiers at Meta. This method involves selecting points for which there is a disagreement between the prediction model and an oracle, as measured by a loss function. Since the true loss of a data point is unknown prior to the acquisition of the label, in practice it is estimated using supervised learning. Loss-based sampling aligns with the spirit of the perceptron algorithm (Rosenblatt, 1958), which updates the model only for falsely-classified points.
Convergence guarantees for some uncertainty sampling strategies have recently been established, such as for margin of confidence sampling (Raj and Bach, 2022). As for loss-based active learning algorithms, there are only limited results on their convergence properties, as they have been started to be studied only recently. e.g. Liu and Li (2023).
The primary focus of this paper is to establish convergence guarantees for loss and uncertainty-based learning algorithms. We delve into stream-based algorithms, where model training entails the use of a stochastic gradient descent algorithm with a loss or an uncertainty-based acquisition function. Our work provides new results on conditions that ensure the convergence of loss- and uncertainty-based active learning algorithms, along with bounds on convergence rates, accommodating various loss functions.
In the case of loss-based acquisition strategies, we assume that the active learner has access to an oracle providing an unbiased estimate of the conditional expected loss for a point, given the feature vector of the point and the current prediction model. In practical settings, the loss cannot be evaluated at acquisition time since labels are yet unknown. Instead, a separate prediction model can be devised, which can estimate the loss subject to estimation noise. In our experiments, we assess the impact of the noise in such a loss estimator for various loss-based active learning strategies.
It is noteworthy that uncertainty and loss-based acquisition strategies are also of interest for the data subset selection problem, often referred to as core-set selection or data pruning. This problem involves finding a small subset of training data such that the predictive performance of a classifier trained on it is close to that of a classifier trained on the full training data. Recent studies have explored this problem in the context of training neural networks, as seen in works like Toneva et al. (2019); Coleman et al. (2020); Paul et al. (2021); Sorscher et al. (2022); Mindermann et al. (2022). In such scenarios, the oracle can evaluate an underlying loss function exactly, avoiding the need for a noisy estimator.
Our contributions can be summarized as follows:
$\bullet$ We outline a set of conditions under which a non-asymptotic convergence rate of order $O(1/n)$ holds, where $n$ represents the number of iterations of the algorithm, i.e., the number of unlabeled points presented to the algorithm. These conditions enable us to establish convergence rate results for loss-based sampling in linearly separable datasets, with the loss function taking on various forms such as the squared hinge loss function, generalized hinge loss function, or satisfying other specified conditions. Our results provide bounds for both expected loss and the number of sampled points, encompassing different loss-based strategies.
$\bullet$ The results discussed above for loss-based strategies, concerning linear classifiers and linearly separable datasets, rely on a key lemma that may be of independent interest. The conditions of the lemma can be verified for various loss and uncertainty-based acquisition strategies. The convergence rate established in Raj and Bach (2022) for an uncertainty-based strategy follows from this lemma, with a specific choice of the loss function used for training and the loss function used for evaluation. The lemma is derived by generalizing an argument used in the proof of Theorem 3.1 in Raj and Bach (2022) to accommodate different acquisition strategies and loss functions.
$\bullet$ We introduce a framework for establishing convergence rate bounds when sampling points according to an arbitrary, increasing function of the conditional expected loss of a point. This framework is particularly noteworthy as it extends the applicability to a broader set of loss functions compared to our initial conditions. Leveraging the observation that the algorithm behaves akin to a stochastic gradient descent algorithm with an underlying objective function, as referred to as an ”equivalent loss” in Liu and Li (2023), allows us to apply known convergence rate results for stochastic gradient descent algorithms. We present a condition under which explicit bounds on the convergence rate can be established in terms of the sampling probability function, and we provide conditions for this to hold for smooth convex loss functions.
$\bullet$ We propose a learning algorithm that combines a sampling-based acquisition strategy with an adaptive step-size stochastic gradient descent update, achieving the stochastic Polyak’s step size update Loizou et al. (2021) in expectation. We establish a condition on the sampling strategy under which a non-asymptotic convergence rate of order $O(1/n)$ holds for smooth convex loss functions. We present uncertainty and loss-based strategies that satisfy this condition for binary classification, as well as an uncertainty strategy for multi-class classification.
$\bullet$ We present numerical results that demonstrate the efficiency of sampling with stochastic Polyak’s step size and its robustness to loss estimation noise.
1.1 Related Work
The early proposal of the query-by-committee (QBC) algorithm by (Seung et al., 1992) demonstrated the benefits of active learning, an analysis of which was conducted under the selective sampling model by Freund et al. (1997) and Gilad-bachrach et al. (2005). Dasgupta et al. (2009) showed that the performance of QBC can be efficiently achieved by a modified perceptron algorithm with adaptive filtering. The efficient and label-optimal learning of halfspaces was studied by Yan and Zhang (2017) and, subsequently, by Shen (2021). Online active learning algorithms, studied under the name of selective sampling, include works by (Cesa-Bianchi et al., 2006, 2009; Dekel et al., 2012; Orabona and Cesa-Bianchi, 2011; Cavallanti et al., 2011; Agarwal, 2013). For a survey, refer to Settles (2012).
Uncertainty sampling has been utilized for classification tasks since as early as (Lewis and Gale, 1994), and subsequently in many other works, such as (Schohn and Cohn, 2000; Zhu et al., 2010; Yang et al., 2015; Yang and Loog, 2016; Lughofer and Pratama, 2018). Mussmann and Liang (2018) demonstrated that threshold-based uncertainty sampling on a convex loss can be interpreted as performing a pre-conditioned stochastic gradient step on the population zero-one loss. However, none of these works have provided theoretical convergence guarantees. Different variants of uncertainty sampling include margin of confidence sampling, least confidence sampling, and entropy-based sampling (Nguyen et al., 2022). The convergence of margin of confidence sampling was recently studied by Raj and Bach (2022), who showed linear convergence for the hinge loss function for a family of selection probability functions and an algorithm that performs a stochastic gradient descent update with respect to the squared hinge loss function.
A loss-based active learning algorithm was proposed by Yoo and Kweon (2019), comprising a loss prediction module and a target prediction model. The algorithm uses the loss prediction module to compute a loss estimate and prioritizes sampling points with a high estimated loss under the current prediction model. Lahlou et al. (2022) generalize this idea within a framework for uncertainty prediction. However, neither Yoo and Kweon (2019) nor Lahlou et al. (2022) provided theoretical guarantees for convergence rates.
Recent analysis of convergence for loss and uncertainty-based active learning strategies has been presented by Liu and Li (2023). Specifically, they introduced the concept of an equivalent loss, demonstrating that a gradient descent algorithm employing point sampling can be viewed as a stochastic gradient descent algorithm optimizing an equivalent loss function. While they focused on discussing specific cases like sampling proportional to conditional expected loss for loss-based acquisition strategies, we introduce a framework that accommodates sampling based on any continuous increasing function of expected conditional loss. Our framework provides explicit convergence rate bounds in terms of the underlying sampling probability function.
In addition, Loizou et al. (2021) introduced a stochastic gradient descent algorithm featuring an adaptive stochastic Polyak’s step size, which has theoretical convergence guarantees under various assumptions. This algorithm showcased robust performance in comparison to state-of-the-art optimization methods, especially when training over-parametrized models. Our work proposes a novel sampling method that employs stochastic Polyak’s step size in expectation, offering a convergence rate guarantee for smooth convex loss functions, contingent on a condition related to the sampling probability function. Notably, we demonstrate the fulfillment of this condition for logistic regression and binary cross-entropy loss functions, encompassing both a loss-based strategy involving proportional sampling to absolute error loss and an uncertainty sampling strategy. Furthermore, we extend this condition to hold for an uncertainty sampling strategy designed for multi-class classification.
2 Background
Algorithm
We consider the projected stochastic gradient descent algorithm defined as: given an initial value $\theta_{1}$,
$\theta_{t+1}=\mathcal{P}_{\Theta_{0}}\left(\theta_{t}-z_{t}\nabla_{\theta}\ell(x_{t},y_{t},\theta_{t})\right),\hbox{ for }t\geq 1$ (1)
where $z_{t}$ is a stochastic step size with mean $\zeta(x_{t},y_{t},\theta_{t})$ for some function $\zeta:\mathcal{X}\times\mathcal{Y}\times\Theta\mapsto\mathbb{R}_{+}$, $\Theta_{0}\subseteq\Theta$, and $\mathcal{P}_{\Theta}$ is the projection function, i.e., $\mathcal{P}_{\Theta_{0}}(u)=\arg\min_{v\in\Theta\cap\Theta_{0}}||u-v||$. Unless specified otherwise, we consider the case $\Theta_{0}=\Theta$, which requires no projection.
For the choice of the stochastic step size, we consider two cases: (a) Bernoulli sampling with a constant step size, and (b) stochastic Polyak’s step size. For case (a), $z_{t}$ is the product of a constant step size $\gamma$ and a Bernoulli random variable with mean $\pi(x_{t},y_{t},\theta_{t})$. For case (b), $\zeta(x,y,\theta)$ is the ”stochastic” Polyak’s step size, and $z_{t}$ is equal to $\zeta(x_{t},y_{t},\theta_{t})/\pi(x_{t},y_{t},\theta_{t})$ with probability $\pi(x_{t},y_{t},\theta_{t})$ and is equal to 00 otherwise. Note that using the notation $\pi(x,y,\theta)$ allows for the case when the sampling probability does not depend on the value of the label $y$.
Linearly Separable Data
For binary classification tasks, $\mathcal{Y}=\{-1,1\}$. We say that data is separable if, for every point $(x,y)\in\mathcal{X}\times\mathcal{Y}$, either $y=1$ with probability $1$ or $y=-1$ with probability $1$. The data is linearly separable if there exists $\theta^{*}\in\Theta$ such that $y=\mathrm{sgn}(x^{\top}\theta^{*})$ for every $x\in\mathcal{X}$. Linearly separable data has a $\rho^{*}$-margin if $|x^{\top}\theta^{*}|\geq\rho^{*}$ for every $x\in\mathcal{X}$, for some $\theta^{*}\in\Theta$.
Linear Classifiers
Some of our results are for linear classifiers, where the predicted label of a point $x$ is a function of $x^{\top}\theta$. For example, a model with a predicted label $\mathrm{sgn}(x^{\top}\theta)$ is a linear classifier. For logistic regression, the predicted label is $1$ with probability $\sigma(x^{\top}\theta)$ and $-1$ otherwise, where $\sigma$ is the logistic function defined as $\sigma(z)=1/(1+e^{-z})$.
Smooth Loss Functions
For any given $(x,y)\in\mathcal{X}\times\mathcal{Y}$, the loss function $\ell(x,y,\theta)$ is considered smooth on $\Theta^{\prime}\subseteq\Theta$ if it has a Lipschitz continuous gradient on $\Theta^{\prime}$, i.e., there exists $L_{x,y}$ such that $||\nabla_{\theta}\ell(x,y,\theta_{1})-\nabla_{\theta}\ell(x,y,\theta_{2})||\leq L_{x,y}||\theta_{1}-\theta_{2}||$ for all $\theta_{1},\theta_{2}\in\Theta^{\prime}$. For any distribution $q$ over $\mathcal{X}\times\mathcal{Y}$, $\mathbb{E}_{(x,y)\sim q}[\ell(x,y,\theta)]$ is $\mathbb{E}_{(x,y)\sim q}[L_{x,y}]$-smooth.
3 Convergence Rate Guarantees
In this section, we present conditions on the stochastic step size of algorithm (1) under which we can bound the total expected loss and, in some cases, bound the expected number of samples. For Bernoulli sampling with a constant step size, we first provide conditions that allow us to derive bounds for linear classifiers and linearly separable data (Section 3.1). Next, we provide bounds for loss-based sampling, leveraging convergence results known to hold for stochastic gradient descent algorithms (Section 3.2). For sampling combined with stochastic Polyak’s step size, we offer a condition that allows us to establish convergence bounds for loss and uncertainty-based sampling (Section 3.1).
3.1 Bernoulli Sampling with Constant Step Size: First Set of Convergence Conditions
The convergence rate results presented in this section are established using Lemma A.2, which, for space reasons, is provided in Appendix A.1. These results pertain to linear classifiers and linearly separable datasets. We initially address binary classification and subsequently extend our considerations to multi-class classification. The conditions for convergence involve both the loss function $\ell$ utilized by algorithm (1) and a separate loss function $\tilde{\ell}$ employed for evaluating the algorithm’s performance.
Binary Classification
We consider binary classification based on the sign of $x^{\top}\theta$. With a slight abuse of notation, let $\ell(x,y,\theta)\equiv\ell(u)$, ~$\tilde{\ell}(x,y,\theta)\equiv\tilde{\ell}(u)$, and $\pi(x,y,\theta)\equiv\pi(u)$ where $u=yx^{\top}\theta$. We assume that the domain $\mathcal{X}$ is bounded, i.e., there exists $R$ such that $||x||\leq R$ for all $x\in\mathcal{X}$, $||\theta_{1}-\theta^{*}||\leq S$, and that the data is $\rho^{*}$-margin linearly separable (see Section 2 for the definition).
Under prevailing assumptions, the following conditions imply conditions of Lemma A.2: for all $u\in\mathbb{R}$,
$\pi(u)\ell^{\prime}(u)^{2}R^{2}\leq\alpha\tilde{\ell}(u)$ (2)
and
$\pi(u)(-\ell^{\prime}(u))(\rho^{*}-u)\geq\beta\tilde{\ell}(u).$ (3)
We demonstrate convergence results for a class of loss functions and sampling proportional to either zero-one or absolute error loss, and then for the squared hinge loss function and a smoothed hinge function (which accommodates the hinge loss function as a limit case).
For any $x$, $y$, and $\theta$, the zero-one loss and absolute error loss are defined as $1_{\{y\neq\mathrm{sgn}(x^{\top}\theta)\}}=\ell_{01}(u):=1_{\{u\leq 0\}}$ and $|y-\mathrm{sgn}(x^{\top}\theta)|=\ell_{\mathrm{abs}}(u):=2\left(1_{\{u<0\}}+\frac{1}{2}1_{\{u=0\}}\right)$, respectively, where $u=yx^{\top}\theta$.
We first consider sampling proportional to either zero-one loss or absolute error loss when the loss function satisfies the following conditions:
Assumption 3.1.
Function $\ell$ is continuously differentiable on $(-\infty,0]$, convex, and $\ell^{\prime}(0)\leq-c_{1}$ and $\lim_{u\rightarrow-\infty}\ell^{\prime}(u)\geq-c_{2}$, for some constants $c_{1},c_{2}>0$.
In particular, conditions (2) and (3) hold for the case when the loss function $\ell$ is the hinge loss function. For this case, the sampling probability equal to zero-one loss makes algorithm (1) correspond to the well-known perceptron algorithm. This establishes a precise relation between loss-based sampling and the perceptron algorithm.
Theorem 3.2.
Assume that the loss function $\ell$ satisfies Assumption 3.1 and $\rho^{*}>0$. Then, under sampling proportional to zero-one loss with a constant factor $\omega\in(0,1]$, for any initial value $\theta_{1}$ such that $||\theta_{1}-\theta^{*}||\leq S$ and $\{\theta_{t}\}_{t>1}$ according to algorithm (1) with $\gamma=c_{1}\rho^{*}/(c_{2}^{2}R^{2})$,
$\mathbb{E}\left[\sum_{t=1}^{n}\ell_{01}(y_{t}x_{t}^{\top}\theta_{t})\right]\leq\frac{c_{2}^{2}R^{2}S^{2}}{c_{1}^{2}\omega}\frac{1}{{\rho^{*}}^{2}}.$ (4)
Furthermore, under sampling proportional to absolute error loss with a factor $\omega\in(0,1/2]$ and $\gamma=2c_{1}\rho^{*}/(c_{2}^{2}R^{2})$, the bound in (4) holds but with an additional factor of $2$.
The bound in (4) is a well-known bound on the number of mistakes made by the perceptron algorithm, of the order $O(1/{\rho^{*}}^{2})$. It can be readily observed that under sampling proportional to zero-one loss, the expected number of sampled points is bounded by $(c_{2}/c_{1})^{2}R^{2}S^{2}/{\rho^{*}}^{2}$, which also holds for sampling proportional to absolute error loss but with an additional factor of $4$.
We next consider sampling for the squared hinge loss function $\ell(u)=(1/2)\max\{1-u,0\}^{2}$.
Theorem 3.3.
Assume that $\rho^{*}>1$, the loss function is the squared hinge loss function, and the sampling probability function $\pi$ is such that for all $u\leq 1$, $\pi(u)\leq\beta/2$ and
$\pi(u)\geq\pi^{*}(\ell(u)):=\frac{\beta}{2}\left(1-\frac{1}{1+\mu\sqrt{\ell(u)}}\right)$ (5)
for some constants $0<\beta\leq 2$ and $\mu\geq\sqrt{2}/(\rho^{*}-1)$. Then, for any initial value $\theta_{1}$ such that $||\theta_{1}-\theta^{*}||\leq S$ and $\{\theta_{t}\}_{t>1}$ according to algorithm (1) with $\gamma=1/R^{2}$,
$\mathbb{E}\left[\ell(yx^{\top}\bar{\theta}_{n})\right]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(y_{t}x_{t}^{\top}\theta_{t})\right]\leq\frac{R^{2}S^{2}}{\beta}\frac{1}{n}.$
Moreover, if the sampling is according to $\pi^{*}$, then the expected number of sampled points satisfies
$\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(\ell(y_{t}x_{t}^{\top}\theta_{t}))\right]\\ \leq\min\left\{\frac{1}{2}RS\mu\sqrt{\beta}\sqrt{n},\frac{1}{2}\beta n\right\}.$
Condition (5) requires that the sampling probability function $\pi$ is lower bounded by an increasing, concave function $\pi^{*}$ of the loss value. This fact, along with the expected loss bound, implies the asserted bound for the expected number of samples. The expected number of samples is $O(\sqrt{n})$ concerning the number of iterations and is $O(1/(\rho^{*}-1))$ concerning the margin $\rho^{*}-1$.
Next, we consider the generalized smooth hinge loss function (Rennie, 2005), defined for $a\geq 1$ as follows:
$\ell(u)=\left\{\begin{array}[]{ll}\frac{a}{a+1}-u&\text{ if }u\leq 0\\ \frac{a}{a+1}-u+\frac{1}{a+1}u^{a+1}&\text{ if }0\leq u\leq 1\\ 0&\text{ otherwise}.\end{array}\right.$ (6)
This is a continuously differentiable function that converges to the value of the hinge loss function, $\max\{1-u,0\}$, as $a$ goes to infinity. The family of loss functions parameterized by $a$ accommodates the smooth hinge loss function with $a=1$, introduced by Rennie and Srebro (2005).
Theorem 3.4.
Assume that $\rho^{*}>1$, the loss function is the generalized smooth hinge loss function, and the sampling probability is according to the function $\pi^{*}$, which, for $\beta\in(0,1]$ and $\rho\in(1,\rho^{*}]$, is defined as
$\pi^{*}(u)=\left\{\begin{array}[]{ll}\beta\frac{\frac{a}{a+1}-u}{\rho-u}&\text{ if }u\leq 0\\ \beta\frac{\frac{a}{a+1}\frac{1}{1-u^{a}}(1-u)-\frac{1}{a+1}u}{\rho-u}&\text{ if }0\leq u\leq 1\\ 0&\text{ if }u\geq 1.\end{array}\right.$
Then, for any initial value $\theta_{1}$ such that $||\theta_{1}-\theta^{*}||\leq S$ and $\{\theta_{t}\}_{t>1}$ according to algorithm (1) with $\gamma=1/(c_{a,\rho}R^{2})$, where $c_{a,\rho}=a/(a(\rho-1)+1)$, we have
$\mathbb{E}\left[\ell(yx^{\top}\bar{\theta}_{n})\right]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(y_{t}x_{t}^{\top}\theta_{t})\right]\leq c_{a,\rho}\frac{R^{2}S^{2}}{\beta}\frac{1}{n}.$
Furthermore, the same bound holds for every $\pi$ such that $\pi^{*}(u)\leq\pi(u)\leq\beta$ for all $u\in\mathbb{R}$, with $c_{a,\rho}=2a$.
Note that the sampling probability function $\pi^{*}$ is increasing in $a$ and is upper-bounded by $\pi^{**}(u)=\beta(1-u)/(\rho-1)$, which may be regarded as the limit for the hinge loss function. See Figure 1 for an illustration.
We remark that the expected loss bound in Theorem 3.4 is the same as for the squared hinge loss function in Theorem 3.3, except for an additional factor $c_{a,\rho}=a/(a(\rho-1)+1)$. This factor is increasing in $a$ but always lies in the interval $[1/\rho,1/(\rho-1)]$, where the boundary values are achieved for $a=1$ and $a\rightarrow\infty$, respectively.
Figure 1: Sampling probability function for the family of generalized smooth hinge loss functions.
Theorem 3.5.
Assuming $\rho^{*}>1$, the loss function is the generalized smooth hinge loss, and the sampling probability function $\pi^{*}$ satisfies the assumptions in Theorem 3.4, with $n>((a+1)/a)c_{a,\rho}R^{2}S^{2}/\beta$, the expected number of sampled points is bounded as
$\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(y_{t}x_{t}^{\top}\theta_{t})\right]\leq\kappa\max\left\{\frac{\sqrt{\beta c_{a,\rho}}RS}{(\rho-1)\sqrt{a}}\sqrt{n},\frac{c_{a,\rho}R^{2}S^{2}}{\rho-1}\right\},$
where $\kappa$ is some positive constant.
For any fixed number of iterations $n$ satisfying the condition of the theorem, the expected number of sampled points is bounded by a constant for sufficiently large value of parameter $a$. The bound in Theorem 3.5 depends on how the loss function $\ell(u)$ varies with $u$. When $a$ is large, $\ell(u)$ is approximately $1-u$ (hinge loss), otherwise, it is approximately $\frac{a}{2}(1-u)^{2}$ for $0\leq u\leq 1$ (squared hinge loss).
Multi-class Classification
The conditions for multi-class classification with the set of classes $\mathcal{Y}$ are the same as for binary classification, with $u(x,y,\theta):=x^{\top}\theta_{y}-\max_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}x^{\top}\theta_{y^{\prime}}$, except for an additional factor of $2$ in the left-hand side of the inequality in the first condition (see Lemma A.5 in the Appendix). Hence, all the observations remain valid for the multi-class classification case.
Discussion
The conditions in Assumption 15 allow us to obtain convergence rate bounds for linearly separable data for some other predictors and loss functions. However, for linear classifiers, loss functions must satisfy certain properties related to hinge loss function. Further details are discussed in Appendix A.7. We next present a framework for deriving convergence rate bounds in more general cases.
3.2 Bernoulli Sampling with Constant Step Size: Second Set of Convergence Conditions
We consider algorithm (1) where $z_{t}$ is product of a fixed step size $\gamma$ and a Bernoulli random variable $\zeta_{t}$ with mean $\pi(x,y,\theta)$. Let $g_{t}=\zeta_{t}\nabla_{\theta}\ell(x_{t},y_{t},\theta_{t})$, which is random vector because $\zeta_{t}$ is a random variable and $(x_{t},y_{t})$ is a sampled point. Following Liu and Li (2023), we note that the algorithm (1) is a stochastic gradient descent algorithm with respect to an objective function $\tilde{\ell}$ with gradient
$\nabla_{\theta}\tilde{\ell}(\theta)=\mathbb{E}[\pi(x,y,\theta)\nabla_{\theta}\ell(x,y,\theta)]$ (7)
where the expectation is with respect to $x$ and $y$. This observation allows us to derive convergence rate results by deploying convergence rate results for stochastic gradient algorithm known to hold under various assumptions on function $\tilde{\ell}$, variance of stochastic gradient vector and step size. A function $\tilde{\ell}$ satisfying condition (7) is referred to as equivalent loss in Liu and Li (2023).
Assume that the sampling probability $\pi$ is an increasing function of the conditional expected loss $\ell(x,\theta)=\mathbb{E}_{y}[\ell(x,y,\theta)\mid x]$. With a slight abuse of notation, we denote this probability as $\pi(\ell(x,\theta))$ where $\pi:\mathbb{R}_{+}\rightarrow[0,1]$ is an increasing and continuous function. Let $\Pi$ be the primitive of $\pi$, i.e. $\Pi^{\prime}=\pi$. We then have
~$\tilde{\ell}(\theta)=\mathbb{E}[\Pi(\ell(x,\theta))].$ (8)
The function ~$\tilde{\ell}(\theta)$ inherits some properties of functions $\ell(x,y,\theta)$, for $(x,y)\in\mathcal{X}\times\mathcal{Y}$. For example, if $\ell(x,y,\theta)$ is a convex function, for every $(x,y)\in\mathcal{X}\times\mathcal{Y}$, then $\tilde{\ell}$ is a convex function.
The framework for establishing convergence rates allows us to accommodate different sampling strategies and loss functions. For example, it accommodates binary classification for binary cross-entropy loss function and sampling proportional to absolute error loss, as shown in the following.
For binary classification, we model the prediction probability of the positive label as $\sigma(x^{\top}\theta)$, where $\sigma:\mathbb{R}_{+}\rightarrow[0,1]$ is an increasing function, and $\sigma(-u)+\sigma(u)=1$ for all $u\in\mathbb{R}$. The absolute error loss takes value $1-\sigma(x^{\top}\theta)$ if $y=1$ or value $\sigma(x^{\top}\theta)$ if $y=-1$, which corresponds to $1-\sigma(yx^{\top}\theta)$. The binary cross-entropy loss for a point $(x,y)$ under model parameter $\theta$ can be written as $\ell(x,y,\theta)=-\log(\sigma(yx^{\top}\theta))$. Hence, absolute error loss-based sampling corresponds to the sampling probability function $\pi(x)=1-e^{-x}$.
The next lemma allows to derive convergence rate results for expected loss with respect to loss function $\ell$ by applying convergence rate results for expected loss with respect to loss function $\tilde{\ell}$ (which, recall, is the equivalent loss function).
Lemma 3.6.
Assume that for algorithm (1) with loss-based sampling according to $\pi$, for some functions $f_{1},\ldots,f_{m}$, we have
$\mathbf{E}\left[\frac{1}{n}\sum_{t=1}^{n}\tilde{\ell}(\theta_{t})\right]\leq\inf_{\theta}\tilde{\ell}(\theta)+\sum_{i=1}^{m}f_{i}(n).$ (9)
Then, it holds:
$\mathbf{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]$$\leq$ $\inf_{\theta}\Pi^{-1}(\tilde{\ell}(\theta))+\sum_{i=1}^{m}\Pi^{-1}(f_{i}(n)).$
Table 1: Examples of sampling probability functions.
$\pi(x)$ $\Pi(x)$ $\Pi^{-1}(x)$
$1-e^{-x}$ $x+e^{-x}-1$ $\approx\sqrt{2x}$ for small $x$
$\min\{x,1\}$$\left\{\begin{array}[]{ll}\frac{1}{2}x^{2}&x\leq 1\\ x-\frac{1}{2}&x\geq 1\end{array}\right.$$\left\{\begin{array}[]{ll}\sqrt{2x}&x\leq 1/2\\ x+\frac{1}{2}&x\geq 1/2\end{array}\right.$
$\min\{(x/b)^{a},1\}$, $a>0,b>0$$\left\{\begin{array}[]{ll}\frac{1}{b^{a}(1+a)}x^{1+a}&x\leq a\\ x-\frac{a}{1+a}&x\geq a\end{array}\right.$$\left\{\begin{array}[]{ll}b^{\frac{a}{1+a}}(1+a)^{\frac{1}{1+a}}x^{\frac{1}{1+a}}&x\leq\frac{b}{1+a}\\ x+\frac{a}{1+a}b&x\geq\frac{b}{1+a}\end{array}\right.$
$1-\frac{1}{1+\mu x}$ $x-\frac{1}{\mu}\log(1+\mu x)$ $\approx\sqrt{(2/\mu)x}$ for small $x$
$1-\frac{1}{1+\mu\sqrt{x}}$ $x-\frac{2}{\mu}\sqrt{x}+\frac{2}{\mu^{2}}\log(1+\mu\sqrt{x})$ $\approx(((3/2)/\mu)x)^{2/3}$ for small $x$
We apply Lemma 3.6 to obtain the following result.
Theorem 3.7.
Assume that $\ell$ is a convex function, $\tilde{\ell}$ is $L$-smooth, $\Theta_{0}$ is a convex set, $S=\sup_{\theta\in\Theta_{0}\cap\Theta}||\theta-\theta_{1}||$, and $\mathbb{E}[\pi(\ell(x,\theta))||\nabla_{\theta}\ell(x,y,\theta)||^{2}]-||\nabla_{\theta}\tilde{\ell}(\theta)||^{2}\leq\sigma_{\pi}^{2}$. Then, algorithm (1) with $\gamma=1/(L+(\sigma/R)\sqrt{n/2})$,
$\mathbb{E}[\ell(\bar{\theta}_{n})]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]$
$\leq$ $\inf_{\theta}\Pi^{-1}(\tilde{\ell}(\theta))+\Pi^{-1}\left(\frac{\sqrt{2}S\sigma_{\pi}}{\sqrt{n}}\right)+\Pi^{-1}\left(\frac{LS^{2}}{n}\right).$
Note that the bound on the expected loss in Theorem 3.7 depends on $\pi$ through $\Pi^{-1}$ and $\sigma_{\pi}^{2}$. Obviously, we have a bound depending on $\pi$ only through $\Pi^{-1}$ by upper bounding $\sigma_{\pi}^{2}$ with $\sup_{\theta\in\Theta_{0}}\mathbb{E}[||\nabla_{\theta}\ell(x,y,\theta)||^{2}]$.
For convergence rates for large values of the number of iterations $n$, the bound in Theorem 3.6 crucially depends on how $\Pi^{-1}(x)$ behaves for small values of $x$. In Table 1, we show $\Pi$ and $\Pi^{-1}$ for several examples of sampling probability function $\pi$. For all examples in the table, $\Pi^{-1}(x)$ is sub-linear in $x$ for small $x$. For instance, for absolute error loss sampling under binary cross-entropy loss function, $\pi(x)=1-e^{-x}$, $\Pi^{-1}(x)$ is approximately $\sqrt{2x}$ for small $x$. For this case, we have the following corollary.
Corollary 3.8.
Under assumptions of Theorem 3.7, sampling probability $\pi(x)=1-e^{-x}$, and $n\geq\max\left\{\left(\frac{16}{9}\right)^{2}2S\sigma_{\pi}^{2},\frac{16}{9}LS^{2}\right\}$ it holds
$\mathbb{E}[\ell(\bar{\theta}_{n})]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]$
$\leq$ $\inf_{\theta}\Pi^{-1}(\tilde{\ell}(\theta))+2^{5/4}\sqrt{S\sigma_{\pi}}\frac{1}{\sqrt[4]{n}}+2\sqrt{L}S\frac{1}{\sqrt{n}}.$
By using a bound on the expected total loss, we can bound the expected total number of sampled points under certain conditions as follows.
Lemma 3.9.
The following bounds hold:
  1. 1.
    Assume that $\pi$ is a concave function, then $\mathbb{E}\left[\sum_{t=1}^{n}\pi(\ell(x_{t},\theta_{t}))\right]\leq\pi\left(\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]\right)n.$
  2. 2.
    Assume that $\pi$ is $K$-Lipschitz or that for some $K>0$, $\pi(\ell)\leq\min\{K\ell,1\}$ for all $\ell\geq 0$, then $\mathbb{E}\left[\sum_{t=1}^{n}\pi(\ell(x_{t},\theta_{t}))\right]\leq\min\left\{K\mathbb{E}\left[\sum_{t=1}^{n}\ell(\theta_{t})\right],n\right\}.$
We remark that $\pi$ is a concave function for all examples in Table 1 without any additional conditions, except for $\pi(\ell)=\min\{(\ell/b)^{a},1\}$ which is concave under assumption $0<a\leq 1$. We remark also that for every example in Table 1 except the last one, $\pi(\ell)\leq\min\{K\ell,1\}$ for some $K>0$. Hence, for all examples in Table 1, we have a bound for the expected number of sampled points provided we have a bound for the expected loss.
3.3 Bernoulli Sampling with Stochastic Polyak’s Step Size: A New Algorithm
In this section we propose a new algorithm that combines a Bernoulli sampling and an adaptive stochastic gradient descent update, and provide a convergence rate guarantee. The algorithm is defined as the stochastic gradient descent algorithm (1) with stochastic step size $z_{t}$ being a binary random variable that takes value $\gamma_{t}:=\zeta(x_{t},y_{t},\theta_{t})/\pi(x_{t},y_{t},\theta_{t})$ with probability $\pi(x_{t},y_{t},\theta_{t})$ and takes value 00 otherwise, where $\pi$ is some function and if $||\nabla_{\theta}\ell(x,y,\theta)||>0$,
$\zeta(x,y,\theta)=\beta\min\left\{\frac{1}{\psi(x,y,\theta)},\rho\right\}$
and, $\zeta(x,y,\theta)=0$, otherwise, for constants $\beta,\rho>0$ and
$\psi(x,y,\theta):=\frac{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}{\ell(x,y,\theta)-\inf_{\theta^{\prime}}\ell(x,y,\theta^{\prime})}.$
The expected step size $\zeta(x,y,\theta)$ corresponds to the adaptive step size proposed by Polyak (Loizou et al., 2021). When $\pi(x,y,\theta)=1$ for all $x,y,\theta$, the algorithm reduces to Polyak’s adaptive step size algorithm (Loizou et al., 2021). Our algorithm introduces a sampling component and re-weighting of the update to ensure that the step size remains according to the stochastic Polyak’s step size in expectation. For many loss functions $\inf_{\theta^{\prime}}\ell(x,y,\theta)=0$, for every $x,y$. In these cases, $\psi(x,y,\theta)=||\nabla_{\theta}\ell(x,y,\theta)||^{2}/\ell(x,y,\theta)$. For instance, for binary cross-entropy loss function, $\inf_{\theta^{\prime}}\ell(x,y,\theta)=\inf_{\theta^{\prime}}(-\log(\sigma(yx^{\top}\theta^{\prime})))=0$, for all $x,y$.
We next show a convergence rate guarantee.
Lemma 3.10.
Assume that $\ell$ is a convex, $L$-smooth function, there exists $\Lambda^{*}$ such that $\mathbb{E}[\ell(x,y,\theta^{*})]-\mathbb{E}[\inf_{\theta}\ell(x,y,\theta)]\leq\Lambda^{*}$, and the sampling probability function $\pi$ is such that, for some constant $c\in(0,1)$, for all $x,y,\theta$ such that $||\nabla_{\theta}\ell(x,y,\theta)||>0$,
$\pi(x,y,\theta)\geq\frac{\beta}{2(1-c)}\min\left\{\rho\psi(x,y,\theta),1\right\}.$ (10)
Then, we have
$\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}(\ell(x_{t},y_{t},\theta_{t})-\ell(x_{t},y_{t},\theta^{*}_{t}))\right]$
$\leq$ $\frac{\rho\beta}{c\kappa}\Lambda^{*}+\frac{1}{2c\kappa}||\theta_{1}-\theta^{*}||^{2}\frac{1}{n}$
where $\kappa=\beta\min\{1/(2L),\rho\}$ and $\theta_{t}^{*}$ is a minimizer of $\ell(x_{t},y_{t},\theta^{\prime})$ over $\theta^{\prime}$.
By taking $\rho=1/(2L)$, the bound on the expected average loss boils down to $\Lambda^{*}/c+(L/(c\beta))||\theta_{1}-\theta^{*}||^{2}/n$.
Notably, under the condition on the sampling probability in Lemma 3.10, the convergence rate is of order $O(1/n)$. A similar bound is known to hold for stochastic gradient descent with adaptive stochastic Polyak step size for the finite-sum problem, as seen in Theorem 3.4 of Loizou et al. (2021). Lemma 3.10 applies to stream-based stochastic gradient descent algorithms and, importantly, allows for Bernoulli sampling with a sampling probability function $\pi$.
We next show instances of $\pi$ that satisfy condition (10).
Constant Sampling Probability
Assume that $\kappa\leq 2(1-c)$ and sampling probability function $\pi$ is constant with value at least $\frac{\kappa}{2(1-c)}$. Then, the expected loss bound in Lemma 3.10 holds.
Power Function of the Expected Step Size
Consider $\pi$ defined as
$\pi(x,y,\theta)=\zeta(x,y,\theta)^{\eta}$ (11)
where $\eta\geq 0$ and $\rho\beta\in(0,1]$. This sampling probability function is an increasing function of $1/\psi(x,y,\theta)$, i.e. $(\ell(x,y,\theta)-\inf_{\theta^{\prime}}\ell(x,y,\theta))/||\nabla_{\theta}\ell(x,y,\theta)||^{2}$. In other words, the sampling probability is an increasing function of the loss normalized by the squared norm of the gradient. Assume that
$\beta^{1-\eta}\leq 2(1-c)\left(\frac{1}{2L}\right)^{\eta}.$ (12)
Then, the loss bound in Theorem 3.10 holds.
Linear Classifiers: Binary Classification
For linear classifiers, if $\ell(u)$ is an $L^{\prime}$-smooth function in $u$, then $\ell(yx^{\top}\theta)$ is an $L^{\prime}||x||^{2}$-smooth function in $\theta$. To satisfy condition (10), it suffices that
$\pi(x,y,\theta)\geq\frac{\beta}{2(1-c)}\min\left\{\rho R^{2}h(yx^{\top}\theta),1\right\}$
where $h(u):=\ell^{\prime}(u)^{2}/\ell(u)$ and note that $||\nabla_{\theta}\ell(x,y,\theta)||^{2}/\ell(x,y,\theta)=h(yx^{\top}\theta)||x||^{2}$. By setting $\rho=1/(2L)$ with $L=L^{\prime}R^{2}$, we get $\rho R^{2}=1/(2L^{\prime})$.
For logistic regression and the binary cross-entropy loss function, the function $h(u)$ can be written as follows
$h(u)=\frac{1}{(1+e^{u})^{2}\log(1+e^{-u})}$ (13)
which is increasing in $u$ for $u\leq 0$ and decreasing in $u$ otherwise. Note that $h(u)=(1-e^{-\ell(u)})^{2}/\ell(u)$. Further properties of function $h$, deferred to Appendix A.13, establish two key facts: (a) Condition of Lemma 3.10 is satisfied with $\rho=1/(2L)$ by sampling proportional to absolute error loss $\pi^{*}(u)=\omega(1-\sigma(u))$ with $\beta/(4(1-c)L^{\prime})\leq\omega\leq 1$, and (b) Condition of Lemma 3.10 is satisfied by uncertainty sampling according to, for any fixed $a\in(0,1/2]$,
$\pi^{*}(u)=\frac{\beta}{2(1-c)}\min\left\{\rho R^{2}\frac{1}{H(a)+(1-a)|u|},1\right\}$
where $H(a)=a\log(1/a)+(1-a)\log(1/(1-a))$.
Linear Classifiers: Multi-class Classification
Condition (10) holds for some loss and uncertainty-based sampling methods. Refer to Appendix A.15 for detailed discussion.
4 Numerical Results
Figure 2: Average cross-entropy progressive loss of Polyak’s step size Loizou et al. (2021) compared to SGD with constant step size, for 1% and 10% sampling from the mushrooms data.
Figure 3: Convergence in terms of average cross-entropy progressive loss of random sampling, loss-based sampling based on the absolute error loss, and our proposed adaptive step size variant of absolute error loss sampling.
We adopt the setting from Loizou et al. (2021) and utilize the mushroom classification dataset (Chang and Lin, 2011) for binary classification tasks using RBF kernels and logistic regression. This dataset comprises $n=6499$ instances and $d=6499$ dimensions. In our training approach, we deliberately confine the training process to a single epoch. Throughout this epoch, we sequentially process each data instance, compute the loss for each individual instance, and subsequently update the model’s weights. This approach, known as progressive validation (Blum et al., 1999), enables us to monitor the evolution of the average loss. The constraint to a single epoch ensures that we calculate losses only for instances that haven’t influenced model weights. For each sampling scheme, we conduct a hyper-parameter sweep to minimize the average progressive loss.
Loizou et al. (2021) demonstrated that stochastic gradient descent with a step size corresponding to their stochastic Polyak’s step size (SPS) converges faster than gradient descent. Figure 2 illustrates that these findings extend to scenarios where we selectively sample from the dataset rather than training on the full dataset, and the step size is according to stochastic Polyak’s step size only in expectation.  [^1] We will make the code and the dataset available for download in the final version of the manuscript.
Figure 3 demonstrates that our proposed active learning algorithm, with adaptive Polyak step size according to SPS and sampling according to $\pi(x)=1-e^{-x}$, leads to faster convergence than the traditional loss-based sampling. The figure also shows that the traditional loss-based sampling approach converges more rapidly than random sampling.
Figure 4: Robustness of the proposed sampling approach with adaptive Polyak’s step size for different variance $\mathrm{var}[\hat{\ell}_{\mathrm{abs}}]=\ell^{2}_{\mathrm{abs}}(1-\ell_{\mathrm{abs}})/(\alpha+\ell_{\mathrm{abs}})$ noise levels of absolute error loss estimator: (low) $\alpha=100$, (medium) $\alpha=2.5$, and (high) $\alpha=1$.
We also simulate a noisy estimator of the absolute error loss for the proposed loss-based sampling with adaptive Polyak’s step size. This corresponds to the loss-based active learning setting in practice, where the loss is not known prior to label acquisition and can only be estimated. We model an unbiased noisy estimator ^$\hat{\ell}_{\mathrm{abs}}$ of the absolute error loss $\ell_{\mathrm{abs}}\in[0,1]$ as a random variable following the beta distribution, denoted as ^$\hat{\ell}_{\mathrm{abs}}\sim\mathrm{Beta}(\alpha,\beta)$, where $\alpha$ and $\beta$ are parameters set to ensure $\mathbb{E}[\hat{\ell}_{\mathrm{abs}}]=\ell_{\mathrm{abs}}$. The variance of the noise can be controlled by the tuning parameter $\alpha$, given by $\mathrm{var}[\hat{\ell}_{\mathrm{abs}}]=\frac{\ell_{\mathrm{abs}}(1-\ell_{\mathrm{abs}})}{\alpha+\ell_{\mathrm{abs}}}$. Figure 4). Figure 4 shows that the convergence results remain robust against noise in estimating the absolute error loss for a wide range of values for $\alpha\geq 1$.
5 Conclusion
In this paper, we have provided convergence rate guarantees for both loss and uncertainty-based active learning algorithms under various assumptions. Furthermore, we introduced a novel algorithm that combines sampling with an adaptive size, conforming to stochastic Polyak’s step size in expectation, and demonstrated its convergence rate guarantee, contingent on a condition related to the sampling probability function. For future research, it would be interesting to apply the methodologies presented in this paper to obtain additional convergence rate results. Additionally, exploring a theoretical study on the impact of noise in the loss estimator, used for evaluating the sampling probability function, on the convergence properties of algorithms could open up a valuable avenue for investigation.
6 Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A MISSING PROOFS AND ADDITIONAL RESULTS
A.1 A Lemma used for our First Set of Convergence Conditions
We present a lemma that is used to establish all convergence rate results for the expected loss of algorithm (1) presented in Section 3.1. The lemma relies on the following condition involving the loss function $\ell$ used by algorithm (1) and the loss function $\tilde{\ell}$ used for evaluating the performance of the algorithm.
Assumption A.1.
There exist constants $\alpha,\beta>0$ such that for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and $\theta\in\Theta$,
$\pi(x,y,\theta)||\nabla_{\theta}\ell(x,y,\theta)||^{2}\leq\alpha\tilde{\ell}(x,y,\theta)$ (14)
and
$\pi(x,y,\theta)\nabla_{\theta}\ell(x,y,\theta)^{\top}(\theta-\theta^{*})\geq\beta\tilde{\ell}(x,y,\theta).$ (15)
Lemma A.2.
Under Assumption 15, for any $\theta_{1}$ and $\{\theta_{t}\}_{t>1}$ according to algorithm (1) with $\gamma=\beta/\alpha$,
$\mathbb{E}\left[\sum_{t=1}^{n}\tilde{\ell}(x_{t},y_{t},\theta_{t})\right]\leq||\theta_{1}-\theta^{*}||^{2}\frac{\alpha}{\beta^{2}}.$
Moreover, if for every $(x,y)\in\mathcal{X}\times\mathcal{Y}$, ~$\tilde{\ell}(x,y,\theta)$ is a convex function in $\theta$, then we have
$\mathbb{E}[\tilde{\ell}(x,y,\bar{\theta}_{n})]\leq||\theta_{1}-\theta^{*}||^{2}\frac{\alpha}{\beta^{2}}\frac{1}{n}$
where $\bar{\theta}_{t}=(1/t)\sum_{s=1}^{t}\theta_{s}$.
Proof.
For any $\theta^{*}\in\Theta$ and $t\geq 1$, we have
$||\theta_{t+1}-\theta^{*}||^{2}=||\theta_{t}-\theta^{*}||^{2}-2z_{t}\nabla_{\theta}\ell(x_{t},y_{t},\theta_{t})^{\top}(\theta_{t}-\theta^{*})+z_{t}^{2}||\nabla_{\theta}\ell(x_{t},y_{t},\theta_{t})||^{2}.$
Taking expectation in both sides of the equation, conditional on $x_{t}$, $y_{t}$ and $\theta_{t}$, we have
$\mathbb{E}[||\theta_{t+1}-\theta^{*}||^{2}\mid x_{t},y_{t},\theta_{t}]$ $=$ $||\theta_{t}-\theta^{*}||^{2}-2\gamma\pi(x_{t},y_{t},\theta_{t})\nabla_{\theta}\ell(x_{t},y_{t},\theta_{t})^{\top}(\theta_{t}-\theta^{*})$
$+\gamma^{2}\pi(x_{t},y_{t},\theta_{t})||\nabla_{\theta}\ell(x_{t},y_{t},\theta_{t})||^{2}.$
Under Assumption 15, we have
$\mathbb{E}[||\theta_{t+1}-\theta^{*}||^{2}\mid x_{t},y_{t},\theta_{t}]\leq||\theta_{t}-\theta^{*}||^{2}-2\gamma\beta\tilde{\ell}(x_{t},y_{t},\theta_{t})+\gamma^{2}\alpha\tilde{\ell}(x_{t},y_{t},\theta_{t}).$
Hence, it holds
$(2\gamma\beta-\gamma^{2}\alpha)\mathbb{E}[\tilde{\ell}(x_{t},y_{t},\theta_{t})]\leq\mathbb{E}[||\theta_{t}-\theta^{*}||^{2}]-\mathbb{E}[||\theta_{t+1}-\theta^{*}||^{2}].$
Summing over $t$, we have
$(2\gamma\beta-\gamma^{2}\alpha)\sum_{t=1}^{n}\mathbb{E}[\tilde{\ell}(x_{t},y_{t},\theta_{t})]\leq||\theta_{1}-\theta^{*}||^{2}.$
By taking $\gamma=\beta/\alpha$, we have
$\sum_{t=1}^{n}\mathbb{E}[\tilde{\ell}(x_{t},y_{t},\theta_{t})]\leq||\theta_{1}-\theta^{*}||^{2}\frac{\alpha}{\beta^{2}}.$
The second statement of the lemma follows from the last above inequality and Jensen’s inequality. ∎
A.2 Proof of Theorem 3.2
We first consider the case when sampling is proportional to zero-one loss. We show that conditions of the theorem imply conditions (2) and (3) to hold, for ~$\tilde{\ell}=\ell_{01}$ and $\pi=\omega\ell_{01}$, which in turn imply conditions (14) and (15) and hence we can apply the convergence result of Lemma A.2.
Conditions (2) and (3) are equivalent to:
$\omega\ell^{\prime}(u)^{2}R^{2}\leq\alpha,\hbox{ for every }u\leq 0$
and
$\omega(-\ell^{\prime}(u))(\rho^{*}-u)\geq\beta\hbox{ for every }u\leq 0.$
Since $\ell$ is a convex function $\ell^{\prime}(u)^{2}$ is decreasing in $u$ and $(-\ell^{\prime}(u))(\rho^{*}-u)$ is decreasing in $u$. Therefore, conditions are equivalent to
$\omega(\lim_{u\rightarrow-\infty}(-\ell^{\prime}(u)))^{2}R^{2}\leq\alpha$
and
$\omega(-\ell^{\prime}(0))\rho^{*}\geq\beta.$
These conditions hold true by taking $\alpha=\omega c_{2}^{2}R^{2}$ and $\beta=\omega c_{1}\rho^{*}$.
We next consider the case when sampling is proportional to absolute error loss. We show that conditions of the theorem imply conditions (2) and (3) to hold, for ~$\tilde{\ell}=\ell_{01}$ and $\pi=\omega\ell_{\mathrm{abs}}$, which in turn imply conditions (14) and (15) and hence we can apply the convergence result of Lemma A.2.
Conditions (2) and (3) correspond to
$2\omega(-\ell^{\prime}(u))^{2}R^{2}\leq\alpha\hbox{ for every }u<0\hbox{ and }\omega(-\ell^{\prime}(0))^{2}R^{2}\leq\alpha$
and
$2\omega(-\ell^{\prime}(u))(\rho^{*}-u)\geq\beta\hbox{ for every }u<0\hbox{ and}\omega(-\ell^{\prime}(0))\rho^{*}\geq\beta.$
Again, since $(-\ell^{\prime}(u))^{2}$ is decreasing in $u$ and $(-\ell^{\prime}(u))(\rho^{*}-u)$ is decreasing in $u$, it follows that the conditions are equivalent to
$2\omega(\lim_{u\rightarrow-\infty}(-\ell^{\prime}(u)))^{2}R^{2}\leq\alpha$
and
$\omega(-\ell^{\prime}(0))\rho^{*}\geq\beta.$
Hence, conditions (2) and (3) hold by taking $\alpha=2\omega c_{2}^{2}R^{2}$ and $\beta=\omega c_{1}\rho^{*}$.
A.3 Proof of Theorem 3.3
We show that conditions of the theorem imply conditions (2) and (3) to hold, for $\tilde{\ell}=\ell$, which in turn imply conditions (14) and (15) and hence we can apply the convergence result of Lemma A.2.
We first consider condition (14). For squared hinge loss function $\ell$ and $\tilde{\ell}=\ell$, clearly condition (14) holds for every $u\geq 1$ as in this case both side of the inequality are equal to zero. For every $u\leq 1$,
$\frac{\ell(u)}{\ell^{\prime}(u)^{2}}=\frac{1}{2}.$
Since by assumption $\pi(u)\leq\beta/2$ for all $u$, condition $\alpha/\beta\geq R^{2}$ implies condition (14).
We next consider condition (15). Again, clearly, condition holds for every $u\geq 1$ as in this case both sides of the inequality are equal to zero. For $u\leq 1$, we can write (15) as follows
$\pi(u)\geq\beta\frac{\ell(u)}{(-\ell^{\prime}(u))(\rho^{*}-u)}$ $=$ $\frac{\beta}{2}\frac{1-u}{\rho^{*}-u}$
$=$ $\frac{\beta}{2}\left(1-\frac{1}{1+(1/(\rho^{*}-1))(1-u)}\right)$
$=$ $\frac{\beta}{2}\left(1-\frac{1}{1+(\sqrt{2}/(\rho^{*}-1))\sqrt{\ell(u)}}\right).$
This condition is implied by $\pi(u)\geq\pi^{*}(\ell(u))$ where
$\pi^{*}(\ell)=\frac{\beta}{2}\left(1-\frac{1}{1+\mu\sqrt{\ell}}\right)$
with $\mu\geq\sqrt{2}/(\rho^{*}-1)$.
The result of the theorem follows from Lemma A.2 with $\alpha/\beta=R^{2}$, $0<\beta\leq 2$ and $\pi(u)\geq\pi^{*}(\ell(u))$ for all $u\leq 1$.
For the expected number of samples we proceed as follows. First by concavity and monotonicity of function $\pi^{*}$ and the expected loss bound, we have
$\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(\ell(x_{t},y_{t},\theta_{t}))\right]\leq\pi^{*}\left(\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(x_{t},y_{t},\theta_{t})\right]\right)n\leq\pi^{*}\left(\frac{||\theta_{1}-\theta^{*}||^{2}R^{2}}{\beta n}\right)n.$
Then, combined with the fact $\pi^{*}(v)\leq\frac{\beta\mu}{2}\sqrt{v}$ for all $v\geq 0$, it follows
$\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(\ell(x_{t},y_{t},\theta_{t}))\right]\leq||\theta_{1}-\theta^{*}||\frac{\mu\sqrt{\beta}}{2}\sqrt{n}.$
Since $\pi^{*}(v)\leq\beta/2$ for all $v$, it obviously holds $\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(\ell(x_{t},y_{t},\theta_{t}))\right]\leq(\beta/2)n$, which completes the proof of the theorem.
A.4 Proof of Theorem 3.4
Assume that $\pi^{*}$ is such that for given $\rho\in(1,\rho^{*}]$, $\pi^{*}(u)(-\ell^{\prime}(u))(\rho-u)=\beta\ell(u)$ for all $u\in\mathbb{R}$. Then, $\pi^{*}$ satisfies equation (3) for all $u\in\mathbb{R}$. Note that
$\pi^{*}(u)=\beta\frac{\ell(u)}{(-\ell^{\prime}(u))(\rho-u)}=\left\{\begin{array}[]{ll}\beta\frac{\frac{a}{a+1}-u}{\rho-u}&\hbox{ if }u\leq 0\\ \beta\left(\frac{a}{a+1}\frac{1}{1-u^{a}}(1-u)-\frac{1}{a+1}u\right)\frac{1}{\rho-u}&\hbox{ if }0\leq u\leq 1\\ 0&\hbox{ if }u\geq 1.\end{array}\right.$
By condition (2), we must have
$\frac{-\ell^{\prime}(u)}{\rho-u}\leq\frac{\alpha}{\beta R^{2}}.$
Note that
$\ell^{\prime}(u)=\left\{\begin{array}[]{ll}-1&\hbox{ if }u\leq 0\\ -(1-u^{a})&\hbox{ if }0\leq u\leq 1\\ 0&\hbox{ otherwise}.\end{array}\right.$
Hence, for every $u\leq 0$, it must hold
$\frac{1}{\rho-u}\leq\frac{\alpha}{\beta R^{2}}$
which is equivalent to $\alpha/\beta\leq R^{2}/\rho$. For every $0\leq u\leq 1$, it must hold
$f_{a}(u):=\frac{1-u^{a}}{\rho-u}\leq\frac{\alpha}{\beta R^{2}}.$
Thus, it must hold $\alpha/\beta\geq c_{a,\rho}R^{2}$ where $c_{a,\rho}=\sup_{u\in[0,1]}f_{a}(u)$.
Function $f_{a}$ has boundary values $f_{a}(0)=1/\rho$ and $f_{a}(1)=0$. Furthermore, note
$f_{a}^{\prime}(u)=\frac{1+(a-1)u^{a}-\rho au^{a-1}}{(\rho-u)^{2}}.$
Note $f_{a}(0)=1/{\rho}^{2}>0$ and $f_{a}^{\prime}(1)=-(\rho-1)a<0$. Let $u_{*}$ be such that $f_{a}^{\prime}(u_{*})=0$, which holds if, and only if, $g_{a}(u_{*}):=(a-1){u_{*}}^{a}-\rho a{u_{*}}^{a-1}+1=0$.
Note that
$c_{a,\rho}=\sup_{u\in[0,1]}f_{a}(u)=f_{a}(u_{*})=au_{*}^{a-1}.$
For $a=1$, $f_{1}$ is decreasing on $[0,1]$ hence $c_{1,\rho}=f_{1}(0)=1/\rho$. For $a=2$, $u_{*}$ is a solution of a quadratic equation, and it can be readily shown that $c_{2,\rho}=2\rho(1-\sqrt{{\rho}^{2}-1})$. For every $a\geq 1$, we have $c_{a,\rho}\leq a/(1+a(\rho-1))$. This obviously holds with equality for $a=1$, hence it suffices to show that the inequality holds for $a>1$.
Consider the case $a>1$. Note that $g_{a}(u_{*})=0$ is equivalent to
$au_{*}^{a-1}=\frac{1}{\rho-\frac{a-1}{a}u_{*}}.$ (16)
Combined with the fact $u_{*}\in[0,1]$, it immediately follows
$au_{*}^{a-1}\leq\frac{a}{a(\rho-1)+1}.$
Furthermore, note that $\lim_{a\rightarrow\infty}c_{a,\rho}=1/(\rho-1)$. To see this, consider (16). Note that $u_{*}$ goes to $1$ as $a$ goes to infinity. This can be shown by contradiction as follows. Assume that there exists a constant $c\in[0,1)$ and $a_{0}$ such that $u_{*}\leq c$ for all $a\geq a_{0}$. Then, from (16), $ac^{a-1}\geq 1/\rho^{*}$. The left-hand side in the last inequality goes to 00 as $a$ goes to infinity while the right-hand side is a constant greater than zero, which yields a contradiction. From (16), it follows that $au_{*}^{a-1}$ goes to $1/(\rho-1)$ as $a$ goes to infinity.
We prove the second statement of the theorem as follows. It suffices to show that condition (2) holds true as condition (3) clearly holds for every $\pi$ such that $\pi(u)\geq\pi^{*}(u)$ for every $u\in\mathbb{R}$. For condition (2) to hold, it is sufficient that
$f(u):=\frac{\ell(u)}{\ell^{\prime}(u)^{2}}\geq\frac{\beta R^{2}}{\alpha},\hbox{ for all }u\leq 1.$
Note that
$f(u)=\left\{\begin{array}[]{ll}\frac{a}{a+1}-u&\hbox{ if }u\leq 0\\ \frac{1}{a+1}\frac{u^{a+1}-(a+1)u+a}{(1-u^{a})^{2}}&\hbox{ if }0\leq u\leq 1.\end{array}\right.$
Function $f$ is a decreasing function. This is obviously true for $u\leq 0$. For $0\leq u\leq 1$, we show this as follows. Note that
$f^{\prime}(u)=\frac{2au^{2a}-(2a-1)(a+1)u^{a}+2a^{2}u^{a-1}-(a+1)}{(a+1)(1-u^{a})^{3}}.$
Hence, $f^{\prime}(u)\leq 0$ is equivalent to
$2au^{a-1}(u^{a+1}+a)\leq(a+1)(1-(2a-1)u^{a}).$
In the last inequality, the left-hand side is increasing in $u$ and the right-hand side is decreasing in $u$. Hence the inequality holds for every $u\in[0,1]$ is equivalent to the inequality holding for $u=1$. For $u=1$, the inequality holds with equality.
Note that
$f(1)$ $=$ $\frac{0}{0}$
$=$ $\frac{1}{a+1}\frac{(a+1)u^{a}-(a+1)}{-2(1-u^{a})au^{a-1}}|_{u=1}=\frac{0}{0}$
$=$ $\frac{1}{a+1}\frac{(a+1)au^{a-1}}{2a^{2}u^{2(a-1)}-2(1-u^{a})a(a-1)u^{a-2}}|_{u=1}=\frac{1}{a+1}\frac{(a+1)a}{2a^{2}}$
$=$ $\frac{1}{2a}.$
It follows that $\inf_{u\leq 1}f(u)=f(1)=\frac{1}{2a}\leq\alpha/(\beta R^{2})$, hence it is suffices that $\alpha/\beta\geq R^{2}/(2a)$.
A.5 Proof of Theorem 3.5
Lemma A.3.
For every $a\geq 1$,
$\pi^{*}(u)\leq\pi^{**}(u)=\beta\frac{1-u}{\rho-1}\hbox{ for every }u\leq 1.$
Proof
For $u\leq 1$, $\pi^{*}(u)=\beta(a/(a+1)-u)/(\rho-u)$, so obviously, $\pi^{*}(u)\leq\beta(1-u)/(\rho-u)$. For $0\leq u\leq 1$, we show that next that $a(1-u)/(1-u^{a})-u\leq a(1-u)$, which implies that $\pi^{*}(u)\leq(a/(a+1))\beta(1-u)/(\rho-u)$. To show the asserted inequality, by straightforward calculus it can be shown that the inequality is equivalent to $(1-(1-u))^{1-a}\geq 1-(1-a)(1-u)$ which clearly holds true.
Lemma A.4.
For every $c\in(1,6/5)$, for every $0\leq u\leq 1$, if $(a+1)(1-u)\geq c$, then
$\ell(u)\geq\left(1-\frac{1}{c}\right)(1-u)$
and, otherwise,
$\ell(u)\geq\left(1-\frac{1}{3}\frac{c}{1-c/2}\right)\frac{a}{2}(1-u)^{2}.$
Proof
We first show the first inequality. For $u\leq 1$, it clearly holds $\ell(u)\leq 1-u$, and
$\frac{\frac{a}{a+1}-u}{1-u}=\frac{1-u-\frac{1}{a+1}}{1-u}=1-\frac{1}{(a+1)(1-u)}$
thus, for every $c>1$, $\ell(u)\geq(1-1/c)(1-u)$ whenever $(a+1)(1-u)\geq c$.
For $10\leq u<$ 1, it holds $\ell(u)=1-u-(1/(a+1))(1-u^{a+1})$, hence it clearly holds $\ell(u)\leq 1-u$. Next, note
$\frac{\ell(u)}{1-u}=1\frac{1-u^{a+1}}{(a+1)(1-u)}\geq 1-\frac{1}{(a+1)(1-u)}.$
Hence, again, for every $c>1$, $\ell(u)\geq(1-1/c)(1-u)$ whenever $(a+1)(1-u)\geq c$.
We next show the second inequality. For $0<u\leq 1$, note that $\ell^{\prime}(u)=-(1-u^{a})$, ′′$\ell^{\prime\prime}(u)=au^{a-1}$ and ′′′$\ell^{\prime\prime\prime}(u)=a(a-1)u^{a-2}$. In particular, $\ell^{\prime}(1)=0$, ′′$\ell^{\prime\prime}(1)=a$ and ′′′$\ell^{\prime\prime\prime}(1)=a(a-1)$. By limited Taylor development, for some $u_{0}\in[u,1]$,
$\ell(u)=\frac{a}{2}(1-u)^{2}-\frac{a(a-1)}{6}u_{0}^{a-2}(1-u)^{3}.$
From this, it immediately follows that $\ell(u)\leq\frac{a}{2}(1-u)^{2}$ for every $u\leq 1$. For the case $a\geq 2$, we have
$\ell(u)\geq\frac{a}{2}(1-u)^{2}\left(1-\frac{1}{3}(a-1)(1-u)\right).$
Hence, for every $c\geq 0$, $\ell(u)\geq\frac{a}{2}(1-u)^{2}(1-c/3)$ whenever $(a+1)(1-u)\leq c$. For $1\leq a\leq 2$, we have
$\ell(u)\geq\frac{a}{2}(1-u)^{2}\left(1-\frac{1}{3}\frac{(a-1)(1-u)}{u^{2-a}}\right)\geq\frac{a}{2}(1-u)^{2}\left(1-\frac{1}{3}\frac{(a-1)(1-u)}{u}\right).$
Under $(a+1)(1-u)\leq c$, with $20\leq c<$ 2, we have
$\frac{(a-1)(1-u)}{u}\leq\frac{c}{1-\frac{c}{a+1}}\leq\frac{c}{1-\frac{c}{2}}.$
Hence, it follows
$\ell(u)\geq\frac{a}{2}(1-u)^{2}\left(1-\frac{1}{3}\frac{c}{1-\frac{c}{2}}\right).$
Next, note that for every $x,y,\theta$,
$\pi^{*}(yx^{\top}\theta)\leq\pi^{**}(yx^{\top}\theta)\leq\frac{\beta}{\rho^{*}-1}(1-yx^{\top}\theta)=\frac{\beta}{\rho^{*}-1}(1-\ell^{-1}(\ell(x,y,\theta)))$
where $\ell^{-1}$ is the inverse function of $\ell(u)$ for $u<1$ and $\ell^{-1}(0)=0$.
Hence, we have
$\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(y_{t}x_{t}^{\top}\theta_{t})\right]$$\leq$ $\frac{\beta}{\rho-1}\left(1-\ell^{-1}\left(\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(x_{t},y_{t},\theta_{t})\right]\right)\right)n$
$\leq$ $\frac{\beta}{\rho-1}\left(1-\ell^{-1}\left(c_{a,\rho}\frac{||\theta_{1}-\theta^{*}||^{2}R^{2}}{\beta}\frac{1}{n}\right)\right)n$
where the first inequality is by concavity of the function $1-\ell^{-1}(\ell)$ and the second inequality is by Theorem 3.4.
To apply Lemma A.4, we need that
$\ell^{-1}\left(c_{a,\rho}\frac{||\theta_{1}-\theta^{*}||^{2}R^{2}}{\beta}\frac{1}{n}\right)>0$
which is equivalent to
$c_{a,\rho}\frac{||\theta_{1}-\theta^{*}||^{2}R^{2}}{\beta}\frac{1}{n}<\ell(0)=\frac{a}{a+1}$
i.e.
$n>\frac{a+1}{a}c_{a,\rho}\frac{||\theta_{1}-\theta^{*}||^{2}R^{2}}{\beta}.$
By Lemma A.4, we can distinguish two cases when
$1-\ell^{-1}\left(c_{a,\rho}\frac{||\theta_{1}-\theta^{*}||^{2}R^{2}}{\beta}\frac{1}{n}\right)\geq\frac{c}{a+1}$
or, otherwise, where $c$ is an arbitrary constant in $(1,6/5)$.
In the first case,
$1-\ell^{-1}\left(c_{a,\rho}\frac{||\theta_{1}-\theta^{*}||^{2}R^{2}}{\beta}\frac{1}{n}\right)\leq c_{1}c_{a,\rho}\frac{||\theta_{1}-\theta^{*}||^{2}R^{2}}{\beta}\frac{1}{n}$
where $c_{1}=1/(1-1/c)$ while in the second case
$1-\ell^{-1}\left(c_{a,\rho}\frac{||\theta_{1}-\theta^{*}||^{2}R^{2}}{\beta}\frac{1}{n}\right)\leq c_{2}\sqrt{\frac{2}{a}}\sqrt{c_{a,\rho}\frac{||\theta_{1}-\theta^{*}||^{2}R^{2}}{\beta}\frac{1}{n}}$
where $c_{2}=1/(1-(1/3)c/(1-c/2)$.
It follows that for some constant $\kappa>0$,
$\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(y_{t}x_{t}^{\top}\theta_{t})\right]\leq\kappa\max\left\{\frac{\sqrt{\beta c_{a,\rho}}}{(\rho-1)\sqrt{a}}||\theta_{1}-\theta^{*}||R\sqrt{n},\frac{c_{a,\rho}}{\rho-1}||\theta_{1}-\theta^{*}||^{2}R^{2}\right\}.$
A.6 Multi-class Classification for Linearly Separable Data
We consider multi-class classification with $k\geq 2$ classes. Let $\mathcal{Y}=\{1,\ldots,k\}$ denote the set of classes. For every $y\in\mathcal{Y}$, let $\theta_{y}\in\mathbb{R}^{d}$ and let $\theta=(\theta_{1}^{\top},\ldots,\theta_{k}^{\top})^{\top}\in\mathbb{R}^{kd}$ be the parameter. For given $x$ and $\theta$, predicted class is an element of $\arg\max_{y\in\mathcal{Y}}x^{\top}\theta_{y}$.
The linear separability condition is defined as follows: there exists $\theta^{*}\in\mathbb{R}^{kd}$ such that for some $\rho^{*}>1$, for every $x\in\mathcal{X}$ and $y\in\mathcal{Y}$,
$x^{\top}\theta^{*}_{y}-\max_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}x^{\top}{\theta^{*}_{y^{\prime}}}\geq\rho^{*}.$
Let
$u(x,y,\theta)=x^{\top}\theta_{y}-\max_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}x^{\top}\theta_{y^{\prime}}.$
We consider margin loss functions which are according to a decreasing function of $u(x,y,\theta)$, i.e. $\ell(x,y,\theta)\equiv\ell(u(x,y,\theta))$ and ~$\tilde{\ell}(x,y,\theta)\equiv\tilde{\ell}(u(x,y,\theta))$. For example, this accomodates hinge loss function for multi-class classification Crammer and Singer (2002).
Lemma A.5.
Conditions in Assumption 15 hold provided that for every $x$, $y$ and $\theta$,
$\pi(x,y,\theta)(-\ell^{\prime}(u(x,y,\theta))^{2}2R^{2}\leq\alpha\tilde{\ell}(u(x,y,\theta))$
and
$\pi(x,y,\theta)(-\ell^{\prime}(u(x,y,\theta)))(\rho^{*}-u(x,y,\theta))\geq\beta\tilde{\ell}(u(x,y,\theta)).$
Proof of the lemma
For $x\in\mathcal{X}$ and $y\in\mathcal{Y}$, let $\phi(x,y)=(\phi_{1}(x,y)^{\top},\ldots,\phi_{k}(x,y)^{\top})^{\top}$ where $\phi_{i}(x,y)=x$ if $i=y$ and $\phi_{i}(x,y)$ is the $d$-dimensional null-vector, otherwise. Note that
$u(x,y,\theta)=\theta^{\top}\phi(x,y)-\frac{1}{|\mathcal{Y}^{*}(x,y,\theta)|}\sum_{y^{\prime}\in\mathcal{Y}^{*}(x,y,\theta)}\theta^{\top}\phi(x,y^{\prime})$ (17)
where
$\mathcal{Y}^{*}(x,y,\theta)=\arg\max_{y^{\prime}\in\mathcal{Y}^{*}(x,y,\theta)}\theta_{y^{\prime}}^{\top}x.$
Note that
$||\nabla_{\theta}\ell(x,y,\theta)||^{2}=\ell^{\prime}(u(x,y,\theta))^{2}||\nabla_{\theta}u(x,y,\theta)||^{2}.$
From (17),
$\nabla_{\theta}u(x,y,\theta)=\phi(x,y)-\frac{1}{|\mathcal{Y}^{*}(x,y,\theta)|}\sum_{y^{\prime}\in\mathcal{Y}^{*}(x,y,\theta)}\phi(x,y^{\prime}).$
It can be readily shown that
$||\nabla_{\theta}u(x,y,\theta)||^{2}=\left(1+\frac{1}{|\mathcal{Y}^{*}(x,y,\theta)|}\right)||x||^{2}\leq 2||x||^{2}.$
Hence, we have
$||\nabla_{\theta}\ell(x,y,\theta)||^{2}\leq 2(-\ell^{\prime}(u(x,y,\theta))^{2}||x||^{2}.$ (18)
Next, note that $\nabla_{\theta}\ell(x,y,\theta)=\ell^{\prime}(u(x,y,\theta))\nabla_{\theta}u(x,y,\theta)$,
$\nabla_{\theta}\ell(x,y,\theta)^{\top}(\theta-\theta^{*})=(-\ell^{\prime}(u(x,y,\theta)))\nabla_{\theta}u(x,y,\theta)^{\top}(\theta^{*}-\theta).$
and
$\nabla_{\theta}u(x,y,\theta)^{\top}\theta^{*}=x^{\top}\theta^{*}_{y}-\max_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}x^{\top}{\theta^{*}_{y^{\prime}}}\geq\rho^{*}.$
It follows
$\nabla_{\theta}\ell(x,y,\theta)^{\top}(\theta-\theta^{*})\geq(-\ell^{\prime}(u(x,y,\theta))(\rho^{*}-u(x,y,\theta)).$ (19)
Using (18) and (19), for conditions (14) and (15) to hold, it suffices that
$\pi(x,y,\theta)(-\ell^{\prime}(u(x,y,\theta)))^{2}2R^{2}\leq\alpha\tilde{\ell}(u(x,y,\theta))$
and
$\pi(x,y,\theta)(-\ell^{\prime}(u(x,y,\theta))(\rho^{*}-u(x,y,\theta))\geq\beta\tilde{\ell}(u(x,y,\theta)).$
Note that these conditions are equivalent to those for the binary case in (2) and (3) except for an additional factor $2$ in the first of the last above inequalities.
A.7 Implications of Conditions (2) and (3)
We discuss some implications of conditions (2) and (3). Some of this discussion will tell us about loss functions for which conditions cannot be applied.
First, we note that ~$\tilde{\ell}(u)>0$ implies $\pi(u)>0$. Hence, equivalently, $\pi(u)=0$ implies ~$\tilde{\ell}(u)=0$.
Second, we note that under conditions (2) and (3) it is necessary that for all $u\in\mathbb{R}$, either
$-\ell^{\prime}(u)\leq\frac{\alpha}{\beta R^{2}}\max\{\rho^{*}-u,0\}\hbox{ or }\pi(u)=0.$
Hence, whenever ~$\tilde{\ell}(u)>0$ (and thus $\pi(u)>0$), then $-\ell^{\prime}(u)\leq\frac{\alpha}{\beta R^{2}}\max\{\rho^{*}-u,0\}$. The latter condition means that $\ell^{\prime}(u)=0$ whenever $u\geq\rho^{*}$ and otherwise the derivative of $\ell$ at $u$ is bounded such that $\ell^{\prime}(u)\geq-(\alpha/(\beta R^{2}))(\rho^{*}-u)$. In other words, function $\ell$ must not decrease too fast on $(-\infty,\rho^{*}]$.
Third, assume $\ell$ and $\tilde{\ell}$ are such that $\ell(u)=0$ and ~$\tilde{\ell}(u)=0$ for all $u\geq 1$ and ~$\tilde{\ell}(u)>0$ for all $u<1$. Then, for every $u\leq 1$
[…]
which by integrating is equivalent to
$\ell(u)\leq\frac{\alpha}{\beta R^{2}}\left((\rho^{*}-1)(1-u)+\frac{1}{2}(1-u)^{2}\right).$
This shows that $\ell$ must be upper bounded by a linear combination of hinge and squared hinge loss function.
Forth, assume that $\pi$ is decreasing in $u$, then for every fixed $u_{0}\in\mathbb{R}$,
~$\tilde{\ell}(u)\geq\frac{\pi(u_{0})R^{2}}{\alpha}(-\ell^{\prime}(u))^{2}\hbox{for every }u\leq u_{0}.$
If $\tilde{\ell}=\ell$, then
$\ell(u)\leq\left(\sqrt{\ell(u_{0})}+\frac{1}{2R}\sqrt{\frac{\alpha}{\pi(u_{0})}}(u_{0}-u)\right)^{2}\hbox{ for all }u\leq u_{0}.$
Fifth, and last, assume that $\pi$ is an even function and that there exists $c>0$ and $u_{0}\leq\rho^{*}$ such that ~$\tilde{\ell}(u)\geq c$ for every $u\leq u_{0}$. Then,
$\pi(u)=\Omega\left(\frac{1}{|u|^{2}}\right)$
which limits the rate at which $\pi(u)$ is allowed to decrease with $|u|$. To see, this, from conditions (2) and (3), for every $u\leq\rho^{*}$,
~$\tilde{\ell}(u)\leq\frac{\alpha}{\beta^{2}R^{2}}\pi(u)(\rho^{*}-u)^{2}.$
Hence, $\pi(u)\geq(c\beta^{2}R^{2}/\alpha)/(\rho^{*}-u))$ for every $u\leq u_{0}$. The lower bound is tight in case when $\ell$ is squared hinge loss function and ~$\tilde{\ell}(u)=c$ for every $u\leq u_{0}<1$. In this case, from conditions (2) and (3), for every $u\leq u_{0}$,
$c\beta\frac{1}{(1-u)(\rho^{*}-u)}\leq\pi(u)\leq\frac{c\alpha}{R^{2}}\frac{1}{(1-u)^{2}}$
which implies $\pi(u)=\Theta(1/|u|^{2})$.
A.8 Proof of Lemma 3.6
Function $\Pi$ is a convex function because, by assumption, $\pi$ is an increasing function. By (8 and Jensen’s inequality, we have
$\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\tilde{\ell}(\theta_{t})\right]$ $=$ $\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\Pi(\ell(x_{t},\theta_{t}))\right]$
$\geq$ $\Pi\left(\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(x_{t},\theta_{t})\right]\right)$
$=$ $\Pi\left(\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(x_{t},y_{t},\theta_{t})\right]\right)$
$=$ $\Pi\left(\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]\right).$
Therefore, we have
$\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]\leq\Pi^{-1}\left(\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\tilde{\ell}(\theta_{t})\right]\right).$
Combined with condition (9), we have
$\mathbf{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]$$\leq$ $\Pi^{-1}\left(\inf_{\theta}\tilde{\ell}(\theta)+\sum_{i=1}^{m}f_{i}(n)\right)$
$\leq$ $\inf_{\theta}\Pi^{-1}(\tilde{\ell}(\theta))+\sum_{i=1}^{m}\Pi^{-1}(f_{i}(n))$
where the last inequality holds because $\Pi^{-1}$ is a concave function, and hence, it is a subadditive function.
A.9 Proof of Theorem 3.7
Under assumptions of the theorem, by Theorem 6.3 Bubeck (2015),
$\mathbb{E}\left[\tilde{\ell}(\bar{\theta}_{n})\right]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\tilde{\ell}(\theta_{t})\right]\leq\inf_{\theta}\tilde{\ell}(\theta)+\sqrt{2}S\sigma_{\pi}\frac{1}{\sqrt{n}}+LS^{2}\frac{1}{n}.$
Combining with Lemma 3.6, we obtain the assertion of the theorem.
A.10 Proof of Corollary 3.8
Lemma A.6.
For $\Pi(x)=x-1+e^{-x}$,
$\Pi^{-1}(y)\leq 2\sqrt{y}\hbox{ for }y\in[0,(3/4)^{2}].$
Proof
We consider
$\Pi(x)=x-(1-e^{-x}).$
By limited Taylor development,
$1-e^{-x}\leq x=\frac{1}{2}x^{2}+\frac{1}{6}x^{3}.$
Hence,
$\Pi(x)\geq\frac{1}{2}x^{2}\left(1-\frac{1}{3}x\right).$
Note that $\Pi(x)\geq cx^{2}$ for some constant $c>0$ provided that
$\frac{1}{2}x^{2}\left(1-\frac{1}{3}x\right)\geq cx^{2}$
which is equivalent to $x\leq 3(1-2c)$. Hence, for any fixed $c\in[0,1/2)$, we have
$\Pi(x)\geq cx^{2},\hbox{ for every }x\in[0,3(1-2c)].$
Now, condition $x\leq 3(1-2c)$ is implied by $\sqrt{\Pi(x)/c}\leq 3(1-2c)$, i.e. $\Pi(x)\leq 9c(1-2c)^{2}$. Hence,
$\Pi^{-1}(y)\leq\sqrt{\frac{1}{c}y}\hbox{ for }y\in[0,9c(1-2c)^{2}].$
In particular, by taking $c=1/4$, we have
$\Pi^{-1}(y)\leq 2\sqrt{y}\hbox{ for }y\in[0,(3/4)^{2}].$
We have the bound in Theorem 3.7. Under $\sqrt{2}S\sigma_{\pi}/\sqrt{n}\leq(3/4)^{2}$, we have
$\Pi^{-1}\left(\sqrt{2}S\sigma_{\pi}\frac{1}{\sqrt{n}}\right)\leq 2^{5/4}\sqrt{S\sigma_{\pi}}\frac{1}{\sqrt[4]{n}}.$
Under $LS^{2}/n\leq(3/4)^{2}$, we have
$\Pi^{-1}\left(LS^{2}\frac{1}{n}\right)\leq 2\sqrt{L}S\frac{1}{\sqrt{n}}.$
This completes the proof of the corollary.
A.11 Proof of Lemma 3.10
To simplify notation, we write $\ell_{t}(\theta)\equiv\ell(x_{t},y_{t},\theta)$, $\gamma_{t}=\gamma(x_{t},y_{t},\theta_{t})$ and $\pi_{t}=\pi(x_{t},y_{t},\theta_{t})$.
Since $\ell$ is an $L$-smooth function, we have $\ell(x,y,\theta)-\inf_{\theta^{\prime}}\ell(x,y,\theta^{\prime})\geq 1/(2L^{2})||\nabla_{\theta}\ell(x,y,\theta)||^{2}$. Hence, for any $x,y,\theta$ such that $||\nabla_{\theta}\ell(x,y,\theta)||>0$, we have
$\frac{\ell(x,y,\theta)-\min_{\theta^{\prime}}\ell(x,y,\theta^{\prime})}{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}\geq\frac{1}{2L}.$ (20)
Combined with the definition of $\zeta$ and the fact $\zeta(x_{t},y_{t},\theta_{t})=\gamma_{t}\pi_{t}$, we have
$\kappa:=\beta\min\left\{\frac{1}{2L},\rho\right\}\leq\gamma_{t}\pi_{t}\hbox{ whenever }||\nabla_{\theta}\ell(x,y,\theta)||>0.$ (21)
From the definition of $\zeta$ and the fact $\zeta(x_{t},y_{t},\theta_{t})=\gamma_{t}\pi_{t}$, we have
$\gamma_{t}\pi_{t}\leq\rho\beta.$ (22)
Next, note
$\mathbb{E}[||\theta_{t+1}-\theta^{*}||^{2}\mid x_{t},y_{t},\theta_{t}]$ $=$ $||\theta_{t}-\theta^{*}||^{2}-2\mathbb{E}[z_{t}\mid x_{t},y_{t},\theta_{t}]\nabla_{\theta}\ell_{t}(\theta_{t})^{\top}(\theta_{t}-\theta^{*})+\mathbb{E}[z_{t}^{2}\mid x_{t},y_{t},\theta_{t}]||\nabla_{\theta}\ell_{t}(\theta_{t})||^{2}$
$=$ $||\theta_{t}-\theta^{*}||^{2}-2\gamma_{t}\pi_{t}\nabla_{\theta}\ell_{t}(\theta_{t})^{\top}(\theta_{t}-\theta^{*})+\gamma_{t}^{2}\pi_{t}||\nabla_{\theta}\ell_{t}(\theta_{t})||^{2}$
$\leq$ $||\theta_{t}-\theta^{*}||^{2}-2\gamma_{t}\pi_{t}(\ell_{t}(\theta_{t})-\ell_{t}(\theta^{*}))+\gamma_{t}^{2}\pi_{t}\frac{||\nabla_{\theta}\ell_{t}(\theta_{t})||^{2}}{\ell_{t}(\theta_{t})-\ell_{t}(\theta_{t}^{*})}(\ell_{t}(\theta_{t})-\ell_{t}(\theta^{*}_{t}))$
$=$ $||\theta_{t}-\theta^{*}||^{2}-\gamma_{t}\pi_{t}\left(2-\gamma_{t}\frac{||\nabla_{\theta}\ell_{t}(\theta_{t})||^{2}}{\ell_{t}(\theta_{t})-\ell_{t}(\theta_{t}^{*})}\right)(\ell_{t}(\theta_{t})-\ell_{t}(\theta^{*}_{t}))+2\gamma_{t}\pi_{t}(\ell_{t}(\theta^{*})-\ell_{t}(\theta^{*}_{t}))$
$\leq$ $||\theta_{t}-\theta^{*}||^{2}-2c\kappa(\ell_{t}(\theta_{t})-\ell_{t}(\theta^{*}_{t}))+2\rho\beta(\ell_{t}(\theta^{*})-\ell_{t}(\theta^{*}_{t}))$
where the first inequality is by convexity of $\ell$, the second inequality is by condition (10), and (21) and (22). Hence, we have
$\mathbb{E}[||\theta_{t+1}-\theta^{*}||^{2}]\leq\mathbb{E}[||\theta_{t}-\theta^{*}||^{2}]-2c\kappa_{1}\mathbb{E}[\ell_{t}(\theta_{t})-\ell_{t}(\theta^{*}_{t})]+2\kappa_{0}\mathbb{E}[\ell_{t}(\theta^{*})-\ell_{t}(\theta^{*}_{t})].$
By summing over $t$ from $1$ to $n$, we have
$\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}(\ell_{t}(\theta_{t})-\ell_{t}(\theta^{*}_{t}))\right]$$\leq$ $\frac{\rho\beta}{c\kappa}\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}(\ell_{t}(\theta^{*})-\ell_{t}(\theta^{*}_{t}))\right]+\frac{1}{2c\kappa}||\theta_{1}-\theta^{*}||^{2}\frac{1}{n}$
$\leq$ $\frac{\rho\beta}{c\kappa}(\mathbb{E}[\ell(x,y,\theta^{*})]-\mathbb{E}[\inf_{\theta}\ell(x,y,\theta)])+\frac{1}{2c\kappa}||\theta_{1}-\theta^{*}||^{2}\frac{1}{n}.$
A.12 Proof of Condition (12) for $\pi(x,y,\theta)=\zeta(x,y,\theta)^{\eta}$
It suffices to show that under given conditions, (11) satisfies condition (10). Using (11), condition (10) can be written as follows
$\left(\frac{\ell(x,y,\theta)-\inf_{\theta^{\prime}}\ell(x,y,\theta^{\prime})}{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}\right)^{\eta}\geq\frac{1}{2(1-c)}\min\left\{\beta,\rho\beta\frac{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}{\ell(x,y,\theta)-\inf_{\theta^{\prime}}\ell(x,y,\theta)}\right\}^{1-\eta}.$ (23)
In the inequality (23), by (20), the left-hand side is at least $(1/(2L))^{\eta}$ and clearly the right-hand side is at most $\beta^{1-\eta}/(2(1-c))$. Hence, it follows that it suffices that
$\left(\frac{1}{2L}\right)^{\eta}\geq\frac{1}{2(1-c)}\beta^{1-\eta}.$
A.13 Sampling with Stochastic Polyak’s Step Size for Linear Binary Classifiers
For linear classifiers,
$\frac{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}{\ell(x,y,\theta)}=h(yx^{\top}\theta)||x||^{2}$
where $h(u)=\ell^{\prime}(u)^{2}/\ell(u)$ which plays a pivotal role in condition (10).
For the condition (10) to hold it suffices that
$\pi(x,y,\theta)\geq\frac{\beta}{2(1-c)}\min\left\{\rho R^{2}h(yx^{\top}\theta),1\right\}.$
Note that under assumption that $\ell(u)$ is an $L^{\prime}$-smooth function in $u$, $\ell(yx^{\top}\theta)$ is an $L^{\prime}||x||^{2}$-smooth function in $\theta$. Taking $\rho=1/(2L)$ with $L=L^{\prime}R^{2}$, we have $\rho R^{2}=1/(2L^{\prime})$.
For the binary cross-entropy loss function, we have $h(u)=\sigma^{\prime}(u)^{2}/(\sigma(u)^{2}(-\log(\sigma(u))))$. Specifically, for the logistic regression case
$h(u)=\frac{1}{(1+e^{u})^{2}\log(1+e^{-u})}$ (24)
which is increasing in $u$ for $u\leq 0$ and is decreasing in $u$ otherwise. Note that $h(u)=(1-e^{-\ell(u)})^{2}/\ell(u)$.
We next note some properties of function $h$ which allow us to establish convergence rate guarantees for some loss and uncertainty-based sampling. See Figure 5 for a graphical illustration.
Lemma A.7.
Function $h$, defined in (13), satisfies
$h(u)\leq\frac{1}{1+e^{u}}=1-\sigma(u)\hbox{ for all }u\in\mathbb{R}.$ (25)
Furthermore $h(u)\sim 1-\sigma(u)$ for large $u$.
By Lemma A.7, the condition of Lemma 3.10 is satisfied with $\rho=1/(2L)$, by sampling proportional to absolute error loss $\pi^{*}(u)=\omega(1-\sigma(u))$ with $\beta/(4(1-c)L^{\prime})\leq\omega\leq 1$.
Figure 5: Upper bounds for function $h$ defined in (13): (left) bound of Lemma A.7, (right) bounds of Lemma A.8.
Lemma A.8.
Function $h$, defined in (13), satisifies, for every fixed $a\in(0,1/2]$,
$h(u)\leq\frac{1}{H(a)+(1-a)|u|}\hbox{ for all }u\in\mathbb{R}$ (26)
where $H(a)=a\log(1/a)+(1-a)\log(1/(1-a))$. Furthermore, $h(u)\sim 1/|u|$ as $u$ tends to $\infty$- ∞.
By Lemma A.8, it follows that condition of Lemma 3.10 is satisfied by uncertainty sampling according to
$\pi^{*}(u)=\frac{\beta}{2(1-c)}\min\left\{\rho R^{2}\frac{1}{H(a)+(1-a)|u|},1\right\}.$
A.13.1 Proof of Lemma A.7
We need to prove that for every $u\in\mathbb{R}$,
$(1+e^{u})^{2}\log(1+e^{-u})\geq 1+e^{u}.$
By dividing both sides in the last inequality with $(1+e^{u})^{2}$ and the fact $1/(1+e^{u})=1-1/(1+e^{-u})$, we note that the last above inequality is equivalent to
$\log(1+e^{-u})\geq 1-\frac{1}{1+e^{-u}}.$
By straightforward calculus, this can be rewritten as
$\log\left(1-\left(1-\frac{1}{1+e^{-u}}\right)\right)\leq-\left(1-\frac{1}{1+e^{-u}}\right).$
This clearly holds true because $1-1/(1+e^{-u})\in(0,1)$ and $\log(1-z)\leq-z$ for every $z\in(0,1)$.
It remains only to show that $\lim_{u\rightarrow\infty}h(u)/(1-\sigma(u))=1$. This is clearly true as
$\frac{h(u)}{1-\sigma(u)}=\frac{1}{(1+e^{u})\log(1+e^{-u})}$
which goes to $1$ as $u$ goes to infinity.
A.14 Proof of Lemma A.8
We first consider the case $u\leq 0$. Fix an arbitrary $v\leq 0$. Since $u\mapsto\log(1+e^{u})$ is a convex function it is lower bounded by the tangent passing through $v$, i.e.
$\log(1+e^{-u})\geq\log(1+e^{-v})-\frac{1}{1+e^{v}}(u-v).$
Now, let $a$ be such that $1-a=1/(1+e^{v})$. Since $v\leq 0$, we have $a\in(0,1/2]$. It follows that for any fixed $a\in(0,1/2]$,
$\log(1+e^{-u})\geq H(a)-(1-a)u.$
Using this along with the obvious fact $(1+e^{u})^{2}\geq 1$, we have that for every $u\leq 0$,
$h(u)\leq\frac{1}{\log(1+e^{-u})}\leq\frac{1}{H(a)+(1-a)|u|}.$
We next consider the case $u\geq 0$. It sufficies to show that for every $u\geq 0$, $h(u)\leq h(-u)$, and hence the upper bound established for the previous case applies. The condition $h(u)\leq h(-u)$ is equivalent to
$\frac{1}{(1+e^{u})^{2}\log(1+e^{-u})}\leq\frac{1}{(1+e^{-u})^{2}\log(1+e^{u})}.$
By straightforward calculus, this is equivalent to
$f(u):=(1-e^{-2u})\log(1+e^{u})-u\geq 0.$
This holds because function (i) $f$ is increasing on $[0,u_{0}]$ and decreasing on $[u_{0},\infty)$, for some $u_{0}\geq 0$, (ii) $f(0)=0$ and (iii) $\lim_{u\rightarrow\infty}f(u)=0$. Properties (ii) and (iii) are easy to check. We only show that property (i) holds true. By straightforward calculus,
$f^{\prime}(u)=e^{-2u}(2\log(1+e^{u})-e^{u}).$
It suffices to show that there is a unique $u^{*}\in\mathbb{R}$ such that $f^{\prime}(u^{*})=0$. For any such $u^{*}$ it must hold $2\log(1+e^{u^{*}})-e^{u^{*}}$. Let $v=e^{v^{*}}$. Then, $2\log(1+v)=v$, which is equivalent to
$1+v=e^{\frac{v}{2}}.$
Both sides of the last equation are increasing in $v$, and the left-hand side is larger than the right-hand side for $v=1$. Since the right-hand side is larger than the left-hand side for any large enough $v$, it follows that there is a unique point $v$ at which the sides of the equation are equal. This shows that there is a unique $u^{*}\geq 0$ such that $f^{\prime}(u^{*})=0$.
It remains to show that $\lim_{u\rightarrow-\infty}h(u)/(1/|u|)=1$, i.e.
$\lim_{u\rightarrow-\infty}\frac{-u}{(1+e^{u})^{2}\log(1+e^{-u})}=1$
which clearly holds true as both $1/(1+e^{u})^{2}$ and $-u/\log(1+e^{-u})$ go to $1$ as $u$ goes to $\infty$- ∞.
A.15 Polyak’s Step Size for Multi-class Classification
We consider multi-class classification according to prediction function
$p(y\mid x,\theta)=\frac{e^{x^{\top}\theta_{y}}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{x^{\top}\theta_{y^{\prime}}}},\hbox{ for }y\in\mathcal{Y}.$
Assume that $\ell$ is the cross-entropy function. Let
$u(x,y,\theta)=-\log\left(\sum_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}e^{-(x^{\top}\theta_{y}-x^{\top}\theta_{y^{\prime}})}\right).$
It can be shown that
$\frac{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}{\ell(x,y,\theta)}\leq 2||x||^{2}h(u(x,y,\theta))$
where function $h$ is defined in (13). Hence, condition of Lemma 3.10 holds under
$\pi(u)\geq\frac{\beta}{2(1-c)}\min\left\{2\rho R^{2}h(u),1\right\}.$
For given $\theta$ and $x$, let $\theta_{(1)},\ldots,\theta_{(k)}$ be an ordering of $\theta_{1},\ldots,\theta_{k}$ such that $x^{\top}\theta_{(1)}\geq\cdots\geq x^{\top}\theta_{(k)}$. Sampling according to function $\pi^{*}$ of the gap $g=|x^{\top}\theta_{(1)}-x^{\top}\theta_{(k)}|$,
$\pi^{*}(g)=\frac{\beta}{2(1-c)}\min\left\{2\rho R^{2}h^{*}(g),1\right\},$
where
$h^{*}(g)=\frac{1}{H(a)+(1-a)\max\{g-\log(k-1),0\}},$
satisfies condition of Lemma 3.10.
We next provide proofs for assertions made above. The loss function is assumed to be the cross-entropy loss function, i.e.
$\ell(x,y,\theta)=-\log\left(\frac{e^{x^{\top}\theta_{y}}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{x^{\top}\theta_{y^{\prime}}}}\right).$
Note that we can write
$\ell(x,y,\theta)=-\left(\phi(x,y)^{\top}\theta-\log\left(\sum_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}e^{\phi(x,y^{\prime})^{\top}\theta}\right)\right).$
We consider
$\frac{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}{\ell(x,y,\theta)}$
which is plays a key role in the condition of Lemma 3.10.
It holds
$\nabla_{\theta}\ell(x,y,\theta)=-\left(\phi(x,y)-\frac{\sum_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}e^{\phi(x,y^{\prime})^{\top}\theta}\phi(x,y^{\prime})}{\sum_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}e^{\phi(x,y^{\prime})^{\top}\theta}}\right)$
and
$||\nabla_{\theta}\ell(x,y,\theta)||^{2}$ $=$ $\left(1-\frac{e^{\phi(x,y)^{\top}\theta}}{\sum_{z\in\mathcal{Y}\setminus\{y\}}e^{\phi(x,z)^{\top}\theta}}\right)^{2}||x||^{2}+\sum_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}\left(\frac{e^{\phi(x,y^{\prime})^{\top}\theta}}{\sum_{z\in\mathcal{Y}\setminus\{y\}}e^{\phi(x,z)^{\top}\theta}}\right)^{2}||x||^{2}$
$=$ $\left(\left(1-e^{-\ell(x,y,\theta)}\right)^{2}+\sum_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}\left(e^{-\ell(x,y^{\prime},\theta)}\right)^{2}\right)||x||^{2}.$
From the last equation, it follows
$||x||^{2}\left(1-e^{-\ell(x,y,\theta)}\right)^{2}\leq||\nabla_{\theta}\ell(x,y,\theta)||^{2}\leq 2||x||^{2}\left(1-e^{-\ell(x,y,\theta)}\right)^{2}.$
Note that $\ell(x,y,\theta)=\log(1+e^{-u(x,y,\theta)})$ where
$u(x,y,\theta)=-\log\left(\sum_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}e^{-(x^{\top}\theta_{y}-x^{\top}\theta_{y^{\prime}})}\right).$
It follows
$||x||^{2}h(u(x,y,\theta))\leq\frac{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}{\ell(x,y,\theta)}\leq 2||x||^{2}h(u(x,y,\theta))$
where $h$ is function defined in (13).
The following equation holds
$u(x,y,\theta)=\theta_{y}^{\top}x-\max_{z\in\mathcal{Y}\setminus\{y\}}x^{\top}\theta_{z}-\log\left(\sum_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}e^{-(\max_{z\in\mathcal{Y}\setminus\{y\}}x^{\top}\theta_{z}-x^{\top}\theta_{y^{\prime}})}\right).$
Note that
$|u(x,y,\theta)|$$\geq$ $|x^{\top}\theta_{y}-\max_{z\in\mathcal{Y}\setminus\{y\}}x^{\top}\theta_{z}|-\log\left(\sum_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}e^{-(\max_{z\in\mathcal{Y}\setminus\{y\}}x^{\top}\theta_{z}-x^{\top}\theta_{y^{\prime}})}\right)$
$\geq$ $|x^{\top}\theta_{(1)}-x^{\top}\theta_{(2)}|-\log(k-1).$
Combining with Lemma A.8, for every $a\in(0,1/2]$,
$h(u(x,y,\theta))\leq\frac{1}{H(a)+(1-a)|u|}\leq h^{*}(|x^{\top}\theta_{(1)}-x^{\top}\theta_{(2)}|)$
where
$h^{*}(g)=\left\{\begin{array}[]{ll}\frac{1}{H(a)}&\hbox{ if }g\leq\log(k-1)\\ \frac{1}{H(a)-(1-a)\log(k-1)+(1-a)g}&\hbox{ if }g>\log(k-1).\end{array}\right.$
A.16 Further Details on Experimental Setup
Figure 6: Average cross entropy loss as a function of labeling cost for different sampling methods.
A.16.1 Hyperparameter tuning
We used the Tree-structured Parzen Estimator (TPE) (Bergstra et al., 2011) algorithm in the hyperopt package (Bergstra et al., 2013) to tune the relevant hyperparameters for each method and minimise the average progressive cross entropy loss. For Polyak absloss and Polyak exponent we set the search space of $\eta$ to $[0.01,1]$ and the search space of $\rho$ to $[0,1]$. Note that the values of $\eta$ and $\rho$ influence the rate of sampling. In line with the typical goal of active learning, we aim to learn efficiently and minimise loss under some desired rate of sampling. Therefore, for every configuration of $\eta$ and $\rho$ we use binary search to find the value of $\beta$ that achieves some target empirical sampling rate. Observe that if we would not control for $\beta$, then our hyperparameter tuning setup would simply find values of $\eta$ and $\rho$ that lead to very high sampling rates, which is not in line with the goal of active learning. In the hyperparameter tuning we set the target empirical sampling rate to 50%.
A.16.2 Related to Figure 2
We similarly used binary search to find the value of $\beta$ that correspondingly achieves the two target values 1% and 10% with Polyak power function, while using the values of $\eta$ and $\rho$ that were optimised for a sampling rate of 50%. Therefore, our findings of the gains achieved for selective sampling according to stochastic Polyak’s step size are likely conservative since $\eta$ and $\rho$ were not optimised for specifically these sampling rates.
A.16.3 Related to Figures 3 and 4
We compare Polyak absolute loss sampling to absolute loss sampling and random sampling. In this setting we have no control over the sampling rate of absolute loss sampling. Hence, we first run absolute loss sampling to find an empircal sampling rate of 14.9%. We then again use binary search to find the value of $\beta$ to match this sampling rate with Polyak absolute loss sampling. Again, this setup is conservative with respect to the gains of Polyak absolute loss sampling as $\eta$ and $\rho$ were optimised for a sampling rate of 50%.
Figure 7: Average cross entropy loss on a hold-out testing set for different sampling methods.
Figure 8: Test accuracy for different sampling methods.
A.17 Additional Numerical Results
Below are some additional empirical experiments conducted on the same mushroom classification dataset (Chang and Lin, 2011) described in Section 4. The training loss as a function of the number of labeled instances that were selected for training (i.e., cost) is presented in Figure 6. Performance of different sampling methods on the hold-out testing set is presented in Figure 7 (average test loss) and Figure 8 (test accuracy). As can be seen from the figures, the relative performance of different sampling methods is consistent across different metrics.

### v3

@@
 Abstract
-We consider the convergence rates of loss and uncertainty-based active learning algorithms under various assumptions.
-Firstly, we establish a set of conditions that ensure convergence rates when applied to linear classifiers and linearly separable datasets.
-This includes demonstrating convergence rate guarantees for loss-based sampling with various loss functions.
-Secondly, we introduce a framework that allows us to derive convergence rate bounds for loss-based sampling by leveraging known convergence rate bounds for stochastic gradient descent algorithms.
-Lastly, we propose a new algorithm that combines point sampling and stochastic Polyak’s step size.
-We establish a condition on the sampling process, ensuring a convergence rate guarantee for this algorithm, particularly in the case of smooth convex loss functions.
-Our numerical results showcase the efficiency of the proposed algorithm.
+We investigate the convergence rates and data sample sizes required for training a machine learning model using a stochastic gradient descent (SGD) algorithm, where data points are sampled based on either their loss value or uncertainty value.
+These training methods are particularly relevant for active learning and data subset selection problems.
+For SGD with a constant step size update, we present convergence results for linear classifiers and linearly separable datasets using squared hinge loss and similar training loss functions.
+Additionally, we extend our analysis to more general classifiers and datasets, considering a wide range of loss-based sampling strategies and smooth convex training loss functions.
+We propose a novel algorithm called Adaptive-Weight Sampling (AWS) that utilizes SGD with an adaptive step size that achieves stochastic Polyak’s step size in expectation.
+We establish convergence rate results for AWS for smooth convex training loss functions.
+Our numerical experiments demonstrate the efficiency of AWS on various datasets by using either exact or estimated loss values.
 1 Introduction
-In practice, when training machine learning models for prediction tasks (classification or regression), one often has an abundance of unlabeled data available, but obtaining the corresponding labels may come at a high cost.
+In practice, when training machine learning models for prediction tasks (classification or regression), one often has access to an abundance of unlabeled data, while obtaining the corresponding labels may entail high costs.
 This may especially be the case in fields like computer vision, natural language processing, and speech recognition.
 Active learning algorithms are designed to efficiently learn a prediction model by employing a label acquisition, with the goal of minimizing the number of labels used to train an accurate prediction model.
-Various label acquisition strategies have been proposed, each aiming to select informative points for the underlying model training task.
-For instance, active learning strategies include query-by-committee (Seung et al., 1992), expected model change (Settles et al., 2007), expected error reduction (Roy and McCallum, 2001), expected variance reduction (Wang et al., 2016), and mutual information maximization (Kirsch et al., 2019; Kirsch and Gal, 2022).
+Various label acquisition strategies have been proposed, each aiming to select informative points for the underlying model training task; including query-by-committee (Seung et al., 1992), expected model change (Settles et al., 2007), expected error reduction (Roy and McCallum, 2001), expected variance reduction (Wang et al., 2016), and mutual information maximization (Kirsch et al., 2019; Kirsch and Gal, 2022).
 A common label acquisition strategy involves estimating uncertainty, which can be viewed as self-disagreement about predictions made by a given model.
 Algorithms using an uncertainty acquisition strategy are referred to as uncertainty-based active learning algorithms.
+Different variants of uncertainty strategies include margin of confidence, least confidence, and entropy-based sampling (Nguyen et al., 2022).
 Recently, a loss-based active learning approach gained attention in research Yoo and Kweon (2019), Lahlou et al.
@@
 (2021), and is now applied at scale in industry, such as for training integrity violation classifiers at Meta.
-This method involves selecting points for which there is a disagreement between the prediction model and an oracle, as measured by a loss function.
-Since the true loss of a data point is unknown prior to the acquisition of the label, in practice it is estimated using supervised learning.
+This method involves selecting points for which there is a disagreement between the predicted label and the true label, as measured by a loss function.
+Since the true loss of a data point is unknown prior to the acquisition of the label, in practice, it is estimated using supervised learning.
 Loss-based sampling aligns with the spirit of the perceptron algorithm (Rosenblatt, 1958), which updates the model only for falsely-classified points.
-Convergence guarantees for some uncertainty sampling strategies have recently been established, such as for margin of confidence sampling (Raj and Bach, 2022).
-As for loss-based active learning algorithms, there are only limited results on their convergence properties, as they have been started to be studied only recently.
-e.g.
-Liu and Li (2023).
-The primary focus of this paper is to establish convergence guarantees for loss and uncertainty-based learning algorithms.
-We delve into stream-based algorithms, where model training entails the use of a stochastic gradient descent algorithm with a loss or an uncertainty-based acquisition function.
-Our work provides new results on conditions that ensure the convergence of loss- and uncertainty-based active learning algorithms, along with bounds on convergence rates, accommodating various loss functions.
-In the case of loss-based acquisition strategies, we assume that the active learner has access to an oracle providing an unbiased estimate of the conditional expected loss for a point, given the feature vector of the point and the current prediction model.
-In practical settings, the loss cannot be evaluated at acquisition time since labels are yet unknown.
-Instead, a separate prediction model can be devised, which can estimate the loss subject to estimation noise.
-In our experiments, we assess the impact of the noise in such a loss estimator for various loss-based active learning strategies.
-It is noteworthy that uncertainty and loss-based acquisition strategies are also of interest for the data subset selection problem, often referred to as core-set selection or data pruning.
+Convergence guarantees for some uncertainty-based active learning algorithms have recently been established, such as for margin of confidence sampling (Raj and Bach, 2022).
+By contrast, there are only limited results on the convergence properties of loss-based active learning algorithms, as these only recently been started to be studied, e.g., Liu and Li (2023).
+The primary focus of this paper is to establish convergence guarantees for stochastic gradient descent (SGD) algorithms where points are sampled based on their loss.
+Our work provides new results on conditions that ensure certain convergence rates and bounds on the expected sample size, accommodating various data sampling strategies.
+Our theoretical results are under assumption that the active learner has access to an oracle that provides unbiased estimate of the conditional expected loss for a point, given the feature vector of the point and the current prediction model.
+In practice, the loss cannot be evaluated at acquisition time since labels are yet unknown.
+Instead, a separate prediction model is used for loss estimation.
+In our experiments, we assess the impact of the bias and noise in such a loss estimator.
+Our convergence rate analysis accommodates also uncertainty-based data selection, for which we provide new results.
+Uncertainty and loss-based acquisition strategies are also of interest for the data subset selection problem, often referred to as core-set selection or data pruning.
 This problem involves finding a small subset of training data such that the predictive performance of a classifier trained on it is close to that of a classifier trained on the full training data.
@@
 (2022).
-In such scenarios, the oracle can evaluate an underlying loss function exactly, avoiding the need for a noisy estimator.
-Our contributions can be summarized as follows:
-$\bullet$ We outline a set of conditions under which a non-asymptotic convergence rate of order $O(1/n)$ holds, where $n$ represents the number of iterations of the algorithm, i.e., the number of unlabeled points presented to the algorithm.
-These conditions enable us to establish convergence rate results for loss-based sampling in linearly separable datasets, with the loss function taking on various forms such as the squared hinge loss function, generalized hinge loss function, or satisfying other specified conditions.
+In such scenarios, the oracle can evaluate an underlying loss function exactly, avoiding the need for using a loss estimator.
+There is a large body of work on convergence of SGD algorithms, e.g.
+see Bubeck (2015); Nesterov (2018).
+These results are established for SGD algorithms under either constant, diminishing or adaptive step sizes.
+Recently, Loizou et al.
+(2021), studied SGD with the stochastic Polyak’s step size, depending on the ratio of the loss and the squared gradient of the loss of a point.
+Our work is different as we consider convergence of SGD under sampling of points.
+1.1 Summary of our Contributions
+Our contributions can be summarizes as given in the following points:
+$\bullet$ For SGD with a constant step size, we present conditions under which a non-asymptotic convergence rate of order $O(1/n)$ holds, where $n$ represents the number of iterations of the algorithm, i.e., the number of unlabeled points presented to the algorithm.
+These conditions enable us to establish convergence rate results for loss-based sampling in the case of linear classifiers and linearly separable datasets, with the loss function taking on various forms such as the squared hinge loss function, generalized hinge loss function, or satisfying other specified conditions.
 Our results provide bounds for both expected loss and the number of sampled points, encompassing different loss-based strategies.
-$\bullet$ The results discussed above for loss-based strategies, concerning linear classifiers and linearly separable datasets, rely on a key lemma that may be of independent interest.
-The conditions of the lemma can be verified for various loss and uncertainty-based acquisition strategies.
-The convergence rate established in Raj and Bach (2022) for an uncertainty-based strategy follows from this lemma, with a specific choice of the loss function used for training and the loss function used for evaluation.
-The lemma is derived by generalizing an argument used in the proof of Theorem 3.1 in Raj and Bach (2022) to accommodate different acquisition strategies and loss functions.
-$\bullet$ We introduce a framework for establishing convergence rate bounds when sampling points according to an arbitrary, increasing function of the conditional expected loss of a point.
-This framework is particularly noteworthy as it extends the applicability to a broader set of loss functions compared to our initial conditions.
-Leveraging the observation that the algorithm behaves akin to a stochastic gradient descent algorithm with an underlying objective function, as referred to as an ”equivalent loss” in Liu and Li (2023), allows us to apply known convergence rate results for stochastic gradient descent algorithms.
-We present a condition under which explicit bounds on the convergence rate can be established in terms of the sampling probability function, and we provide conditions for this to hold for smooth convex loss functions.
-$\bullet$ We propose a learning algorithm that combines a sampling-based acquisition strategy with an adaptive step-size stochastic gradient descent update, achieving the stochastic Polyak’s step size update Loizou et al.
-(2021) in expectation.
-We establish a condition on the sampling strategy under which a non-asymptotic convergence rate of order $O(1/n)$ holds for smooth convex loss functions.
+These results are established by using a convergence rate lemma that may be of independent interest.
+$\bullet$ For SGD with a constant step size, we provide new convergence rate results for more general classifiers and datasets, with sampling of points according to an increasing function $\pi$ of the conditional expected loss of a point.
+In this case, we present conditions for smooth convex training loss functions under which a non-asymptotic convergence rate of order $O(\Pi^{-1}(1/\sqrt{n}))$ holds, where $\Pi$ is the primitive function of $\pi$.
+These results are established by leveraging the fact that the algorithm behaves akin to a SGD algorithm with an underlying objective function, as referred to as an equivalent loss in Liu and Li (2023), allowing us to apply known convergence rate results for SGD algorithms.
+$\bullet$ We propose Adaptive-Weight Sampling (AWS), a novel learning algorithm that combines a sampling-based acquisition strategy with an adaptive step-size SGD update, achieving the stochastic Polyak’s step size update in expectation, which can be used with any differentiable loss function.
+We establish a condition under which a non-asymptotic convergence rate of order $O(1/n)$ holds for AWS with smooth convex loss functions.
 We present uncertainty and loss-based strategies that satisfy this condition for binary classification, as well as an uncertainty strategy for multi-class classification.
-$\bullet$ We present numerical results that demonstrate the efficiency of sampling with stochastic Polyak’s step size and its robustness to loss estimation noise.
+$\bullet$ We present numerical results that demonstrate the efficiency of AWS on various datasets.
 1.2 Related Work
@@
 However, none of these works have provided theoretical convergence guarantees.
-Different variants of uncertainty sampling include margin of confidence sampling, least confidence sampling, and entropy-based sampling (Nguyen et al., 2022).
-The convergence of margin of confidence sampling was recently studied by Raj and Bach (2022), who showed linear convergence for the hinge loss function for a family of selection probability functions and an algorithm that performs a stochastic gradient descent update with respect to the squared hinge loss function.
+The convergence of margin of confidence sampling was recently studied by Raj and Bach (2022), who demonstrated linear convergence for linear classifiers and linearly separable datasets, specifically for the hinge loss function, for a family of selection probability functions and an algorithm that performs a SGD update with respect to the squared hinge loss function.
+However, our results for linear classifiers and linearly separable datasets differ, as our focus lies on loss-based sampling strategies and providing bounds on the convergence rate of a loss function and the expected number of sampled points.
+These results are established using a convergence rate lemma, which may be of independent interest.
+It is noteworthy that the convergence rate for uncertainty-based sampling, as in Theorem 3.1 of Raj and Bach (2022), can be derived by checking the conditions of the convergence rate lemma.
 A loss-based active learning algorithm was proposed by Yoo and Kweon (2019), comprising a loss prediction module and a target prediction model.
@@
 Recent analysis of convergence for loss and uncertainty-based active learning strategies has been presented by Liu and Li (2023).
-Specifically, they introduced the concept of an equivalent loss, demonstrating that a gradient descent algorithm employing point sampling can be viewed as a stochastic gradient descent algorithm optimizing an equivalent loss function.
-While they focused on discussing specific cases like sampling proportional to conditional expected loss for loss-based acquisition strategies, we introduce a framework that accommodates sampling based on any continuous increasing function of expected conditional loss.
-Our framework provides explicit convergence rate bounds in terms of the underlying sampling probability function.
+Specifically, they introduced the concept of an equivalent loss, demonstrating that a gradient descent algorithm employing point sampling can be viewed as a SGD algorithm optimizing an equivalent loss function.
+While they focused on specific cases like sampling proportional to conditional expected loss, our results allow for sampling based on any continuous increasing function of expected conditional loss, and provide explicit convergence rate bounds in terms of the underlying sampling probability function.
 In addition, Loizou et al.
-(2021) introduced a stochastic gradient descent algorithm featuring an adaptive stochastic Polyak’s step size, which has theoretical convergence guarantees under various assumptions.
+(2021) introduced a SGD algorithm featuring an adaptive stochastic Polyak’s step size, which has theoretical convergence guarantees under various assumptions.
 This algorithm showcased robust performance in comparison to state-of-the-art optimization methods, especially when training over-parametrized models.
@@
 Furthermore, we extend this condition to hold for an uncertainty sampling strategy designed for multi-class classification.
-2 Background
-Algorithm
-We consider the projected stochastic gradient descent algorithm defined as: given an initial value $\theta_{1}$,
+2 Problem Statement
+We consider the projected SGD algorithm defined as: given an initial value $\theta_{1}\in\Theta$,
 $\theta_{t+1}=\mathcal{P}_{\Theta_{0}}\left(\theta_{t}-z_{t}\nabla_{\theta}\ell(x_{t},y_{t},\theta_{t})\right),\hbox{ for }t\geq 1$ (1)
-where $z_{t}$ is a stochastic step size with mean $\zeta(x_{t},y_{t},\theta_{t})$ for some function $\zeta:\mathcal{X}\times\mathcal{Y}\times\Theta\mapsto\mathbb{R}_{+}$, $\Theta_{0}\subseteq\Theta$, and $\mathcal{P}_{\Theta}$ is the projection function, i.e., $\mathcal{P}_{\Theta_{0}}(u)=\arg\min_{v\in\Theta\cap\Theta_{0}}||u-v||$.
+where $\ell:\mathcal{X}\times\mathcal{Y}\times\Theta\rightarrow\mathbb{R}$ is a training loss function, $z_{t}$ is a stochastic step size with mean $\zeta(x_{t},y_{t},\theta_{t})$ for some function $\zeta:\mathcal{X}\times\mathcal{Y}\times\Theta\mapsto\mathbb{R}_{+}$, $\Theta_{0}\subseteq\Theta$, and $\mathcal{P}_{\Theta_{0}}$ is the projection function, i.e., $\mathcal{P}_{\Theta_{0}}(u)=\arg\min_{v\in\Theta_{0}}||u-v||$.
 Unless specified otherwise, we consider the case $\Theta_{0}=\Theta$, which requires no projection.
-For the choice of the stochastic step size, we consider two cases: (a) Bernoulli sampling with a constant step size, and (b) stochastic Polyak’s step size.
+For binary classification tasks, we assume $\mathcal{Y}=\{-,1,1\}$.
+For every $t>0$, we define $\bar{\theta}_{t}=(1/t)\sum_{s=1}^{t}\theta_{s}$.
+For the choice of the stochastic step size, we consider two cases: (a) Constant-Weight Sampling: a Bernoulli sampling with a constant step size, and (b) Adaptive-Weight Sampling: a sampling that achieves stochastic Polyak’s step size in expectation.
 For case (a), $z_{t}$ is the product of a constant step size $\gamma$ and a Bernoulli random variable with mean $\pi(x_{t},y_{t},\theta_{t})$.
-For case (b), $\zeta(x,y,\theta)$ is the ”stochastic” Polyak’s step size, and $z_{t}$ is equal to $\zeta(x_{t},y_{t},\theta_{t})/\pi(x_{t},y_{t},\theta_{t})$ with probability $\pi(x_{t},y_{t},\theta_{t})$ and is equal to 00 otherwise.
+For case (b), $\zeta(x,y,\theta)$ is the "stochastic" Polyak’s step size, and $z_{t}$ is equal to $\zeta(x_{t},y_{t},\theta_{t})/\pi(x_{t},y_{t},\theta_{t})$ with probability $\pi(x_{t},y_{t},\theta_{t})$ and is equal to 00 otherwise.
 Note that using the notation $\pi(x,y,\theta)$ allows for the case when the sampling probability does not depend on the value of the label $y$.
-Linearly Separable Data
-For binary classification tasks, $\mathcal{Y}=\{-1,1\}$.
-We say that data is separable if, for every point $(x,y)\in\mathcal{X}\times\mathcal{Y}$, either $y=1$ with probability $1$ or $y=-1$ with probability $1$.
+For a loss-based sampling, $\pi$ is an increasing function of some loss function $\ell^{\star}$, which does not necessarily correspond to the training loss function $\ell$.
+Specifically, for a binary classifier with $p(x,y,\theta)$ denoting the expected prediction label, sampling proportional to the absolute error loss is defined as $\pi(\ell^{*})=\omega\ell^{*}$ where $\ell^{*}(x,y,\theta)=|y-p(x,y,\theta)|$ and $\omega\in(0,1/2]$.
+For an uncertainty-based sampling, $\pi$ is a function of some quantity reflecting the uncertainty of the prediction model.
+Our focus is on finding convergence conditions for algorithm (1) and convergence rates under these conditions, as well as bounds on the expected number of points sampled by the algorithm.
+Additional Assumptions and Notation
+For binary classification, we say that data is separable if, for every point $(x,y)\in\mathcal{X}\times\mathcal{Y}$, either $y=1$ with probability $1$ or $y=-1$ with probability $1$.
 The data is linearly separable if there exists $\theta^{*}\in\Theta$ such that $y=\mathrm{sgn}(x^{\top}\theta^{*})$ for every $x\in\mathcal{X}$.
 Linearly separable data has a $\rho^{*}$-margin if $|x^{\top}\theta^{*}|\geq\rho^{*}$ for every $x\in\mathcal{X}$, for some $\theta^{*}\in\Theta$.
-Linear Classifiers
 Some of our results are for linear classifiers, where the predicted label of a point $x$ is a function of $x^{\top}\theta$.
@@
 For logistic regression, the predicted label is $1$ with probability $\sigma(x^{\top}\theta)$ and $-1$ otherwise, where $\sigma$ is the logistic function defined as $\sigma(z)=1/(1+e^{-z})$.
-Smooth Loss Functions
+For binary classification, we model the prediction probability of the positive label as $\sigma(x^{\top}\theta)$, where $\sigma:\mathbb{R}_{+}\rightarrow[0,1]$ is an increasing function, and $\sigma(-u)+\sigma(u)=1$ for all $u\in\mathbb{R}$.
+The absolute error loss takes value $1-\sigma(x^{\top}\theta)$ if $y=1$ or value $\sigma(x^{\top}\theta)$ if $y=-1$, which corresponds to $1-\sigma(yx^{\top}\theta)$.
+The binary cross-entropy loss for a point $(x,y)$ under model parameter $\theta$ can be written as $\ell(x,y,\theta)=-\log(\sigma(yx^{\top}\theta))$.
+Hence, absolute error loss-based sampling corresponds to the sampling probability function $\pi(\ell)=1-e^{-\ell}$.
 For any given $(x,y)\in\mathcal{X}\times\mathcal{Y}$, the loss function $\ell(x,y,\theta)$ is considered smooth on $\Theta^{\prime}\subseteq\Theta$ if it has a Lipschitz continuous gradient on $\Theta^{\prime}$, i.e., there exists $L_{x,y}$ such that $||\nabla_{\theta}\ell(x,y,\theta_{1})-\nabla_{\theta}\ell(x,y,\theta_{2})||\leq L_{x,y}||\theta_{1}-\theta_{2}||$ for all $\theta_{1},\theta_{2}\in\Theta^{\prime}$.
@@
 3 Convergence Rate Guarantees
-In this section, we present conditions on the stochastic step size of algorithm (1) under which we can bound the total expected loss and, in some cases, bound the expected number of samples.
-For Bernoulli sampling with a constant step size, we first provide conditions that allow us to derive bounds for linear classifiers and linearly separable data (Section 3.1).
-Next, we provide bounds for loss-based sampling, leveraging convergence results known to hold for stochastic gradient descent algorithms (Section 3.2).
-For sampling combined with stochastic Polyak’s step size, we offer a condition that allows us to establish convergence bounds for loss and uncertainty-based sampling (Section 3.1).
-3.1 Bernoulli Sampling with Constant Step Size: First Set of Convergence Conditions
-The convergence rate results presented in this section are established using Lemma A.2, which, for space reasons, is provided in Appendix A.1.
-These results pertain to linear classifiers and linearly separable datasets.
-We initially address binary classification and subsequently extend our considerations to multi-class classification.
-The conditions for convergence involve both the loss function $\ell$ utilized by algorithm (1) and a separate loss function $\tilde{\ell}$ employed for evaluating the algorithm’s performance.
-Binary Classification
-We consider binary classification based on the sign of $x^{\top}\theta$.
-With a slight abuse of notation, let $\ell(x,y,\theta)\equiv\ell(u)$, ~$\tilde{\ell}(x,y,\theta)\equiv\tilde{\ell}(u)$, and $\pi(x,y,\theta)\equiv\pi(u)$ where $u=yx^{\top}\theta$.
-We assume that the domain $\mathcal{X}$ is bounded, i.e., there exists $R$ such that $||x||\leq R$ for all $x\in\mathcal{X}$, $||\theta_{1}-\theta^{*}||\leq S$, and that the data is $\rho^{*}$-margin linearly separable (see Section 2 for the definition).
-Under prevailing assumptions, the following conditions imply conditions of Lemma A.2: for all $u\in\mathbb{R}$,
-$\pi(u)\ell^{\prime}(u)^{2}R^{2}\leq\alpha\tilde{\ell}(u)$ (2)
-and
-$\pi(u)(-\ell^{\prime}(u))(\rho^{*}-u)\geq\beta\tilde{\ell}(u).$ (3)
-We demonstrate convergence results for a class of loss functions and sampling proportional to either zero-one or absolute error loss, and then for the squared hinge loss function and a smoothed hinge function (which accommodates the hinge loss function as a limit case).
-For any $x$, $y$, and $\theta$, the zero-one loss and absolute error loss are defined as $1_{\{y\neq\mathrm{sgn}(x^{\top}\theta)\}}=\ell_{01}(u):=1_{\{u\leq 0\}}$ and $|y-\mathrm{sgn}(x^{\top}\theta)|=\ell_{\mathrm{abs}}(u):=2\left(1_{\{u<0\}}+\frac{1}{2}1_{\{u=0\}}\right)$, respectively, where $u=yx^{\top}\theta$.
-We first consider sampling proportional to either zero-one loss or absolute error loss when the loss function satisfies the following conditions:
-Assumption 3.1.
-Function $\ell$ is continuously differentiable on $(-\infty,0]$, convex, and $\ell^{\prime}(0)\leq-c_{1}$ and $\lim_{u\rightarrow-\infty}\ell^{\prime}(u)\geq-c_{2}$, for some constants $c_{1},c_{2}>0$.
-In particular, conditions (2) and (3) hold for the case when the loss function $\ell$ is the hinge loss function.
-For this case, the sampling probability equal to zero-one loss makes algorithm (1) correspond to the well-known perceptron algorithm.
-This establishes a precise relation between loss-based sampling and the perceptron algorithm.
-Theorem 3.2.
-Assume that the loss function $\ell$ satisfies Assumption 3.1 and $\rho^{*}>0$.
-Then, under sampling proportional to zero-one loss with a constant factor $\omega\in(0,1]$, for any initial value $\theta_{1}$ such that $||\theta_{1}-\theta^{*}||\leq S$ and $\{\theta_{t}\}_{t>1}$ according to algorithm (1) with $\gamma=c_{1}\rho^{*}/(c_{2}^{2}R^{2})$,
-$\mathbb{E}\left[\sum_{t=1}^{n}\ell_{01}(y_{t}x_{t}^{\top}\theta_{t})\right]\leq\frac{c_{2}^{2}R^{2}S^{2}}{c_{1}^{2}\omega}\frac{1}{{\rho^{*}}^{2}}.$ (4)
-Furthermore, under sampling proportional to absolute error loss with a factor $\omega\in(0,1/2]$ and $\gamma=2c_{1}\rho^{*}/(c_{2}^{2}R^{2})$, the bound in (4) holds but with an additional factor of $2$.
-The bound in (4) is a well-known bound on the number of mistakes made by the perceptron algorithm, of the order $O(1/{\rho^{*}}^{2})$.
-It can be readily observed that under sampling proportional to zero-one loss, the expected number of sampled points is bounded by $(c_{2}/c_{1})^{2}R^{2}S^{2}/{\rho^{*}}^{2}$, which also holds for sampling proportional to absolute error loss but with an additional factor of $4$.
-We next consider sampling for the squared hinge loss function $\ell(u)=(1/2)\max\{1-u,0\}^{2}$.
+In this section, we present conditions on the stochastic step size of algorithm (1) under which we can bound the total expected loss and the expected number of samples.
+For the Constant-Weight Sampling, we provide conditions that allow us to derive bounds for linear classifiers and linearly separable datasets and more general cases.
+For Adaptive-Weight Sampling, we offer a condition that allows us to establish convergence bounds for both loss and uncertainty-based sampling.
+3.1 Constant-Weight Sampling
+Linear Classifiers and Linearly Separable Datasets
+We focus on binary classification and briefly discuss extension to multi-class classification.
+We consider the linear classifier with the predicted label $\hbox{sgn}(x^{\top}\theta)$.
+With a slight abuse of notation, let $\ell(x,y,\theta)\equiv\ell(u)$ and $\pi(x,y,\theta)\equiv\pi(u)$ where $u=yx^{\top}\theta$.
+We assume that the domain $\mathcal{X}$ is bounded, i.e., there exists $R$ such that $||x||\leq R$ for all $x\in\mathcal{X}$, $||\theta_{1}-\theta^{*}||\leq S$ for some $S\geq 0$, and that the data is $\rho^{*}$-margin linearly separable.
+We present convergence rate results for the training loss function corresponding to the squared hinge loss function, i.e.
+$\ell(u)=(1/2)\max\{1-u,0\}^{2}$.
+Our additional results also cover other cases, including a class of smooth convex loss functions and a generalized smooth hinge loss function, which are presented in the Appendix.
 Theorem 3.1.
@@
 The expected number of samples is $O(\sqrt{n})$ concerning the number of iterations and is $O(1/(\rho^{*}-1))$ concerning the margin $\rho^{*}-1$.
-Next, we consider the generalized smooth hinge loss function (Rennie, 2005), defined for $a\geq 1$ as follows:
-$\ell(u)=\left\{\begin{array}[]{ll}\frac{a}{a+1}-u&\text{ if }u\leq 0\\ \frac{a}{a+1}-u+\frac{1}{a+1}u^{a+1}&\text{ if }0\leq u\leq 1\\ 0&\text{ otherwise}.\end{array}\right.$ (6)
-This is a continuously differentiable function that converges to the value of the hinge loss function, $\max\{1-u,0\}$, as $a$ goes to infinity.
-The family of loss functions parameterized by $a$ accommodates the smooth hinge loss function with $a=1$, introduced by Rennie and Srebro (2005).
-Theorem 3.4.
-Assume that $\rho^{*}>1$, the loss function is the generalized smooth hinge loss function, and the sampling probability is according to the function $\pi^{*}$, which, for $\beta\in(0,1]$ and $\rho\in(1,\rho^{*}]$, is defined as
-$\pi^{*}(u)=\left\{\begin{array}[]{ll}\beta\frac{\frac{a}{a+1}-u}{\rho-u}&\text{ if }u\leq 0\\ \beta\frac{\frac{a}{a+1}\frac{1}{1-u^{a}}(1-u)-\frac{1}{a+1}u}{\rho-u}&\text{ if }0\leq u\leq 1\\ 0&\text{ if }u\geq 1.\end{array}\right.$
-Then, for any initial value $\theta_{1}$ such that $||\theta_{1}-\theta^{*}||\leq S$ and $\{\theta_{t}\}_{t>1}$ according to algorithm (1) with $\gamma=1/(c_{a,\rho}R^{2})$, where $c_{a,\rho}=a/(a(\rho-1)+1)$, we have
-$\mathbb{E}\left[\ell(yx^{\top}\bar{\theta}_{n})\right]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(y_{t}x_{t}^{\top}\theta_{t})\right]\leq c_{a,\rho}\frac{R^{2}S^{2}}{\beta}\frac{1}{n}.$
-Furthermore, the same bound holds for every $\pi$ such that $\pi^{*}(u)\leq\pi(u)\leq\beta$ for all $u\in\mathbb{R}$, with $c_{a,\rho}=2a$.
-Note that the sampling probability function $\pi^{*}$ is increasing in $a$ and is upper-bounded by $\pi^{**}(u)=\beta(1-u)/(\rho-1)$, which may be regarded as the limit for the hinge loss function.
-See Figure 1 for an illustration.
-We remark that the expected loss bound in Theorem 3.4 is the same as for the squared hinge loss function in Theorem 3.3, except for an additional factor $c_{a,\rho}=a/(a(\rho-1)+1)$.
-This factor is increasing in $a$ but always lies in the interval $[1/\rho,1/(\rho-1)]$, where the boundary values are achieved for $a=1$ and $a\rightarrow\infty$, respectively.
-Figure 1: Sampling probability function for the family of generalized smooth hinge loss functions.
-Theorem 3.5.
-Assuming $\rho^{*}>1$, the loss function is the generalized smooth hinge loss, and the sampling probability function $\pi^{*}$ satisfies the assumptions in Theorem 3.4, with $n>((a+1)/a)c_{a,\rho}R^{2}S^{2}/\beta$, the expected number of sampled points is bounded as
-$\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(y_{t}x_{t}^{\top}\theta_{t})\right]\leq\kappa\max\left\{\frac{\sqrt{\beta c_{a,\rho}}RS}{(\rho-1)\sqrt{a}}\sqrt{n},\frac{c_{a,\rho}R^{2}S^{2}}{\rho-1}\right\},$
-where $\kappa$ is some positive constant.
-For any fixed number of iterations $n$ satisfying the condition of the theorem, the expected number of sampled points is bounded by a constant for sufficiently large value of parameter $a$.
-The bound in Theorem 3.5 depends on how the loss function $\ell(u)$ varies with $u$.
-When $a$ is large, $\ell(u)$ is approximately $1-u$ (hinge loss), otherwise, it is approximately $\frac{a}{2}(1-u)^{2}$ for $0\leq u\leq 1$ (squared hinge loss).
-Multi-class Classification
-The conditions for multi-class classification with the set of classes $\mathcal{Y}$ are the same as for binary classification, with $u(x,y,\theta):=x^{\top}\theta_{y}-\max_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}x^{\top}\theta_{y^{\prime}}$, except for an additional factor of $2$ in the left-hand side of the inequality in the first condition (see Lemma A.5 in the Appendix).
+Theorem 3.1, and our other results for linear classifiers and linearly separable datasets, are established using a convergence rate lemma, which is presented in Appendix A.2, along with its proof.
+This lemma generalizes the conditions used to establish the convergence rate for an uncertainty-based sampling algorithm by Raj and Bach (2022), with the sampling probability function $\pi(u)=1/(1+\mu|u|)$, for some constant $\mu>0$.
+It can be readily shown that Theorem 3.1 in Raj and Bach (2022) follows from our convergence rate lemma with the training loss function corresponding to the squared hinge loss function and the evaluation loss function (used for convergence rate guarantee) corresponding to the hinge loss function.
+Further details on the convergence rate lemma are discussed in Appendix A.2.1.
+The convergence rate conditions for multi-class classification with the set of classes $\mathcal{Y}$ are the same as for binary classification, with $u(x,y,\theta):=x^{\top}\theta_{y}-\max_{y^{\prime}\in\mathcal{Y}\setminus\{y\}}x^{\top}\theta_{y^{\prime}}$, except for an additional factor of $2$ in one of the conditions (see Lemma A.10 in the Appendix).
 Hence, all the observations remain valid for the multi-class classification case.
-Discussion
-The conditions in Assumption 15 allow us to obtain convergence rate bounds for linearly separable data for some other predictors and loss functions.
-However, for linear classifiers, loss functions must satisfy certain properties related to hinge loss function.
-Further details are discussed in Appendix A.7.
-We next present a framework for deriving convergence rate bounds in more general cases.
-3.2 Bernoulli Sampling with Constant Step Size: Second Set of Convergence Conditions
+More General Classifiers and Datasets
 We consider algorithm (1) where $z_{t}$ is product of a fixed step size $\gamma$ and a Bernoulli random variable $\zeta_{t}$ with mean $\pi(x,y,\theta)$.
 Let $g_{t}=\zeta_{t}\nabla_{\theta}\ell(x_{t},y_{t},\theta_{t})$, which is random vector because $\zeta_{t}$ is a random variable and $(x_{t},y_{t})$ is a sampled point.
-Following Liu and Li (2023), we note that the algorithm (1) is a stochastic gradient descent algorithm with respect to an objective function $\tilde{\ell}$ with gradient
+Following Liu and Li (2023), we note that the algorithm (1) is an SGD algorithm with respect to an objective function $\tilde{\ell}$ with gradient
 $\nabla_{\theta}\tilde{\ell}(\theta)=\mathbb{E}[\pi(x,y,\theta)\nabla_{\theta}\ell(x,y,\theta)]$ (3)
 where the expectation is with respect to $x$ and $y$.
-This observation allows us to derive convergence rate results by deploying convergence rate results for stochastic gradient algorithm known to hold under various assumptions on function $\tilde{\ell}$, variance of stochastic gradient vector and step size.
-A function $\tilde{\ell}$ satisfying condition (7) is referred to as equivalent loss in Liu and Li (2023).
+This observation allows us to derive convergence rate results by deploying convergence rate results that are known to hold for SGD under various assumptions on function $\tilde{\ell}$, variance of stochastic gradient vector and step size.
+A function $\tilde{\ell}$ satisfying condition (3) is referred to as an equivalent loss in Liu and Li (2023).
 Assume that the sampling probability $\pi$ is an increasing function of the conditional expected loss $\ell(x,\theta)=\mathbb{E}_{y}[\ell(x,y,\theta)\mid x]$.
@@
 ~$\tilde{\ell}(\theta)=\mathbb{E}[\Pi(\ell(x,\theta))].$ (4)
-The function ~$\tilde{\ell}(\theta)$ inherits some properties of functions $\ell(x,y,\theta)$, for $(x,y)\in\mathcal{X}\times\mathcal{Y}$.
-For example, if $\ell(x,y,\theta)$ is a convex function, for every $(x,y)\in\mathcal{X}\times\mathcal{Y}$, then $\tilde{\ell}$ is a convex function.
-The framework for establishing convergence rates allows us to accommodate different sampling strategies and loss functions.
-For example, it accommodates binary classification for binary cross-entropy loss function and sampling proportional to absolute error loss, as shown in the following.
-For binary classification, we model the prediction probability of the positive label as $\sigma(x^{\top}\theta)$, where $\sigma:\mathbb{R}_{+}\rightarrow[0,1]$ is an increasing function, and $\sigma(-u)+\sigma(u)=1$ for all $u\in\mathbb{R}$.
-The absolute error loss takes value $1-\sigma(x^{\top}\theta)$ if $y=1$ or value $\sigma(x^{\top}\theta)$ if $y=-1$, which corresponds to $1-\sigma(yx^{\top}\theta)$.
-The binary cross-entropy loss for a point $(x,y)$ under model parameter $\theta$ can be written as $\ell(x,y,\theta)=-\log(\sigma(yx^{\top}\theta))$.
-Hence, absolute error loss-based sampling corresponds to the sampling probability function $\pi(x)=1-e^{-x}$.
-The next lemma allows to derive convergence rate results for expected loss with respect to loss function $\ell$ by applying convergence rate results for expected loss with respect to loss function $\tilde{\ell}$ (which, recall, is the equivalent loss function).
+If $\ell(x,y,\theta)$ is a convex function, for every $(x,y)\in\mathcal{X}\times\mathcal{Y}$, then $\tilde{\ell}$ is a convex function.
+This framework for establishing convergence rates allows us to accommodate different sampling strategies and loss functions.
+The next lemma allows us to derive convergence rate results for expected loss with respect to loss function $\ell$ by applying convergence rate results for expected loss with respect to loss function $\tilde{\ell}$ (which, recall, is the equivalent loss function).
 Lemma 3.2.
@@
 Theorem 3.3.
-Assume that $\ell$ is a convex function, $\tilde{\ell}$ is $L$-smooth, $\Theta_{0}$ is a convex set, $S=\sup_{\theta\in\Theta_{0}\cap\Theta}||\theta-\theta_{1}||$, and $\mathbb{E}[\pi(\ell(x,\theta))||\nabla_{\theta}\ell(x,y,\theta)||^{2}]-||\nabla_{\theta}\tilde{\ell}(\theta)||^{2}\leq\sigma_{\pi}^{2}$.
-Then, algorithm (1) with $\gamma=1/(L+(\sigma/R)\sqrt{n/2})$,
-$\mathbb{E}[\ell(\bar{\theta}_{n})]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]$
-$\leq$ $\inf_{\theta}\Pi^{-1}(\tilde{\ell}(\theta))+\Pi^{-1}\left(\frac{\sqrt{2}S\sigma_{\pi}}{\sqrt{n}}\right)+\Pi^{-1}\left(\frac{LS^{2}}{n}\right).$
+Assume that $\ell$ is a convex function, $\tilde{\ell}$ is $L$-smooth, $\Theta_{0}$ is a convex set, $S=\sup_{\theta\in\Theta_{0}}||\theta-\theta_{1}||$, and $\mathbb{E}[\pi(\ell(x,\theta))||\nabla_{\theta}\ell(x,y,\theta)||^{2}]-||\nabla_{\theta}\tilde{\ell}(\theta)||^{2}\leq\sigma_{\pi}^{2}$.
+Then, for algorithm (1) with $\gamma=1/(L+(\sigma/R)\sqrt{n/2})$,
+$\mathbb{E}[\ell(\bar{\theta}_{n})]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]\leq\inf_{\theta}\Pi^{-1}(\tilde{\ell}(\theta))+\Pi^{-1}\left(\frac{\sqrt{2}S\sigma_{\pi}}{\sqrt{n}}\right)+\Pi^{-1}\left(\frac{LS^{2}}{n}\right).$
 Note that the bound on the expected loss in Theorem 3.3 depends on $\pi$ through $\Pi^{-1}$ and $\sigma_{\pi}^{2}$.
-Obviously, we have a bound depending on $\pi$ only through $\Pi^{-1}$ by upper bounding $\sigma_{\pi}^{2}$ with $\sup_{\theta\in\Theta_{0}}\mathbb{E}[||\nabla_{\theta}\ell(x,y,\theta)||^{2}]$.
+Specifically, we have a bound depending on $\pi$ only through $\Pi^{-1}$ by upper bounding $\sigma_{\pi}^{2}$ with $\sup_{\theta\in\Theta_{0}}\mathbb{E}[||\nabla_{\theta}\ell(x,y,\theta)||^{2}]$.
 For convergence rates for large values of the number of iterations $n$, the bound in Theorem 3.2 crucially depends on how $\Pi^{-1}(x)$ behaves for small values of $x$.
@@
 Under assumptions of Theorem 3.3, sampling probability $\pi(x)=1-e^{-x}$, and $n\geq\max\left\{\left(\frac{16}{9}\right)^{2}2S\sigma_{\pi}^{2},\frac{16}{9}LS^{2}\right\}$ it holds
-$\mathbb{E}[\ell(\bar{\theta}_{n})]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]$
-$\leq$ $\inf_{\theta}\Pi^{-1}(\tilde{\ell}(\theta))+2^{5/4}\sqrt{S\sigma_{\pi}}\frac{1}{\sqrt[4]{n}}+2\sqrt{L}S\frac{1}{\sqrt{n}}.$
+$\mathbb{E}[\ell(\bar{\theta}_{n})]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(\theta_{t})\right]\leq\inf_{\theta}\Pi^{-1}(\tilde{\ell}(\theta))+2^{5/4}\sqrt{S\sigma_{\pi}}\frac{1}{\sqrt[4]{n}}+2\sqrt{L}S\frac{1}{\sqrt{n}}.$
 By using a bound on the expected total loss, we can bound the expected total number of sampled points under certain conditions as follows.
@@
 Hence, for all examples in Table 1, we have a bound for the expected number of sampled points provided we have a bound for the expected loss.
-3.3 Bernoulli Sampling with Stochastic Polyak’s Step Size: A New Algorithm
-In this section we propose a new algorithm that combines a Bernoulli sampling and an adaptive stochastic gradient descent update, and provide a convergence rate guarantee.
-The algorithm is defined as the stochastic gradient descent algorithm (1) with stochastic step size $z_{t}$ being a binary random variable that takes value $\gamma_{t}:=\zeta(x_{t},y_{t},\theta_{t})/\pi(x_{t},y_{t},\theta_{t})$ with probability $\pi(x_{t},y_{t},\theta_{t})$ and takes value 00 otherwise, where $\pi$ is some function and if $||\nabla_{\theta}\ell(x,y,\theta)||>0$,
+3.2 Adaptive-Weight Sampling
+In this section we propose the Adaptive-Weight Sampling (AWS) algorithm that combines Bernoulli sampling and an adaptive SGD update, and provide a convergence rate guarantee.
+The algorithm is defined by (1) with the stochastic step size $z_{t}$ being a binary random variable that takes value $\gamma_{t}:=\zeta(x_{t},y_{t},\theta_{t})/\pi(x_{t},y_{t},\theta_{t})$ with probability $\pi(x_{t},y_{t},\theta_{t})$ and takes value 00 otherwise, where $\pi$ is some sampling probability function.
+Here, $\zeta(x,y,\theta)$ is the expected SGD (1) step size, defined as
 $\zeta(x,y,\theta)=\beta\min\left\{\frac{1}{\psi(x,y,\theta)},\rho\right\}$
-and, $\zeta(x,y,\theta)=0$, otherwise, for constants $\beta,\rho>0$ and
+whenever $||\nabla_{\theta}\ell(x,y,\theta)||>0$ and $\zeta(x,y,\theta)=0$ otherwise, for constants $\beta,\rho>0$, where
 $\psi(x,y,\theta):=\frac{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}{\ell(x,y,\theta)-\inf_{\theta^{\prime}}\ell(x,y,\theta^{\prime})}.$
-The expected step size $\zeta(x,y,\theta)$ corresponds to the adaptive step size proposed by Polyak (Loizou et al., 2021).
-When $\pi(x,y,\theta)=1$ for all $x,y,\theta$, the algorithm reduces to Polyak’s adaptive step size algorithm (Loizou et al., 2021).
-Our algorithm introduces a sampling component and re-weighting of the update to ensure that the step size remains according to the stochastic Polyak’s step size in expectation.
-For many loss functions $\inf_{\theta^{\prime}}\ell(x,y,\theta)=0$, for every $x,y$.
+The expected step size $\zeta(x,y,\theta)$ corresponds to the stochastic Polyak’s step size used by a gradient descent algorithm proposed by (Loizou et al., 2021), which is accommodated as a special case when $\pi(x,y,\theta)=1$ for all $x,y,\theta$.
+AWS introduces a sampling component and re-weighting of the update to ensure that the step size remains according to the stochastic Polyak’s step size in expectation.
+For many loss functions, $\inf_{\theta^{\prime}}\ell(x,y,\theta)=0$, for every $x,y$.
 In these cases, $\psi(x,y,\theta)=||\nabla_{\theta}\ell(x,y,\theta)||^{2}/\ell(x,y,\theta)$.
 For instance, for binary cross-entropy loss function, $\inf_{\theta^{\prime}}\ell(x,y,\theta)=\inf_{\theta^{\prime}}(-\log(\sigma(yx^{\top}\theta^{\prime})))=0$, for all $x,y$.
-We next show a convergence rate guarantee.
-Lemma 3.10.
+We next show a convergence rate guarantee for AWS.
+Theorem 3.6.
 Assume that $\ell$ is a convex, $L$-smooth function, there exists $\Lambda^{*}$ such that $\mathbb{E}[\ell(x,y,\theta^{*})]-\mathbb{E}[\inf_{\theta}\ell(x,y,\theta)]\leq\Lambda^{*}$, and the sampling probability function $\pi$ is such that, for some constant $c\in(0,1)$, for all $x,y,\theta$ such that $||\nabla_{\theta}\ell(x,y,\theta)||>0$,
@@
 Then, we have
-$\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}(\ell(x_{t},y_{t},\theta_{t})-\ell(x_{t},y_{t},\theta^{*}_{t}))\right]$
-$\leq$ $\frac{\rho\beta}{c\kappa}\Lambda^{*}+\frac{1}{2c\kappa}||\theta_{1}-\theta^{*}||^{2}\frac{1}{n}$
+$\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}(\ell(x_{t},y_{t},\theta_{t})-\ell(x_{t},y_{t},\theta^{*}_{t}))\right]\leq\frac{\rho\beta}{c\kappa}\Lambda^{*}+\frac{1}{2c\kappa}||\theta_{1}-\theta^{*}||^{2}\frac{1}{n}$
 where $\kappa=\beta\min\{1/(2L),\rho\}$ and $\theta_{t}^{*}$ is a minimizer of $\ell(x_{t},y_{t},\theta^{\prime})$ over $\theta^{\prime}$.
-By taking $\rho=1/(2L)$, the bound on the expected average loss boils down to $\Lambda^{*}/c+(L/(c\beta))||\theta_{1}-\theta^{*}||^{2}/n$.
-Notably, under the condition on the sampling probability in Lemma 3.10, the convergence rate is of order $O(1/n)$.
-A similar bound is known to hold for stochastic gradient descent with adaptive stochastic Polyak step size for the finite-sum problem, as seen in Theorem 3.4 of Loizou et al.
+The bound on the expected average loss in Theorem 3.6 boils down to $\Lambda^{*}/c+(L/(c\beta))||\theta_{1}-\theta^{*}||^{2}/n$ by taking $\rho=1/(2L)$.
+Notably, under the condition on the sampling probability in Theorem 3.6, the convergence rate is of order $O(1/n)$.
+A similar bound is known to hold for SGD with adaptive stochastic Polyak step size for the finite-sum problem, as seen in Theorem 3.4 of Loizou et al.
 (2021).
-Lemma 3.10 applies to stream-based stochastic gradient descent algorithms and, importantly, allows for Bernoulli sampling with a sampling probability function $\pi$.
-We next show instances of $\pi$ that satisfy condition (10).
-Constant Sampling Probability
-Assume that $\kappa\leq 2(1-c)$ and sampling probability function $\pi$ is constant with value at least $\frac{\kappa}{2(1-c)}$. Then, the expected loss bound in Lemma 3.10 holds.
-Power Function of the Expected Step Size
-Consider $\pi$ defined as
-$\pi(x,y,\theta)=\zeta(x,y,\theta)^{\eta}$ (11)
-where $\eta\geq 0$ and $\rho\beta\in(0,1]$.
-This sampling probability function is an increasing function of $1/\psi(x,y,\theta)$, i.e.
-$(\ell(x,y,\theta)-\inf_{\theta^{\prime}}\ell(x,y,\theta))/||\nabla_{\theta}\ell(x,y,\theta)||^{2}$.
-In other words, the sampling probability is an increasing function of the loss normalized by the squared norm of the gradient.
-Assume that
-$\beta^{1-\eta}\leq 2(1-c)\left(\frac{1}{2L}\right)^{\eta}.$ (12)
-Then, the loss bound in Theorem 3.10 holds.
-Linear Classifiers: Binary Classification
-For linear classifiers, if $\ell(u)$ is an $L^{\prime}$-smooth function in $u$, then $\ell(yx^{\top}\theta)$ is an $L^{\prime}||x||^{2}$-smooth function in $\theta$.
-To satisfy condition (10), it suffices that
-$\pi(x,y,\theta)\geq\frac{\beta}{2(1-c)}\min\left\{\rho R^{2}h(yx^{\top}\theta),1\right\}$
-where $h(u):=\ell^{\prime}(u)^{2}/\ell(u)$ and note that $||\nabla_{\theta}\ell(x,y,\theta)||^{2}/\ell(x,y,\theta)=h(yx^{\top}\theta)||x||^{2}$.
-By setting $\rho=1/(2L)$ with $L=L^{\prime}R^{2}$, we get $\rho R^{2}=1/(2L^{\prime})$.
-For logistic regression and the binary cross-entropy loss function, the function $h(u)$ can be written as follows
-$h(u)=\frac{1}{(1+e^{u})^{2}\log(1+e^{-u})}$ (13)
-which is increasing in $u$ for $u\leq 0$ and decreasing in $u$ otherwise.
-Note that $h(u)=(1-e^{-\ell(u)})^{2}/\ell(u)$.
-Further properties of function $h$, deferred to Appendix A.13, establish two key facts: (a) Condition of Lemma 3.10 is satisfied with $\rho=1/(2L)$ by sampling proportional to absolute error loss $\pi^{*}(u)=\omega(1-\sigma(u))$ with $\beta/(4(1-c)L^{\prime})\leq\omega\leq 1$, and (b) Condition of Lemma 3.10 is satisfied by uncertainty sampling according to, for any fixed $a\in(0,1/2]$,
-$\pi^{*}(u)=\frac{\beta}{2(1-c)}\min\left\{\rho R^{2}\frac{1}{H(a)+(1-a)|u|},1\right\}$
-where $H(a)=a\log(1/a)+(1-a)\log(1/(1-a))$.
-Linear Classifiers: Multi-class Classification
-Condition (10) holds for some loss and uncertainty-based sampling methods. Refer to Appendix A.15 for detailed discussion.
+A difference is that Theorem 3.6 allows for sampling of the points.
+Loss and Uncertainty-based Sampling for Linear Binary Classifiers
+We consider linear binary classifiers, focusing particularly on logistic regression and the binary cross-entropy training loss function.
+The following corollaries of Theorem 3.6 hold for sampling proportional to absolute error loss and an uncertainty-based sampling probability function, respectively.
+Corollary 3.7.
+For sampling proportional to absolute error loss, $\pi(u)=\omega(1-\sigma(u))$, with $\beta/(4(1-c)L^{\prime})\leq\omega\leq 1$ and $\rho=1/(2L)$, the bound on the expected loss in Theorem 3.6 holds.
+Corollary 3.8.
+For the uncertainty-based sampling according to
+$\pi(u)=\frac{\beta}{2(1-c)}\min\left\{\rho R^{2}\frac{1}{H(a)+(1-a)|u|},1\right\}$
+where $a\in(0,1/2]$ and $H(a)=a\log(1/a)+(1-a)\log(1/(1-a))$, the bound on the expected loss in Theorem 3.6 holds.
+Other Cases
+For a constant sampling probability function with a value of at least $\kappa/(2(1-c))$, condition (6) holds when $\kappa\leq 2(1-c)$.
+When $\pi(x,y,\theta)=\zeta(x,y,\theta)^{\eta}$, where $\eta\geq 0$ and $\rho\beta\in(0,1]$, condition (6) holds under $\beta^{1-\eta}\leq 2(1-c)(1/(2L))^{\eta}$, as shown in Appendix A.14.
+Condition (6) also holds for an uncertainty-based sampling in multi-class classification, as shown in Appendix A.15.
 4 Numerical Results
-Figure 2: Average cross-entropy progressive loss of Polyak’s step size Loizou et al. (2021) compared to SGD with constant step size, for 1% and 10% sampling from the mushrooms data.
-Figure 3: Convergence in terms of average cross-entropy progressive loss of random sampling, loss-based sampling based on the absolute error loss, and our proposed adaptive step size variant of absolute error loss sampling.
-We adopt the setting from Loizou et al.
-(2021) and utilize the mushroom classification dataset (Chang and Lin, 2011) for binary classification tasks using RBF kernels and logistic regression.
-This dataset comprises $n=6499$ instances and $d=6499$ dimensions.
-In our training approach, we deliberately confine the training process to a single epoch.
+(a) Mushrooms dataset
+(b) MNIST classification of 3s vs 5s
+(c) Parkinsons dataset
+(d) Splice dataset
+(e) Tictactoe dataset
+(f) Credit dataset
+Figure 1: Convergence in terms of average cross-entropy progressive loss of random sampling, loss-based sampling based on the absolute error loss, and our proposed algorithm (loss-based sampling with stochastic Polyak’s step size).
+Our proposed algorithm outperforms the baselines in most cases.
+In this section we evaluate our AWS algorithm, defined in Section 3.2.
+In particular, we focus on an instance of AWS with stochastic Polyak’s expected step size for logistic regression and the loss-based sampling proportional to absolute error loss, which we refer to as Adaptive-Weight Sampling - Polyak Absloss (AWS-PA).
+By, Corollary 3.7, AWS-PA converges according to Theorem 3.6.
+Here we demonstrate convergence on real-world datasets and compare with other algorithms.
+We use a modified version of the mushroom binary classification dataset (Chang and Lin, 2011) that was used by Loizou et al.
+(2021) for evaluation of their algorithm.
+This modification uses RBF kernel features, resulting in a linearly separable dataset for a linear classifier like logistic regression.
+Furthermore, we include five datasets that we selected at random from the 44 real-world datasets that were used in Yang and Loog (2018), a benchmark study of active learning for logistic regression: MNIST 3 vs 5 LeCun et al.
+(1998), parkinsons Little et al.
+(2007), splice Noordewier et al.
+(1990), tictactoe Aha (1991), and credit Quinlan (1987).
+While these datasets are not necessarily linearly separable, Yang and Loog (2018) has shown that logistic regression achieves a good quality-of-fit.
+In our evaluation, we deliberately confine the training to a single epoch.
 Throughout this epoch, we sequentially process each data instance, compute the loss for each individual instance, and subsequently update the model’s weights.
@@
 The constraint to a single epoch ensures that we calculate losses only for instances that haven’t influenced model weights.
-For each sampling scheme, we conduct a hyper-parameter sweep to minimize the average progressive loss.
-Loizou et al.
-(2021) demonstrated that stochastic gradient descent with a step size corresponding to their stochastic Polyak’s step size (SPS) converges faster than gradient descent.
-Figure 2 illustrates that these findings extend to scenarios where we selectively sample from the dataset rather than training on the full dataset, and the step size is according to stochastic Polyak’s step size only in expectation.
-[^1] We will make the code and the dataset available for download in the final version of the manuscript.
-Figure 3 demonstrates that our proposed active learning algorithm, with adaptive Polyak step size according to SPS and sampling according to $\pi(x)=1-e^{-x}$, leads to faster convergence than the traditional loss-based sampling.
-The figure also shows that the traditional loss-based sampling approach converges more rapidly than random sampling.
-Figure 4: Robustness of the proposed sampling approach with adaptive Polyak’s step size for different variance $\mathrm{var}[\hat{\ell}_{\mathrm{abs}}]=\ell^{2}_{\mathrm{abs}}(1-\ell_{\mathrm{abs}})/(\alpha+\ell_{\mathrm{abs}})$ noise levels of absolute error loss estimator: (low) $\alpha=100$, (medium) $\alpha=2.5$, and (high) $\alpha=1$.
-We also simulate a noisy estimator of the absolute error loss for the proposed loss-based sampling with adaptive Polyak’s step size.
-This corresponds to the loss-based active learning setting in practice, where the loss is not known prior to label acquisition and can only be estimated.
-We model an unbiased noisy estimator ^$\hat{\ell}_{\mathrm{abs}}$ of the absolute error loss $\ell_{\mathrm{abs}}\in[0,1]$ as a random variable following the beta distribution, denoted as ^$\hat{\ell}_{\mathrm{abs}}\sim\mathrm{Beta}(\alpha,\beta)$, where $\alpha$ and $\beta$ are parameters set to ensure $\mathbb{E}[\hat{\ell}_{\mathrm{abs}}]=\ell_{\mathrm{abs}}$.
-The variance of the noise can be controlled by the tuning parameter $\alpha$, given by $\mathrm{var}[\hat{\ell}_{\mathrm{abs}}]=\frac{\ell_{\mathrm{abs}}(1-\ell_{\mathrm{abs}})}{\alpha+\ell_{\mathrm{abs}}}$.
-Figure 4).
-Figure 4 shows that the convergence results remain robust against noise in estimating the absolute error loss for a wide range of values for $\alpha\geq 1$.
+For each sampling scheme, we conduct a hyper-parameter sweep to minimize the average progressive loss and apply a procedure to ensure that all algorithms sample comparable numbers of instances.
+In Appendix B.1 we include further details on this procedure, the hyper-parameter tuning, and other aspects of the experimental setup.
+Figure 1 demonstrates that AWS-PA leads to faster convergence than the traditional loss-based sampling with a constant step size (akin to Yoo and Kweon (2019)).
+It also shows that the traditional loss-based sampling approach converges more rapidly than random sampling on five of the six datasets.
+These results are obtained under a hyper-parameter tuning such that different algorithms have comparable data sampling rates.
+We provide additional experimental results that demonstrate the efficiency of AWS-PA in Appendix B.2.
+(a) Mushrooms dataset
+(b) MNIST classification of 3s vs 5s
+(c) Parkinsons dataset
+(d) Splice dataset
+(e) Tictactoe dataset
+(f) Credit dataset
+Figure 2: Active learning sampling based on an estimator of the absolute error loss performs on par with the sampling based on the ground truth value of absolute error loss.
+In active learning applications, the true loss of a point cannot be computed before the corresponding label is obtained.
+Hence, in practice we do not know the true loss at the moment of making the sampling decision.
+Therefore, we assess the effect of using a loss estimator, instead of using the true loss values.
+We use a Random Forest regressor to estimate absolute error loss based on the same set of features as the target model and the target’s model prediction as an extra feature.
+We retrain this estimator on every sampling step using the labeled points observed so far.
+Figure 2 demonstrates that AWS-PA with the estimated absolute error losses performs similarly on all datasets to AWS-PA with the true absolute error losses.
+Moreover, for a majority of the datasets, the two variants of AWS-PA achieve similar data sampling rates; this is shown Appendix B.3 along with further discussion.
 5 Conclusion
 In this paper, we have provided convergence rate guarantees for both loss and uncertainty-based active learning algorithms under various assumptions.
-Furthermore, we introduced a novel algorithm that combines sampling with an adaptive size, conforming to stochastic Polyak’s step size in expectation, and demonstrated its convergence rate guarantee, contingent on a condition related to the sampling probability function.
+Furthermore, we introduced the novel Adaptive-Weight Sampling (AWS) algorithm that combines sampling with an adaptive size, conforming to stochastic Polyak’s step size in expectation, and demonstrated its convergence rate guarantee, contingent on a condition related to the sampling probability function.
 For future research, it would be interesting to apply the methodologies presented in this paper to obtain additional convergence rate results.
-Additionally, exploring a theoretical study on the impact of noise in the loss estimator, used for evaluating the sampling probability function, on the convergence properties of algorithms could open up a valuable avenue for investigation.
-6 Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A MISSING PROOFS AND ADDITIONAL RESULTS
-A.1 A Lemma used for our First Set of Convergence Conditions
-We present a lemma that is used to establish all convergence rate results for the expected loss of algorithm (1) presented in Section 3.1.
-The lemma relies on the following condition involving the loss function $\ell$ used by algorithm (1) and the loss function $\tilde{\ell}$ used for evaluating the performance of the algorithm.
+Additionally, exploring a theoretical study on the impact of bias and noise in the loss estimator, used for evaluating the sampling probability function, on the convergence properties of algorithms could open up a valuable avenue for investigation.
+Appendix A Appendix: Proofs, Discussion, and Additional Results
+A.1 Limitations & Discussion
+It remains an open problem to investigate the tightness of convergence rate bounds for constant-weight sampling under the assumptions outlined in Theorem 3.3.
+The convergence rate results presented in Theorems 3.3 and 3.6 pertain to smooth convex training loss functions. Future research may explore weaker assumptions regarding the training loss function.
+Regarding loss-based sampling strategies, our theoretical analysis assumes an unbiased and noiseless loss estimator.
+Extending this to account for estimation bias and noise would be a valuable avenue for further investigation.
+Our numerical results demonstrate the effectiveness of our proposed algorithm and the robustness of our theoretical findings to loss estimation bias and noise across different datasets, utilizing the logistic regression model as a binary classifier.
+Future work could explore the application of other classification models, such as multi-layer neural networks.
+A.2 A Set of Convergence Rate Conditions
+We present and prove two convergence rate lemmas for algorithm (1): the first providing sufficient conditions for a certain convergence rate and the second restricted to linear classifiers and linearly separable datasets.
+The first lemma relies on the following condition involving the loss function $\ell$ used by algorithm (1) and the loss function $\tilde{\ell}$ used for evaluating the performance of the algorithm.
 Assumption A.1.
@@
 $\mathbb{E}[\tilde{\ell}(x,y,\bar{\theta}_{n})]\leq||\theta_{1}-\theta^{*}||^{2}\frac{\alpha}{\beta^{2}}\frac{1}{n}$
-where $\bar{\theta}_{t}=(1/t)\sum_{s=1}^{t}\theta_{s}$.
+where $\bar{\theta}_{n}=(1/n)\sum_{t=1}^{n}\theta_{t}$.
 Proof.
@@
 The second statement of the lemma follows from the last above inequality and Jensen’s inequality. ∎
-A.2 Proof of Theorem 3.2
+For the case of linear classifiers and linearly separable datasets, it can be readily checked that Lemma A.2 implies the following lemma.
+Lemma A.3.
+Assume that there exist constants $\alpha,\beta>0$ such that for all $u\in\mathbb{R}$,
+$\pi(u)\ell^{\prime}(u)^{2}R^{2}\leq\alpha\tilde{\ell}(u)$ (9)
+and
+$\pi(u)(-\ell^{\prime}(u))(\rho^{*}-u)\geq\beta\tilde{\ell}(u).$ (10)
+Then, for any $\theta_{1}$ such that $||\theta_{1}-\theta^{*}||\leq S$ and $\{\theta_{t}\}_{t>1}$ according to algorithm (1) with $\gamma=\beta/\alpha$,
+$\mathbb{E}\left[\sum_{t=1}^{n}\tilde{\ell}(y_{t}x_{t}^{\top}\theta_{t})\right]\leq S^{2}\frac{\alpha}{\beta^{2}}.$
+Moreover, if $\tilde{\ell}$ is a convex function, then $\mathbb{E}\left[\tilde{\ell}(yx^{\top}\bar{\theta}_{n})\right]\leq S^{2}\alpha/(\beta^{2}n)$ where $\bar{\theta}_{n}=(1/n)\sum_{t=1}^{n}\theta_{t}$.
+A.2.1 Discussion
+We discuss some implications of conditions (9) and (10). Some of this discussion will help us identify the types of loss functions to which the conditions cannot be applied.
+First, we note that ~$\tilde{\ell}(u)>0$ implies $\pi(u)>0$. Hence, equivalently, $\pi(u)=0$ implies ~$\tilde{\ell}(u)=0$.
+Second, we note that under conditions (9) and (10) it is necessary that for all $u\in\mathbb{R}$, either
+$-\ell^{\prime}(u)\leq\frac{\alpha}{\beta R^{2}}\max\{\rho^{*}-u,0\}\hbox{ or }\pi(u)=0.$
+Hence, whenever ~$\tilde{\ell}(u)>0$ (and thus $\pi(u)>0$), then $-\ell^{\prime}(u)\leq\frac{\alpha}{\beta R^{2}}\max\{\rho^{*}-u,0\}$.
+The latter condition means that $\ell^{\prime}(u)=0$ whenever $u\geq\rho^{*}$ and otherwise the derivative of $\ell$ at $u$ is bounded such that $\ell^{\prime}(u)\geq-(\alpha/(\beta R^{2}))(\rho^{*}-u)$.
+In other words, function $\ell$ must not decrease too fast on $(-\infty,\rho^{*}]$.
+Third, assume $\ell$ and $\tilde{\ell}$ are such that $\ell(u)=0$ and ~$\tilde{\ell}(u)=0$ for all $u\geq 1$ and ~$\tilde{\ell}(u)>0$ for all $u<1$. Then, for every $u\leq 1$
+[…]
+which by integrating is equivalent to
+$\ell(u)\leq\frac{\alpha}{\beta R^{2}}\left((\rho^{*}-1)(1-u)+\frac{1}{2}(1-u)^{2}\right).$
+This shows that $\ell$ must be upper bounded by a linear combination of hinge and squared hinge loss function.
+Forth, assume that $\pi$ is decreasing in $u$, then for every fixed $u_{0}\in\mathbb{R}$,
+~$\tilde{\ell}(u)\geq\frac{\pi(u_{0})R^{2}}{\alpha}(-\ell^{\prime}(u))^{2}\hbox{for every }u\leq u_{0}.$
+If $\tilde{\ell}=\ell$, then
+$\ell(u)\leq\left(\sqrt{\ell(u_{0})}+\frac{1}{2R}\sqrt{\frac{\alpha}{\pi(u_{0})}}(u_{0}-u)\right)^{2}\hbox{ for all }u\leq u_{0}.$
+Fifth, and last, assume that $\pi$ is an even function and that there exists $c>0$ and $u_{0}\leq\rho^{*}$ such that ~$\tilde{\ell}(u)\geq c$ for every $u\leq u_{0}$. Then,
+$\pi(u)=\Omega\left(\frac{1}{|u|^{2}}\right)$
+which limits the rate at which $\pi(u)$ is allowed to decrease with $|u|$. To see, this, from conditions (9) and (10), for every $u\leq\rho^{*}$,
+~$\tilde{\ell}(u)\leq\frac{\alpha}{\beta^{2}R^{2}}\pi(u)(\rho^{*}-u)^{2}.$
+Hence, $\pi(u)\geq(c\beta^{2}R^{2}/\alpha)/(\rho^{*}-u))$ for every $u\leq u_{0}$.
+The lower bound is tight in case when $\ell$ is squared hinge loss function and ~$\tilde{\ell}(u)=c$ for every $u\leq u_{0}<1$.
+In this case, from conditions (9) and (10), for every $u\leq u_{0}$,
+$c\beta\frac{1}{(1-u)(\rho^{*}-u)}\leq\pi(u)\leq\frac{c\alpha}{R^{2}}\frac{1}{(1-u)^{2}}$
+which implies $\pi(u)=\Theta(1/|u|^{2})$.
+A.3 Proof of Theorem 3.1
+We show that conditions of the theorem imply conditions (9) and (10) to hold, for $\tilde{\ell}=\ell$, which in turn imply conditions (7) and (8) and hence we can apply the convergence result of Lemma A.3.
+We first consider condition (7).
+For squared hinge loss function $\ell$ and $\tilde{\ell}=\ell$, clearly condition (7) holds for every $u\geq 1$ as in this case both side of the inequality are equal to zero.
+For every $u\leq 1$,
+$\frac{\ell(u)}{\ell^{\prime}(u)^{2}}=\frac{1}{2}.$
+Since by assumption $\pi(u)\leq\beta/2$ for all $u$, condition $\alpha/\beta\geq R^{2}$ implies condition (7).
+We next consider condition (8). Again, clearly, condition holds for every $u\geq 1$ as in this case both sides of the inequality are equal to zero. For $u\leq 1$, we can write (8) as follows
+$\pi(u)\geq\beta\frac{\ell(u)}{(-\ell^{\prime}(u))(\rho^{*}-u)}$ $=$ $\frac{\beta}{2}\frac{1-u}{\rho^{*}-u}$
+$=$ $\frac{\beta}{2}\left(1-\frac{1}{1+(1/(\rho^{*}-1))(1-u)}\right)$
+$=$ $\frac{\beta}{2}\left(1-\frac{1}{1+(\sqrt{2}/(\rho^{*}-1))\sqrt{\ell(u)}}\right).$
+This condition is implied by $\pi(u)\geq\pi^{*}(\ell(u))$ where
+$\pi^{*}(\ell)=\frac{\beta}{2}\left(1-\frac{1}{1+\mu\sqrt{\ell}}\right)$
+with $\mu\geq\sqrt{2}/(\rho^{*}-1)$.
+The result of the theorem follows from Lemma A.3 with $\alpha/\beta=R^{2}$, $0<\beta\leq 2$ and $\pi(u)\geq\pi^{*}(\ell(u))$ for all $u\leq 1$.
+For the expected number of samples we proceed as follows. First by concavity and monotonicity of function $\pi^{*}$ and the expected loss bound, we have
+$\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(\ell(x_{t},y_{t},\theta_{t}))\right]\leq\pi^{*}\left(\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(x_{t},y_{t},\theta_{t})\right]\right)n\leq\pi^{*}\left(\frac{||\theta_{1}-\theta^{*}||^{2}R^{2}}{\beta n}\right)n.$
+Then, combined with the fact $\pi^{*}(v)\leq\frac{\beta\mu}{2}\sqrt{v}$ for all $v\geq 0$, it follows
+$\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(\ell(x_{t},y_{t},\theta_{t}))\right]\leq||\theta_{1}-\theta^{*}||\frac{\mu\sqrt{\beta}}{2}\sqrt{n}.$
+Since $\pi^{*}(v)\leq\beta/2$ for all $v$, it obviously holds $\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(\ell(x_{t},y_{t},\theta_{t}))\right]\leq(\beta/2)n$, which completes the proof of the theorem.
+A.4 Linear Classifiers: Zero-one Loss and Absolute Error Loss-based Sampling
+Here we consider training loss functions satisfying:
+Assumption A.4.
+Function $\ell$ is continuously differentiable on $(-\infty,0]$, convex, and $\ell^{\prime}(0)\leq-c_{1}$ and $\lim_{u\rightarrow-\infty}\ell^{\prime}(u)\geq-c_{2}$, for some constants $c_{1},c_{2}>0$.
+We consider sampling proportional to either zero-one loss or absolute error loss.
+The zero-one loss is defined as $\ell_{01}(u):=\mathbbm{1}_{\{u\leq 0\}}=\mathbbm{1}_{\{y\neq\mathrm{sgn}(x^{\top}\theta)\}}$.
+The absolute error loss is defined as $\ell_{\mathrm{abs}}(u):=2\left(\mathbbm{1}_{\{u<0\}}+\frac{1}{2}\mathbbm{1}_{\{u=0\}}\right)=|y-\mathrm{sgn}(x^{\top}\theta)|$.
+Sampling proportional to zero-one loss is defined as $\pi(u)=\omega\ell_{01}(u)$ for some constant $\omega\in(0,1]$, while sampling proportional to absolute error loss is defined as $\pi(u)=\omega\ell_{\mathrm{abs}}(u)$ for some constant $\omega\in(0,1/2]$.
+Theorem A.5.
+Assume that the loss function $\ell$ satisfies Assumption A.4 and $\rho^{*}>0$.
+Then, under sampling proportional to zero-one loss, for any initial value $\theta_{1}$ such that $||\theta_{1}-\theta^{*}||\leq S$ and $\{\theta_{t}\}_{t>1}$ according to algorithm (1) with $\gamma=c_{1}\rho^{*}/(c_{2}^{2}R^{2})$,
+$\mathbb{E}\left[\sum_{t=1}^{n}\ell_{01}(y_{t}x_{t}^{\top}\theta_{t})\right]\leq\frac{c_{2}^{2}R^{2}S^{2}}{c_{1}^{2}\omega}\frac{1}{{\rho^{*}}^{2}}.$ (11)
+Furthermore, under sampling proportional to absolute error loss and $\gamma=2c_{1}\rho^{*}/(c_{2}^{2}R^{2})$, the bound in (11) holds but with an additional factor of $2$.
+Proof is provided in Appendix A.4.1.
+The bound in (11) is a well-known bound on the number of mistakes made by the perceptron algorithm, of the order $O(1/{\rho^{*}}^{2})$.
+It can be readily observed that under sampling proportional to zero-one loss, the expected number of sampled points is bounded by $(c_{2}/c_{1})^{2}R^{2}S^{2}/{\rho^{*}}^{2}$, which also holds for sampling proportional to absolute error loss but with an additional factor of $4$.
+A.4.1 Proof of Theorem A.5
 We first consider the case when sampling is proportional to zero-one loss.
@@
 Hence, conditions (9) and (10) hold by taking $\alpha=2\omega c_{2}^{2}R^{2}$ and $\beta=\omega c_{1}\rho^{*}$.
-A.3 Proof of Theorem 3.3
-We show that conditions of the theorem imply conditions (2) and (3) to hold, for $\tilde{\ell}=\ell$, which in turn imply conditions (14) and (15) and hence we can apply the convergence result of Lemma A.2.
-We first consider condition (14).
-For squared hinge loss function $\ell$ and $\tilde{\ell}=\ell$, clearly condition (14) holds for every $u\geq 1$ as in this case both side of the inequality are equal to zero.
-For every $u\leq 1$,
-$\frac{\ell(u)}{\ell^{\prime}(u)^{2}}=\frac{1}{2}.$
-Since by assumption $\pi(u)\leq\beta/2$ for all $u$, condition $\alpha/\beta\geq R^{2}$ implies condition (14).
-We next consider condition (15). Again, clearly, condition holds for every $u\geq 1$ as in this case both sides of the inequality are equal to zero. For $u\leq 1$, we can write (15) as follows
-$\pi(u)\geq\beta\frac{\ell(u)}{(-\ell^{\prime}(u))(\rho^{*}-u)}$ $=$ $\frac{\beta}{2}\frac{1-u}{\rho^{*}-u}$
-$=$ $\frac{\beta}{2}\left(1-\frac{1}{1+(1/(\rho^{*}-1))(1-u)}\right)$
-$=$ $\frac{\beta}{2}\left(1-\frac{1}{1+(\sqrt{2}/(\rho^{*}-1))\sqrt{\ell(u)}}\right).$
-This condition is implied by $\pi(u)\geq\pi^{*}(\ell(u))$ where
-$\pi^{*}(\ell)=\frac{\beta}{2}\left(1-\frac{1}{1+\mu\sqrt{\ell}}\right)$
-with $\mu\geq\sqrt{2}/(\rho^{*}-1)$.
-The result of the theorem follows from Lemma A.2 with $\alpha/\beta=R^{2}$, $0<\beta\leq 2$ and $\pi(u)\geq\pi^{*}(\ell(u))$ for all $u\leq 1$.
-For the expected number of samples we proceed as follows. First by concavity and monotonicity of function $\pi^{*}$ and the expected loss bound, we have
-$\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(\ell(x_{t},y_{t},\theta_{t}))\right]\leq\pi^{*}\left(\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(x_{t},y_{t},\theta_{t})\right]\right)n\leq\pi^{*}\left(\frac{||\theta_{1}-\theta^{*}||^{2}R^{2}}{\beta n}\right)n.$
-Then, combined with the fact $\pi^{*}(v)\leq\frac{\beta\mu}{2}\sqrt{v}$ for all $v\geq 0$, it follows
-$\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(\ell(x_{t},y_{t},\theta_{t}))\right]\leq||\theta_{1}-\theta^{*}||\frac{\mu\sqrt{\beta}}{2}\sqrt{n}.$
-Since $\pi^{*}(v)\leq\beta/2$ for all $v$, it obviously holds $\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(\ell(x_{t},y_{t},\theta_{t}))\right]\leq(\beta/2)n$, which completes the proof of the theorem.
-A.4 Proof of Theorem 3.4
+A.5 Linear Classifiers: Generalized Smooth Hinge Loss Function
+Here we consider the training loss function corresponding to the generalized smooth hinge loss function [Rennie, 2005], defined for $a\geq 1$ as follows:
+$\ell(u)=\left\{\begin{array}[]{ll}\frac{a}{a+1}-u&\hbox{ if }u\leq 0\\ \frac{a}{a+1}-u+\frac{1}{a+1}u^{a+1}&\hbox{ if }0\leq u\leq 1\\ 0&\hbox{ otherwise}.\end{array}\right.$ (12)
+This is a continuously differentiable function that converges to the value of the hinge loss function, $\max\{1-u,0\}$, as $a$ goes to infinity.
+The family of loss functions parameterized by $a$ accommodates the smooth hinge loss function with $a=1$, introduced by Rennie and Srebro [2005].
+Theorem A.6.
+Assume that $\rho^{*}>1$, the loss function is the generalized smooth hinge loss function, and the sampling probability is according to the function $\pi^{*}$, which, for $\beta\in(0,1]$ and $\rho\in(1,\rho^{*}]$, is defined as
+$\pi^{*}(u)=\left\{\begin{array}[]{ll}\beta\frac{\frac{a}{a+1}-u}{\rho-u}&\hbox{ if }u\leq 0\\ \beta\frac{\frac{a}{a+1}\frac{1}{1-u^{a}}(1-u)-\frac{1}{a+1}u}{\rho-u}&\hbox{ if }0\leq u\leq 1\\ 0&\hbox{ if }u\geq 1.\end{array}\right.$
+Then, for any initial value $\theta_{1}$ such that $||\theta_{1}-\theta^{*}||\leq S$ and $\{\theta_{t}\}_{t>1}$ according to algorithm (1) with $\gamma=1/(c_{a,\rho}R^{2})$, where $c_{a,\rho}=a/(a(\rho-1)+1)$, we have
+$\mathbb{E}\left[\ell(yx^{\top}\bar{\theta}_{n})\right]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(y_{t}x_{t}^{\top}\theta_{t})\right]\leq c_{a,\rho}\frac{R^{2}S^{2}}{\beta}\frac{1}{n}.$
+Furthermore, the same bound holds for every $\pi$ such that $\pi^{*}(u)\leq\pi(u)\leq\beta$ for all $u\in\mathbb{R}$, with $c_{a,\rho}=2a$.
+Proof is provided in Appendix A.5.1.
+Note that $\pi^{*}$ is increasing in $a$ and is upper-bounded by $\pi^{**}(u)=\beta(1-u)/(\rho-1)$, which may be regarded as the limit for the hinge loss function. See Figure 3 for an illustration.
+We remark that the expected loss bound in Theorem A.6 is the same as for the squared hinge loss function in Theorem 3.1, except for an additional factor $c_{a,\rho}=a/(a(\rho-1)+1)$.
+This factor is increasing in $a$ but always lies in the interval $[1/\rho,1/(\rho-1)]$, where the boundary values are achieved for $a=1$ and $a\rightarrow\infty$, respectively.
+Figure 3: Sampling probability function for the family of generalized smooth hinge loss functions.
+Theorem A.7.
+Assuming $\rho^{*}>1$, the loss function is the generalized smooth hinge loss, and the sampling probability function $\pi^{*}$ satisfies the assumptions in Theorem A.6, with $n>((a+1)/a)c_{a,\rho}R^{2}S^{2}/\beta$, the expected number of sampled points is bounded as
+$\mathbb{E}\left[\sum_{t=1}^{n}\pi^{*}(y_{t}x_{t}^{\top}\theta_{t})\right]\leq\kappa\max\left\{\frac{\sqrt{\beta c_{a,\rho}}RS}{(\rho-1)\sqrt{a}}\sqrt{n},\frac{c_{a,\rho}R^{2}S^{2}}{\rho-1}\right\},$
+where $\kappa$ is some positive constant.
+Proof is provided in Appendix A.5.2.
+For any fixed number of iterations $n$ satisfying the condition of the theorem, the expected number of sampled points is bounded by a constant for sufficiently large value of parameter $a$.
+The bound in Theorem A.7 depends on how the loss function $\ell(u)$ varies with $u$.
+When $a$ is large, $\ell(u)$ is approximately $1-u$ (hinge loss), otherwise, it is approximately $\frac{a}{2}(1-u)^{2}$ for $0\leq u\leq 1$ (squared hinge loss).
+A.5.1 Proof of Theorem A.6
 Assume that $\pi^{*}$ is such that for given $\rho\in(1,\rho^{*}]$, $\pi^{*}(u)(-\ell^{\prime}(u))(\rho-u)=\beta\ell(u)$ for all $u\in\mathbb{R}$.
@@
 It follows that $\inf_{u\leq 1}f(u)=f(1)=\frac{1}{2a}\leq\alpha/(\beta R^{2})$, hence it is suffices that $\alpha/\beta\geq R^{2}/(2a)$.
-A.5 Proof of Theorem 3.5
+A.5.2 Proof of Theorem A.7
 Lemma A.8.
@@
 $\pi^{*}(u)\leq\pi^{**}(u)=\beta\frac{1-u}{\rho-1}\hbox{ for every }u\leq 1.$
-Proof
+Proof.
 For $u\leq 1$, $\pi^{*}(u)=\beta(a/(a+1)-u)/(\rho-u)$, so obviously, $\pi^{*}(u)\leq\beta(1-u)/(\rho-u)$.
@@
 To show the asserted inequality, by straightforward calculus it can be shown that the inequality is equivalent to $(1-(1-u))^{1-a}\geq 1-(1-a)(1-u)$ which clearly holds true.
+∎
 Lemma A.9.
@@
 $\ell(u)\geq\left(1-\frac{1}{3}\frac{c}{1-c/2}\right)\frac{a}{2}(1-u)^{2}.$
-Proof
+Proof.
 We first show the first inequality. For $u\leq 1$, it clearly holds $\ell(u)\leq 1-u$, and
@@
 $\leq$ $\frac{\beta}{\rho-1}\left(1-\ell^{-1}\left(c_{a,\rho}\frac{||\theta_{1}-\theta^{*}||^{2}R^{2}}{\beta}\frac{1}{n}\right)\right)n$
-where the first inequality is by concavity of the function $1-\ell^{-1}(\ell)$ and the second inequality is by Theorem 3.4.
+where the first inequality is by concavity of the function $1-\ell^{-1}(\ell)$ and the second inequality is by Theorem A.6.
 To apply Lemma A.9, we need that
@@
 $\ell(x,y,\theta)\equiv\ell(u(x,y,\theta))$ and ~$\tilde{\ell}(x,y,\theta)\equiv\tilde{\ell}(u(x,y,\theta))$.
-For example, this accomodates hinge loss function for multi-class classification Crammer and Singer (2002).
+For example, this accomodates hinge loss function for multi-class classification Crammer and Singer [2002].
 Lemma A.10.
@@
 $\pi(x,y,\theta)(-\ell^{\prime}(u(x,y,\theta)))(\rho^{*}-u(x,y,\theta))\geq\beta\tilde{\ell}(u(x,y,\theta)).$
-Proof of the lemma
+Proof.
 For $x\in\mathcal{X}$ and $y\in\mathcal{Y}$, let $\phi(x,y)=(\phi_{1}(x,y)^{\top},\ldots,\phi_{k}(x,y)^{\top})^{\top}$ where $\phi_{i}(x,y)=x$ if $i=y$ and $\phi_{i}(x,y)$ is the $d$-dimensional null-vector, otherwise.
@@
 $\pi(x,y,\theta)(-\ell^{\prime}(u(x,y,\theta))(\rho^{*}-u(x,y,\theta))\geq\beta\tilde{\ell}(u(x,y,\theta)).$
-Note that these conditions are equivalent to those for the binary case in (2) and (3) except for an additional factor $2$ in the first of the last above inequalities.
-A.7 Implications of Conditions (2) and (3)
-We discuss some implications of conditions (2) and (3). Some of this discussion will tell us about loss functions for which conditions cannot be applied.
-First, we note that ~$\tilde{\ell}(u)>0$ implies $\pi(u)>0$. Hence, equivalently, $\pi(u)=0$ implies ~$\tilde{\ell}(u)=0$.
-Second, we note that under conditions (2) and (3) it is necessary that for all $u\in\mathbb{R}$, either
-$-\ell^{\prime}(u)\leq\frac{\alpha}{\beta R^{2}}\max\{\rho^{*}-u,0\}\hbox{ or }\pi(u)=0.$
-Hence, whenever ~$\tilde{\ell}(u)>0$ (and thus $\pi(u)>0$), then $-\ell^{\prime}(u)\leq\frac{\alpha}{\beta R^{2}}\max\{\rho^{*}-u,0\}$.
-The latter condition means that $\ell^{\prime}(u)=0$ whenever $u\geq\rho^{*}$ and otherwise the derivative of $\ell$ at $u$ is bounded such that $\ell^{\prime}(u)\geq-(\alpha/(\beta R^{2}))(\rho^{*}-u)$.
-In other words, function $\ell$ must not decrease too fast on $(-\infty,\rho^{*}]$.
-Third, assume $\ell$ and $\tilde{\ell}$ are such that $\ell(u)=0$ and ~$\tilde{\ell}(u)=0$ for all $u\geq 1$ and ~$\tilde{\ell}(u)>0$ for all $u<1$. Then, for every $u\leq 1$
-[…]
-which by integrating is equivalent to
-$\ell(u)\leq\frac{\alpha}{\beta R^{2}}\left((\rho^{*}-1)(1-u)+\frac{1}{2}(1-u)^{2}\right).$
-This shows that $\ell$ must be upper bounded by a linear combination of hinge and squared hinge loss function.
-Forth, assume that $\pi$ is decreasing in $u$, then for every fixed $u_{0}\in\mathbb{R}$,
-~$\tilde{\ell}(u)\geq\frac{\pi(u_{0})R^{2}}{\alpha}(-\ell^{\prime}(u))^{2}\hbox{for every }u\leq u_{0}.$
-If $\tilde{\ell}=\ell$, then
-$\ell(u)\leq\left(\sqrt{\ell(u_{0})}+\frac{1}{2R}\sqrt{\frac{\alpha}{\pi(u_{0})}}(u_{0}-u)\right)^{2}\hbox{ for all }u\leq u_{0}.$
-Fifth, and last, assume that $\pi$ is an even function and that there exists $c>0$ and $u_{0}\leq\rho^{*}$ such that ~$\tilde{\ell}(u)\geq c$ for every $u\leq u_{0}$. Then,
-$\pi(u)=\Omega\left(\frac{1}{|u|^{2}}\right)$
-which limits the rate at which $\pi(u)$ is allowed to decrease with $|u|$. To see, this, from conditions (2) and (3), for every $u\leq\rho^{*}$,
-~$\tilde{\ell}(u)\leq\frac{\alpha}{\beta^{2}R^{2}}\pi(u)(\rho^{*}-u)^{2}.$
-Hence, $\pi(u)\geq(c\beta^{2}R^{2}/\alpha)/(\rho^{*}-u))$ for every $u\leq u_{0}$.
-The lower bound is tight in case when $\ell$ is squared hinge loss function and ~$\tilde{\ell}(u)=c$ for every $u\leq u_{0}<1$.
-In this case, from conditions (2) and (3), for every $u\leq u_{0}$,
-$c\beta\frac{1}{(1-u)(\rho^{*}-u)}\leq\pi(u)\leq\frac{c\alpha}{R^{2}}\frac{1}{(1-u)^{2}}$
-which implies $\pi(u)=\Theta(1/|u|^{2})$.
-A.8 Proof of Lemma 3.6
-Function $\Pi$ is a convex function because, by assumption, $\pi$ is an increasing function. By (8 and Jensen’s inequality, we have
+Note that these conditions are equivalent to those for the binary case in (9) and (10) except for an additional factor $2$ in the first of the last above inequalities. ∎
+A.7 Proof of Lemma 3.2
+Function $\Pi$ is a convex function because, by assumption, $\pi$ is an increasing function. By (4 and Jensen’s inequality, we have
 $\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\tilde{\ell}(\theta_{t})\right]$ $=$ $\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\Pi(\ell(x_{t},\theta_{t}))\right]$
@@
 where the last inequality holds because $\Pi^{-1}$ is a concave function, and hence, it is a subadditive function.
-A.9 Proof of Theorem 3.7
-Under assumptions of the theorem, by Theorem 6.3 Bubeck (2015),
+A.8 Proof of Theorem 3.3
+Under assumptions of the theorem, by Theorem 6.3 Bubeck [2015],
 $\mathbb{E}\left[\tilde{\ell}(\bar{\theta}_{n})\right]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\tilde{\ell}(\theta_{t})\right]\leq\inf_{\theta}\tilde{\ell}(\theta)+\sqrt{2}S\sigma_{\pi}\frac{1}{\sqrt{n}}+LS^{2}\frac{1}{n}.$
 Combining with Lemma 3.2, we obtain the assertion of the theorem.
-A.10 Proof of Corollary 3.8
+A.9 Proof of Corollary 3.4
 Lemma A.11.
@@
 $\Pi^{-1}(y)\leq 2\sqrt{y}\hbox{ for }y\in[0,(3/4)^{2}].$
-Proof
+Proof.
 We consider
@@
 This completes the proof of the corollary.
-A.11 Proof of Lemma 3.10
+A.10 Proof of Theorem 3.6
 To simplify notation, we write $\ell_{t}(\theta)\equiv\ell(x_{t},y_{t},\theta)$, $\gamma_{t}=\gamma(x_{t},y_{t},\theta_{t})$ and $\pi_{t}=\pi(x_{t},y_{t},\theta_{t})$.
@@
 Next, note
-$\mathbb{E}[||\theta_{t+1}-\theta^{*}||^{2}\mid x_{t},y_{t},\theta_{t}]$ $=$ $||\theta_{t}-\theta^{*}||^{2}-2\mathbb{E}[z_{t}\mid x_{t},y_{t},\theta_{t}]\nabla_{\theta}\ell_{t}(\theta_{t})^{\top}(\theta_{t}-\theta^{*})+\mathbb{E}[z_{t}^{2}\mid x_{t},y_{t},\theta_{t}]||\nabla_{\theta}\ell_{t}(\theta_{t})||^{2}$
+$\mathbb{E}[||\theta_{t+1}-\theta^{*}||^{2}\mid x_{t},y_{t},\theta_{t}]$
+$=$ $||\theta_{t}-\theta^{*}||^{2}-2\mathbb{E}[z_{t}\mid x_{t},y_{t},\theta_{t}]\nabla_{\theta}\ell_{t}(\theta_{t})^{\top}(\theta_{t}-\theta^{*})+\mathbb{E}[z_{t}^{2}\mid x_{t},y_{t},\theta_{t}]||\nabla_{\theta}\ell_{t}(\theta_{t})||^{2}$
 $=$ $||\theta_{t}-\theta^{*}||^{2}-2\gamma_{t}\pi_{t}\nabla_{\theta}\ell_{t}(\theta_{t})^{\top}(\theta_{t}-\theta^{*})+\gamma_{t}^{2}\pi_{t}||\nabla_{\theta}\ell_{t}(\theta_{t})||^{2}$
@@
 $\leq$ $\frac{\rho\beta}{c\kappa}(\mathbb{E}[\ell(x,y,\theta^{*})]-\mathbb{E}[\inf_{\theta}\ell(x,y,\theta)])+\frac{1}{2c\kappa}||\theta_{1}-\theta^{*}||^{2}\frac{1}{n}.$
-A.12 Proof of Condition (12) for $\pi(x,y,\theta)=\zeta(x,y,\theta)^{\eta}$
-It suffices to show that under given conditions, (11) satisfies condition (10). Using (11), condition (10) can be written as follows
-$\left(\frac{\ell(x,y,\theta)-\inf_{\theta^{\prime}}\ell(x,y,\theta^{\prime})}{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}\right)^{\eta}\geq\frac{1}{2(1-c)}\min\left\{\beta,\rho\beta\frac{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}{\ell(x,y,\theta)-\inf_{\theta^{\prime}}\ell(x,y,\theta)}\right\}^{1-\eta}.$ (23)
-In the inequality (23), by (20), the left-hand side is at least $(1/(2L))^{\eta}$ and clearly the right-hand side is at most $\beta^{1-\eta}/(2(1-c))$. Hence, it follows that it suffices that
-$\left(\frac{1}{2L}\right)^{\eta}\geq\frac{1}{2(1-c)}\beta^{1-\eta}.$
-A.13 Sampling with Stochastic Polyak’s Step Size for Linear Binary Classifiers
+A.11 Proofs of Corollaries 3.7 and 3.8
 For linear classifiers,
@@
 which is increasing in $u$ for $u\leq 0$ and is decreasing in $u$ otherwise. Note that $h(u)=(1-e^{-\ell(u)})^{2}/\ell(u)$.
-We next note some properties of function $h$ which allow us to establish convergence rate guarantees for some loss and uncertainty-based sampling. See Figure 5 for a graphical illustration.
+A.11.1 Proof of Corollary 3.7
+We first note the following lemma, whose proof is provided in Appendix A.12.
 Lemma A.12.
@@
 Furthermore $h(u)\sim 1-\sigma(u)$ for large $u$.
-By Lemma A.7, the condition of Lemma 3.10 is satisfied with $\rho=1/(2L)$, by sampling proportional to absolute error loss $\pi^{*}(u)=\omega(1-\sigma(u))$ with $\beta/(4(1-c)L^{\prime})\leq\omega\leq 1$.
+See See Figure 4, left, for a graphical illustration.
+By Lemma A.12, condition (6) in Theorem 3.6 is satisfied with $\rho=1/(2L)$, by sampling proportional to absolute error loss $\pi^{*}(u)=\omega(1-\sigma(u))$ with $\beta/(4(1-c)L^{\prime})\leq\omega\leq 1$.
 Figure 4: Upper bounds for function $h$ defined in (20): (left) bound of Lemma A.12, (right) bounds of Lemma A.13.
+A.11.2 Proof of Corollary 3.8
+We have the following lemma, whose proof is provided in A.13.
 Lemma A.13.
-Function $h$, defined in (13), satisifies, for every fixed $a\in(0,1/2]$,
+Function $h$, defined in (20), satisfies, for every fixed $a\in(0,1/2]$,
 $h(u)\leq\frac{1}{H(a)+(1-a)|u|}\hbox{ for all }u\in\mathbb{R}$ (22)
 where $H(a)=a\log(1/a)+(1-a)\log(1/(1-a))$. Furthermore, $h(u)\sim 1/|u|$ as $u$ tends to $\infty$- ∞.
-By Lemma A.8, it follows that condition of Lemma 3.10 is satisfied by uncertainty sampling according to
+See See Figure 4, right, for a graphical illustration.
+By Lemma A.13, it follows that condition (6) in Theorem 3.6 is satisfied by uncertainty sampling according to
 $\pi^{*}(u)=\frac{\beta}{2(1-c)}\min\left\{\rho R^{2}\frac{1}{H(a)+(1-a)|u|},1\right\}.$
-A.13.1 Proof of Lemma A.7
+A.12 Proof of Lemma A.12
 We need to prove that for every $u\in\mathbb{R}$,
@@
 which goes to $1$ as $u$ goes to infinity.
-A.14 Proof of Lemma A.8
+A.13 Proof of Lemma A.13
 We first consider the case $u\leq 0$. Fix an arbitrary $v\leq 0$. Since $u\mapsto\log(1+e^{u})$ is a convex function it is lower bounded by the tangent passing through $v$, i.e.
@@
 We next consider the case $u\geq 0$.
-It sufficies to show that for every $u\geq 0$, $h(u)\leq h(-u)$, and hence the upper bound established for the previous case applies.
+It suffices to show that for every $u\geq 0$, $h(u)\leq h(-u)$, and hence the upper bound established for the previous case applies.
 The condition $h(u)\leq h(-u)$ is equivalent to
@@
 which clearly holds true as both $1/(1+e^{u})^{2}$ and $-u/\log(1+e^{-u})$ go to $1$ as $u$ goes to $\infty$- ∞.
-A.15 Polyak’s Step Size for Multi-class Classification
+A.14 Convergence Conditions for $\pi(x,y,\theta)=\zeta(x,y,\theta)^{\eta}$
+It suffices to show that under given conditions, the sampling probability function satisfies condition (6).
+Using the definition of the sampling probability function, condition (6) can be written as follows
+$\left(\frac{\ell(x,y,\theta)-\inf_{\theta^{\prime}}\ell(x,y,\theta^{\prime})}{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}\right)^{\eta}\geq\frac{1}{2(1-c)}\min\left\{\beta,\rho\beta\frac{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}{\ell(x,y,\theta)-\inf_{\theta^{\prime}}\ell(x,y,\theta)}\right\}^{1-\eta}.$ (23)
+In the inequality (23), by (17), the left-hand side is at least $(1/(2L))^{\eta}$ and clearly the right-hand side is at most $\beta^{1-\eta}/(2(1-c))$. Hence, it follows that it suffices that
+$\left(\frac{1}{2L}\right)^{\eta}\geq\frac{1}{2(1-c)}\beta^{1-\eta}.$
+A.15 Uncertainty-based Sampling for Multi-class Classification
 We consider multi-class classification according to prediction function
@@
 $\frac{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}{\ell(x,y,\theta)}\leq 2||x||^{2}h(u(x,y,\theta))$
-where function $h$ is defined in (13). Hence, condition of Lemma 3.10 holds under
+where function $h$ is defined in (20). Hence, condition of Theorem 3.6 holds under
 $\pi(u)\geq\frac{\beta}{2(1-c)}\min\left\{2\rho R^{2}h(u),1\right\}.$
@@
 $h^{*}(g)=\frac{1}{H(a)+(1-a)\max\{g-\log(k-1),0\}},$
-satisfies condition of Lemma 3.10.
+satisfies condition of Theorem 3.6.
 We next provide proofs for assertions made above. The loss function is assumed to be the cross-entropy loss function, i.e.
@@
 $\frac{||\nabla_{\theta}\ell(x,y,\theta)||^{2}}{\ell(x,y,\theta)}$
-which is plays a key role in the condition of Lemma 3.10.
+which is plays a key role in the condition of Theorem 3.6.
 It holds
@@
 $h^{*}(g)=\left\{\begin{array}[]{ll}\frac{1}{H(a)}&\hbox{ if }g\leq\log(k-1)\\ \frac{1}{H(a)-(1-a)\log(k-1)+(1-a)g}&\hbox{ if }g>\log(k-1).\end{array}\right.$
-A.16 Further Details on Experimental Setup
-Figure 6: Average cross entropy loss as a function of labeling cost for different sampling methods.
-A.16.1 Hyperparameter tuning
-We used the Tree-structured Parzen Estimator (TPE) (Bergstra et al., 2011) algorithm in the hyperopt package (Bergstra et al., 2013) to tune the relevant hyperparameters for each method and minimise the average progressive cross entropy loss.
+Appendix B Appendix: Additional Material for Numerical Experiments
+B.1 Further Details on Experimental Setup
+Hyperparameter Tuning
+We used the Tree-structured Parzen Estimator (TPE) [Bergstra et al., 2011] algorithm in the hyperopt package [Bergstra et al., 2013] to tune the relevant hyperparameters for each method and minimize the average progressive cross entropy loss.
 For Polyak absloss and Polyak exponent we set the search space of $\eta$ to $[0.01,1]$ and the search space of $\rho$ to $[0,1]$.
 Note that the values of $\eta$ and $\rho$ influence the rate of sampling.
-In line with the typical goal of active learning, we aim to learn efficiently and minimise loss under some desired rate of sampling.
+In line with the typical goal of active learning, we aim to learn efficiently and minimize loss under some desired rate of sampling.
 Therefore, for every configuration of $\eta$ and $\rho$ we use binary search to find the value of $\beta$ that achieves some target empirical sampling rate.
@@
 In the hyperparameter tuning we set the target empirical sampling rate to 50%.
-A.16.2 Related to Figure 2
-We similarly used binary search to find the value of $\beta$ that correspondingly achieves the two target values 1% and 10% with Polyak power function, while using the values of $\eta$ and $\rho$ that were optimised for a sampling rate of 50%.
-Therefore, our findings of the gains achieved for selective sampling according to stochastic Polyak’s step size are likely conservative since $\eta$ and $\rho$ were not optimised for specifically these sampling rates.
-A.16.3 Related to Figures 3 and 4
-We compare Polyak absolute loss sampling to absolute loss sampling and random sampling.
+Compute Resources
+All experiments were performed on a single machine with 72 CPU cores and 228 GB RAM.
+It took us around 2,000 seconds to complete a training run for an AWS-PA with an absloss estimator on Mushrooms dataset, our slowest experiment.
+The training runs for other datasets and algorithms were considerably faster.
+B.2 Further Details on Numerical Experiments with Different Algorithms
+In Section 4, we presented numerical results for comparing AWS-PA with other algorithms. These results are shown in Figure 1. Below are some additional details for these experiments.
+Tuning Sampling Rate
+In Figure 1 we compare Polyak absolute loss sampling to absolute loss sampling and random sampling.
 In this setting we have no control over the sampling rate of absolute loss sampling.
-Hence, we first run absolute loss sampling to find an empircal sampling rate of 14.9%.
+Hence, we first run absolute loss sampling to find an empirical sampling rate of 14.9%.
 We then again use binary search to find the value of $\beta$ to match this sampling rate with Polyak absolute loss sampling.
-Again, this setup is conservative with respect to the gains of Polyak absolute loss sampling as $\eta$ and $\rho$ were optimised for a sampling rate of 50%.
+Again, this setup is conservative with respect to the gains of Polyak absolute loss sampling as $\eta$ and $\rho$ were optimized for a sampling rate of 50%.
+Sampling Efficiency of AWS-PA
+In Figure 1 we had demonstrated on various datasets that AWS-PA leads to faster convergence than the traditional loss-based sampling Yoo and Kweon [2019].
+Figure 5 presents results as a function of the number of sampled instances, i.e., the number of labeled instances that were selected for training (i.e., cost).
+This contrasts Figure 1, which showed on the X-axis the total number of iterations.
+The results confirm that sampling with AWS-PA not only leads to faster convergence than traditional loss-based sampling when expressed in terms of number of iterations, but also when expressed in the number of sampled instances.
+Figure 5: Average cross entropy loss as a function of labeling cost for different sampling methods.
+AWS-PA Results on a Holdout Test Set
+The results in Figure 1 were obtained using a progressive validation Blum et al.
+[1999] procedure where the average loss is measured during an online learning procedure where for each instance the loss is calculated prior to the weight update.
+Figures 6 and 7 show that our finding that AWS-PA leads to faster convergence than traditional loss-based sampling and than random sampling also holds true on a separate hold out test set.
 Figure 6: Average cross entropy loss on a hold-out testing set for different sampling methods.
 Figure 7: Test accuracy for different sampling methods.
-A.17 Additional Numerical Results
-Below are some additional empirical experiments conducted on the same mushroom classification dataset (Chang and Lin, 2011) described in Section 4.
-The training loss as a function of the number of labeled instances that were selected for training (i.e., cost) is presented in Figure 6.
-Performance of different sampling methods on the hold-out testing set is presented in Figure 7 (average test loss) and Figure 8 (test accuracy).
-As can be seen from the figures, the relative performance of different sampling methods is consistent across different metrics.
+B.3 Further Details on the Robustness of AWS-PA to Loss Estimation
+In Section 4, we presented numerical results for comparison of the training loss achieved by our AWS-PA algorithm using the ground truth absolute error loss and estimated absolute error loss.
+These results are shown in Figure 2.
+Here, we provide more details for the underlying setup of experiments, and the number of sampled points.
+Details on the Absloss Estimator
+For the experiments in Figure 2 we use a separate Random Forest (RF) regressor which estimates absolute error loss based on the same set of features as the target model with an addition of the target’s model prediction as an extra feature.
+The estimator is retrained on every sampling step using the labeled points observed so far.
+We used the scikit-learn implementation of the RF regressor and manually tuned two hyperparameters for different datasets: (a) number of tree estimators (b) number of "warm-up" steps during which we sample content with a constant probability until we collect enough samples to train an RF estimator.
+We parallelized training of the RF estimator across all available CPU cores and used default values for all other hyperparameters.
+The statistics of the absloss as well as the parameters of the RF estimators for different datasets are summarized in Table 2.
+From the table, we note that the mean ground truth values of the absloss are largely in line with the mean estimated absloss.
+This suggests that it is possible to train absloss estimator with low bias or is even unbiased.
+Sampling efficiency of the absloss estimator
+In Figure 8 we compare the cost of sampling based on the ground truth absolute loss versus sampling based on the estimated absloss.
+We note that in 4 out 6 datasets, the sampling cost closely matches that of sampling based on the ground truth absloss.
+However, in one of the cases (Splice) the sampling cost is lower and in one of the cases (Credit) it is higher than the baseline.
+Table 2: Hyperparameters of the absloss estimator and the comparison of the mean of ground truth and the mean of estimated absolute loss values.
+Dataset Number of trees Warm-up steps Mean absloss Mean estimated absloss
+Mushrooms 25 1 0.100 0.104
+MNIST 3s vs 5s 25 50 0.089 0.087
+Parkinsons 100 5 0.448 0.443
+Splice 100 25 0.163 0.160
+Tictactoe 100 1 0.435 0.416
+Credit 100 50 0.277 0.294
+(a) Mushrooms dataset
+(b) MNIST classification of 3s vs 5s
+(c) Parkinsons dataset
+(d) Splice dataset
+(e) Tictactoe dataset
+(f) Credit dataset
+Figure 8: Sampling efficiency for sampling based on the ground truth of absolute loss v.s. on estimated absolute loss.
+B.4 Additional experiments
+Experiments with different sampling rates
+Loizou et al.
+[2021] demonstrated that stochastic gradient descent with a step size corresponding to their stochastic Polyak’s step size converges faster than gradient descent.
+Figure 9 illustrates that these findings extend to scenarios where we selectively sample from the dataset rather than training on the full dataset, and the step size is according to stochastic Polyak’s step size only in expectation.
+Figure 9: Average cross-entropy progressive loss of Polyak’s step size compared to SGD with constant step size, for 1% and 10% sampling from the mushrooms data.
+To perform these experiments, similarly to the procedure described in Appendix B.1, we used binary search to find the value of $\beta$ that correspondingly achieves the two target values 1% and 10% with Polyak power function, while using the values of $\eta$ and $\rho$ that were optimised for a sampling rate of 50%.
+Therefore, our findings of the gains achieved for selective sampling according to stochastic Polyak’s step size are likely conservative since $\eta$ and $\rho$ were not optimised for specifically these sampling rates.
+Experiments with synthetic absloss estimator
+We simulate a noisy estimator of the absolute error loss in AWS-PA.
+We model an unbiased noisy estimator ^$\hat{\ell}_{\mathrm{abs}}$ of the absolute error loss $\ell_{\mathrm{abs}}\in[0,1]$ as a random variable following the beta distribution, denoted as ^$\hat{\ell}_{\mathrm{abs}}\sim\mathrm{Beta}(\alpha,\beta)$, where $\alpha$ and $\beta$ are parameters set to ensure $\mathbb{E}[\hat{\ell}_{\mathrm{abs}}]=\ell_{\mathrm{abs}}$.
+The variance of the noise can be controlled by the tuning parameter $\alpha$, given by
+$\mathrm{var}[\hat{\ell}_{\mathrm{abs}}]=\frac{\ell_{\mathrm{abs}}(1-\ell_{\mathrm{abs}})}{\alpha+\ell_{\mathrm{abs}}}.$
+Figure 10 shows that the convergence results are robust against estimation noise of the absolute error loss for a wide range of values for $\alpha\geq 1$.
+Figure 10: Robustness of the proposed sampling approach with adaptive Polyak’s step size for different variance $\mathrm{var}[\hat{\ell}_{\mathrm{abs}}]=\ell^{2}_{\mathrm{abs}}(1-\ell_{\mathrm{abs}})/(\alpha+\ell_{\mathrm{abs}})$ noise levels of absolute error loss estimator: (low) $\alpha=100$, (medium) $\alpha=2.5$, and (high) $\alpha=1$.

### v4

@@
 There is a large body of work on convergence of SGD algorithms, e.g.
-see Bubeck (2015); Nesterov (2018).
+see Bubeck (2015) and Nesterov (2018).
 These results are established for SGD algorithms under either constant, diminishing or adaptive step sizes.
@@
 (2021), studied SGD with the stochastic Polyak’s step size, depending on the ratio of the loss and the squared gradient of the loss of a point.
-Our work is different as we consider convergence of SGD under sampling of points.
+Our work proposes an adaptive-window sampling algorithm and provides its convergence analysis, with the algorithm defined as SGD with a sampling of points and an adaptive step size update that conform to the stochastic Polyak’s step size in expectation.
+This is unlike to the adaptive step size SGD algorithm by Loizou et al.
+(2021) which does not use sampling.
 1.1 Summary of our Contributions
@@
 2 Problem Statement
-We consider the projected SGD algorithm defined as: given an initial value $\theta_{1}\in\Theta$,
+We consider the setting of streaming algorithms where a machine learning model parameter $\theta_{t}$ is updated sequentially, upon encountering each data point, with $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\in\mathcal{X}\times\mathcal{Y}$ denoting the sequence of data points with the corresponding labels, assumed to be independent and identically distributed with distribution $\mathcal{D}$.
+Specifically, we consider the class of projected SGD algorithms defined as: given an initial value $\theta_{1}\in\Theta$,
 $\theta_{t+1}=\mathcal{P}_{\Theta_{0}}\left(\theta_{t}-z_{t}\nabla_{\theta}\ell(x_{t},y_{t},\theta_{t})\right),\hbox{ for }t\geq 1$ (1)
@@
 Unless specified otherwise, we consider the case $\Theta_{0}=\Theta$, which requires no projection.
-For binary classification tasks, we assume $\mathcal{Y}=\{-,1,1\}$.
+For binary classification tasks, we assume $\mathcal{Y}=\{-1,1\}$.
 For every $t>0$, we define $\bar{\theta}_{t}=(1/t)\sum_{s=1}^{t}\theta_{s}$.
+By defining the distribution of the stochastic step size $z_{t}$ in Equation (1) appropriately, we can accommodate different active learning and data subset selection algorithms.
+In the context of active learning algorithms, at each step $t$, the algorithm observes the value of $x_{t}$ and decides whether or not to observe the value of the label $y_{t}$.
+The value of $z_{t}$ determine whether or not we observe the label $y_{t}$.
+Deciding not to observe the value of the label $y_{t}$ implies the step size $z_{t}$ of value zero (not updating the machine learning model).
 For the choice of the stochastic step size, we consider two cases: (a) Constant-Weight Sampling: a Bernoulli sampling with a constant step size, and (b) Adaptive-Weight Sampling: a sampling that achieves stochastic Polyak’s step size in expectation.
@@
 Then, for any initial value $\theta_{1}$ such that $||\theta_{1}-\theta^{*}||\leq S$ and $\{\theta_{t}\}_{t>1}$ according to algorithm (1) with $\gamma=1/R^{2}$,
-$\mathbb{E}\left[\ell(yx^{\top}\bar{\theta}_{n})\right]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(y_{t}x_{t}^{\top}\theta_{t})\right]\leq\frac{R^{2}S^{2}}{\beta}\frac{1}{n}.$
+$\mathbb{E}\left[\ell(yx^{\top}\bar{\theta}_{n})\right]\leq\mathbb{E}\left[\frac{1}{n}\sum_{t=1}^{n}\ell(y_{t}x_{t}^{\top}\theta_{t})\right]\leq\frac{R^{2}S^{2}}{\beta}\frac{1}{n},$
+where $(x,y)$ is an independent sample of a labeled data point from $\mathcal{D}$.
 Moreover, if the sampling is according to $\pi^{*}$, then the expected number of sampled points satisfies
@@
 Condition (6) also holds for an uncertainty-based sampling in multi-class classification, as shown in Appendix A.15.
-4 Numerical Results
 (a) Mushrooms dataset
-(b) MNIST classification of 3s vs 5s
+(b) MNIST binary classification of 3s vs 5s
 (c) Parkinsons dataset
@@
 Our proposed algorithm outperforms the baselines in most cases.
+4 Numerical Results
 In this section we evaluate our AWS algorithm, defined in Section 3.2.
@@
 Here we demonstrate convergence on real-world datasets and compare with other algorithms.
+The implementation of AWS-PA algorithm along with all the other code run the experimental setup that is described in this section is available at https://www.github.com/facebookresearch/AdaptiveWeightSampling.
 We use a modified version of the mushroom binary classification dataset (Chang and Lin, 2011) that was used by Loizou et al.
@@
 (a) Mushrooms dataset
-(b) MNIST classification of 3s vs 5s
+(b) MNIST binary classification of 3s vs 5s
 (c) Parkinsons dataset
@@
 5 Conclusion
-In this paper, we have provided convergence rate guarantees for both loss and uncertainty-based active learning algorithms under various assumptions.
+We have provided convergence rate guarantees for loss and uncertainty-based active learning algorithms under various assumptions.
 Furthermore, we introduced the novel Adaptive-Weight Sampling (AWS) algorithm that combines sampling with an adaptive size, conforming to stochastic Polyak’s step size in expectation, and demonstrated its convergence rate guarantee, contingent on a condition related to the sampling probability function.
-For future research, it would be interesting to apply the methodologies presented in this paper to obtain additional convergence rate results.
+For future research, it would be interesting to establish tight convergence rates for the training loss function and the sampling cost, especially comparing policies using sampling with a constant probability with those using adaptive loss-based sampling probabilities.
+It would be interesting to explore adaptive-weight sampling algorithms with adaptive sizes different than those studied in this paper.
 Additionally, exploring a theoretical study on the impact of bias and noise in the loss estimator, used for evaluating the sampling probability function, on the convergence properties of algorithms could open up a valuable avenue for investigation.
@@
 (a) Mushrooms dataset
-(b) MNIST classification of 3s vs 5s
+(b) MNIST binary classification of 3s vs 5s
 (c) Parkinsons dataset
@@
 Figure 8: Sampling efficiency for sampling based on the ground truth of absolute loss v.s. on estimated absolute loss.
+Alternative absloss estimators
+In Figure 2 we shared results of AWS-PA using a Random Forest regressor to estimate the absolute loss.
+Figure 9 shows the results the credit dataset of an otherwise identical experimental setup where we have replaced the Random Forest regressor absloss estimator with an MLP neural network.
+Figure 9: Credit dataset with an MLP neural network loss estimator.
 B.4 Additional experiments